A Bibliography of Publications by, and about, Charles Babbage

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

04 April 2024 Version 1.44

Abstract

This bibliography records publications of Charles Babbage.

Title word cross-reference

#3 [Her99].

0 [Tee94]. 1 [Bab27d, Bab31c, Bab15]. 108 000 [Bab31c, Bab15]. 108000 [Bab27d]. 1791 + 200 = 1991 [Sti91]. $19.95 [Dis91]. $21.50 [Mad86]. $25 [O’H82]. $26.50 [Enr80a, Enr80b]. $27.95 [L.90]. $28 [Hun96]. $35.00 [Ano91c]. $37.50 [Ano91c]. $45.00 [Ano91c]. q [And99a, And99b].

-analogs [And99a, And99b].

0 [Bar96, CK01b]. 0-201-50814-1 [Ano91c]. 0-262-01121-2 [Ano91c]. 0-262-12146-8 [Ano91c, Twe91]. 0-262-13278-8 [Twe93]. 0-262-14046-2 [Twe92]. 0-262-16123-0 [Ano91c]. 0-316-64847-7 [Cro04b, CK01b]. 0-571-17242-3 [Bar96].

1 [Bab97, BRG^{+}87, Mar25, Mar86, Rob87a, Rob87b, Tur91]. 1-85196-005-8 [Twe89b]. 100th [Sen71]. 108-bit [Bar00]. 1784 [MB89]. 1792/1871 [Ynt77]. 17th [Hun96]. 1800s [Mar08]. 1800s-Style [Mar08]. 1828 [Bab29a]. 1835 [Van83]. 1851 [Bab51a, CK89d, CK89i, She54, She60]. 1852 [Bab69]. 1853 [She54, She60]. 1871 [Ano71b, Ano91a]. 1873 [Dod00]. 18th [Bab29a]. 1947 [Ano48]. 1961 [Pan63, Wil64]. 1990 [CW91]. 1991 [Ano90, GG92a]. 19th [Gre06, Gre01, GST01, Swa03, Syd09]. 19th-century [Gre01, Syd09].

2 [Ano08, Bab35a, BL53, Don92, Swa96b, Swa05]. 2.1 [DH94]. 200th [Ano90]. 2016 [Ano16]. 2017 [Ano17]. 2019 [Ano18a, Ano18b, Ano18c]. 2022 [Ano21a]. 2023 [Ano22]. 20th [BHB^{+}89, CW91]. 21st [MN88]. 28a [Wil00c, Bro00].

3 [Bab97, Bar96]. 35.95 [Twe92].

4 [Bro86, Seg08]. 40.50 [Twe91]. 44.95 [Twe93].

500 [Twe89b]. 5000 [BRG^{+}87].

7 [CK01b]. 75th [JL11]. 7880s [WSC^{+}88]. 7950s [WSC^{+}88].

85a [Ber92b]. 87 [Bro86]. 87-872-0086-4 [Bro86].

9. [Bab67a]. 90s [Car69]. 978 [Seg08]. 978-88-02-07715-4 [Seg08].

A. [VBCC83]. abandonment [Tur91]. Abraham [Bab35a]. Abstract [BH30, Bab30e, Bab30a, Bab30d, Bab30c, Bab30h, Bab34a]. academies [BBKC32]. Academy [Seg03]. according [BdL80]. Account [Bab16b, Bab17a, BH25, Bab29a, BH30, BCKS^{+}83, Fra81, Bab38c, Byr38, WB76]. Accounting [Com36]. Acid [Bab43b]. ACM [Ano91c, Bel84, BHB^{+}89]. Act [BH25, BH30, BW00, Wil00b]. Action [Bab26b, Bab30d, Bab56d, Wil64, NJ05]. Ada [AFB^{+}89, HG79, Mad86, Ess14a, Ess14b, EB19, FFT^{+}03, FF03b, HR16, Hoo12, KT99, KM12, Lea20, Lov12, Men21a, Men21b, Swa11a, Swa13]. Adam [O’B93]. Added [Bab16b, Byr38]. Addison [Ano91c]. Addison-Wesley [Ano91c]. Addition [Bab43a]. Additions [MU13]. Addressed [Bab56b, Bab69, Bab56c, Bab89g]. Adelle [OLCJ91]. Admiral [Fra93]. Advancement [MT81]. Advantage [Bab32b, Bab32a, Bab33a, Bab35b, Bab56e]. advantages [Bab34g]. Adventures [Pad15, Sus11]. Advertisement [Ran83]. advisability [Mer82]. affect [Tuk72]. AFIPS [WSC^{+}88]. after [Ano91e, HR03]. afterlife [Cla03]. afterwards [Byr38]. Age [Cur12, Ess04, WSC^{+}88, Ess14a, Ess14b, HR16, CW91, Hun96]. Agency [Swa89, Fis17, Rei07]. agents [Coo05]. Agusta [Men21a]. Aiken [BHW^{+}90, Coh88]. Air [WSC^{+}88, Sus11]. Airy [Swa03]. Alan [Sus11, Swa13, THWV88]. Aldrich [Ano00, Twe93]. Alfred [Nim08]. Algebra [Bec80, Dub77]. algebraic [DR92a, DR92b, NZ06]. algébrique [DR92a]. Algol [HP16]. Algorithm [Bar00, Ess14a]. Algorithmic [GG92b]. alienation [Wen09]. Aliens [Kil10]. Allan [Wil00c]. alleged [BM70]. alliance [Bla02]. also [Byr38]. Alternative [Sus94]. Ambivalence [Spe99]. America [Ano08, Bla02]. American [All91, Ano00, Fri05]. Ames [L.90]. amico [Bul68]. analitica [For75]. analogs [And99a, And99b]. Analogy [Bab17c, Bab30c]. Analysis [Bab17c, Bab23a, Bab30c, Bab56a, Ash96, NJ05, Pat60, Rim09]. Analytical [Ano77, Bro90, Bro00, Enr83, For75, GC10, Lov43, Pay66, Wil00c, Bab10, Bab89a, CK89l, Gre01, Ort07a, WB76, Ald91, Bab43a, Bab49, Bab89a, Bab82b, Ber64, Ber81, Bro82, Bro98, DH90, DH94, DH98, DH04, DR92a, DR92b, Har49b, Men21a, Men21b, Mer82, Pad17, Ran82a, Seg03]. Analytique [Men43, DR92a]. ancient [Pin06]. Andrea [Seg08]. Anecdotes [Nag80a, Roe09, SHH81, THWV88, TW87]. Angeles [Twe92, Twe93]. anglais [BdL80]. anglaise [DR92a]. Animals [Bab60a]. Ann. [Ber92b]. Anniversary [Bab56b, BHB^{+}89, Smi71, Ano90, Sen71]. Annotated [Ran79]. Annual [Dod00]. Annuities [Far64]. Anthony [Asp82, Enr88, Hil85, O’D85a, O’H82, Ros90, Smi83, Tro84, W.86, CK89a, Ric83]. Anticipation [BHW^{+}90]. Antique [Ano91a, JC90]. apercus [Poi82]. APL [Fra84, Fra85, Bro86, Vam85, Wil86, Wor87]. Apparatus [Ric39]. appendix [BR89, Byr38, CK89q, CK89r, BL53]. appertain [Byr38]. Apple [CW91]. Application [Bab22b, Bab22c, Bab22d, Bab23a, Bab29d, Byr38, HB13]. Applications [Com36, Cat81, PHB20]. appointed [Mer82, Sto55, Wro56]. appointment [Ric97]. approach [Joh92]. Arago [BH25, BH30]. Archaeological [Ald91]. Architects [Cur12]. architectures [HP16]. Archive [WSC^{+}88]. Archives [WSC^{+}88, All91]. Archon [Mad86]. Art [Bab32b, Bab32a, Bab33a, Bab35b, Bab56e, Bab60a, Sus11, CH02]. Arthur [Asp07, Mis07, WWCP92]. Article [Ano91a]. Articles [Bab34b]. Artifacts [All91]. Arts [Bab29d, HBC^{+}55, All91]. Ascent [CW91]. Asimov [BCKS^{+}83]. Aspects [Sch81]. Aspray [Ano91c, Hun96, L.90, CW91]. Association [Ano91c, Bab39, MT81]. Assurance [Bab24b, Bab26a, Bab67b, Sib95, Bab89b, CK89o, CK89p, CK94]. Astronomical [Bab22c, Bab47c, Bai89, Mar22]. astronomy [Ash94]. Atanasoff [BHW^{+}90, Smi10]. attending [Nic43a]. Attwood [O’B93]. Augustus [Ric97]. ausgewählte [Dot96]. Australasia [Tee83]. Australia [Tee82]. Author [Arc47]. authors [Bro66]. Autobiografia [Seg08, BV07]. autobiography [Bab91a, BV07]. Automate [Dot96]. Automated [Bul08]. Automatic [GT80, Swa10, HP16, Jon91, Mar25, Mar86, Com47]. Automation [WSC^{+}88]. automatischer [Mar25, Mar86]. Autopsy [Bab91b]. avantages [Bab34g]. Award [Ano21a, Bel84, Ano16, Ano17, Ano18a, Ano18b, Ano18c, Ano21b, Ano21a, Ano22]. awards [Bel84].

B [BRG^{+}87, BHB^{+}89, Mer82]. B-5000 [BHB^{+}89]. Babbage [All91, Ano26, Ano71b, Ano33, Ano80, Ano90, Ano91a, Ano91c, Ano00, Ano16, Ano17, Ano18a, Ano18b, Ano18c, Ano21b, Ano21a, Ano22, Ash02, Asp82, AB86, AFB^{+}89, Asp07, Ben16, Ber92b, Bla91, Boo62, Bow15, BCKS^{+}83, Bro86, BHW^{+}90, Bül88, BH88, Cal68, CK92, CK01b, CK01c, Can66, Car69, CW91, Cro04a, Cro04b, Dot97, Fin52, Enr88, Fah91, For75, Fra63a, Fra63b, Gre06, GG92a, Gul80, Gul81, Her31a, Her57, Hil85, Hor09, Hym87, Lee95, MN88, Mis07, Nof15, O’B93, O’D85a, O’H82, O’N91, Ogb65, OS65, Pan63, Ric83, Schxx, Seg08, Sei95, Sib95, Smi83, SBET85, SHH81, Tro84, Twe89b, Twe91, Twe92, Twe93, Ugl96, Vam85, W.62, W.86, Wil64, Wil00c, Wil86, WSC^{+}88, Wor87, Yos21, Ait16, And99a, And99b, Ano32a]. Babbage [Ano32b, Ano71c, Ano71d, Ano71a, Ano61, Ano74, Ano77, Ano81, Ano82a, Ano82b, Ano91b, Ano91e, Ano91d, Ano93, Ano03, Ano06, Ano08, Ano10a, Arc35, Ash94, Ash96, Bab34b, Bab35e, Bab39, Bab49, Bab56c, Bab56g, BdL80, Bab89b, Bab10, Bab48, Bab82b, Bab82c, BR89, BH89, BCKR89, Bab91b, Bab91a, Bab97, Bab10c, Bai89, Bar00, Bar71, Bar72, Bec80, Ber92a, Ber92b, Bow93, Bra91, Bro82, Bro85, Bro87, Bro91, Bro98, Bro00, Bro92, BRG^{+}87, BWV^{+}88, BHB^{+}89, BHW^{+}90, BDH90, BW00, Bul68, Bul71, Bul08, Byr38, CK88, CK89a, CK89b, CK89e, CK89h, CK89i, CK89j, CK89k, CK89l, CK89m, CK89n, CK89o, CK89p, CK89q, CK89r, CK89s, CK89t, CK89f, CK89g, CK94]. Babbage [CK00, CK01a, CSW90, Cat81, Cha33, Cha01, Coh88, Col06, Col98, Col71, Col90, CM98, Com31, Com46, Coo05, Cro85, Das14, De 06, Dij86, Dod74, Dod00, Dol98, Don92, Dot96, Dot97, Dub77, Dub78, DR92a, DR92b, DS21, Edi92, Fis14, Fra81, Fra84, Fra85, Fra86, Fra93, Fra63b, Fri92, FZZ08, FF03a, FF15, Gar69, GP08, Goo02, GC10, GG79a, GG79b, GG92b, Gre01, Gri06, Gri10, HP16, Hal70, Har49b, Her89, HR03, HH80, HG79, Hym82a, Hym82b, Hym83, Hym87, Hym88, Hym92, Hym02, Jar03, JC90, Jon91, Jon16, Kan94, Kem79, KT99, Kirxx, Kou14, Lar34, Lar89, Lav14, Lew03, Lew07, Lin87, Lin90, Los72, Lov43, Lyo40, Mac07, MÈ78, Mar22, Men43, Men21b, Mer82, MU13, MM52, MM61a, MM61b, Mos83]. Babbage [Mos64a, Mos64b, Mos65, Mul48, Nag80a, Nag80b, Nic43a, Nic43b, Nim08, Nor01, Nud69, O’D85b, OLCJ91, Ort07a, Ort07b, Ozg10, Pad15, Pad17, Par96, PP79, Poi82, Por07, Pur93, Ran91, Rat52, Rei68, Rei07, Ric39, Ric02, Rit16, Rob87a, Rob87b, Rob93, Roe09, Roj21, Ros88, Ros94a, Ros90, Schxx, Sch02, Sco05, Seg03, Sen71, She54, She60, She06, SU96, Ste92, Sti91, Sus94, Sus11, Swa91b, Swa91c, Swa91d, Swa93, Swa96a, Swa96b, Swa00a, Swa00b, Swa01a, Swa01b, Swa02, SB02, Swa03, Swa04, Swa05, Swa09, Swa10, Swa13, Swa17, Syd09, Tay92, Tee82, Tee83, Tee84, Tee86, Tee94, Tee95, Tom02, THWV88, Tri72, TW87, Tur91, Twe89a, Van80, Van83, VBCC83, VW91, Wil64, Wil75, Wil77, Wil81a, Wil90, Wil91a, Wil91b, Wil91c, Wil92a, Wil95, Wil00a]. Babbage [Wil00b, Wil02, Wil80, Wil81b, Wil87, Wil88, Wil98, WB76, Ynt77, Aga18, Bar96, Bey79a, Bey79b, Bey79c, Enr78, Enr80a, Enr80b, Hil79, Opp71, Rob79, Ros90, Twe89a, For75, GG79b, Gri78, Swa91a].

Babbage/Faraday [GG92a]. Babbages [Dot96]. Babel [Fri92]. Bache [Bab69]. Back [Bra91, CW91, DeF22]. Backus [Ran91]. bad [Thi00]. Baily [Ash94]. Banks [Bab27a, Ano91d]. Bantam [Dis91]. Baptiste [O’B93]. Barlow [Ano26]. Barometer [Bab24a]. Barometrical [HB22]. Barry [CW91]. Bart [Bab22b]. Basic [Hun96]. Basics [CW91]. Battery [Bab55b, Bab55c]. Baum [Mad86, VBCC83]. Be [Bar00, Bab32b, Bab35b, Bab38c, Bab56e, Mor95]. Beaufort [Fra93]. Bebidzh [AMÈ81]. been [Byr38]. Before [L.90, ABCK^{+}90, Bab27a]. Began [Lav14, Nof15, Bow15, Das14]. beginning [Bar71]. beginnings [Lew03]. behavior [Rat52]. Being [Bab10c, Bab89b]. Bell [Bab26f]. Bentham [O’B93, Sco05]. Berlin [Bab29a]. Bernstein [Pay66]. besides [Jon91]. best [Ano11]. Better [O’N91, Kou14]. between [Ano06, Bab17c, Bab30c, Bla02, Gre06, Kan94, KT99]. beyond [Fri92, WSC^{+}88]. Bibliographia [Arc35]. Bibliographic [Yos02]. Bibliographical [Arc48]. Bibliography [Arc47, Ran79, MU13]. Bicentenary [GG92a]. Bicentennial [Ano90]. Biddell [Swa03]. binomial [And99a, And99b]. Bio [Arc48]. Bio-Bibliographical [Arc48]. Biographical [BCKS^{+}83]. Biographies [Cor87, Gri06, Wei88, WWCP92, Wil92b, BHW^{+}90]. Biography [Arc47, But74, Pad17, Smi10]. Birkbeck [Por07]. Birkerød [Wor87]. Birth [Ess04, Fah91, Ano90, Joh91]. Birthday [JL11, VW91]. Births [Bab29c]. Bit [Aug84, Bar00]. Bits [CK01b, BCKS^{+}83]. black [Ros94b]. Blaug [O’B93, O’B93]. BLOG [HMB22]. BNF [BHW^{+}90]. board [Mat90, She54, She60]. body [Bab91b, Mus06]. Bones [Bab60a]. Book [Aga18, Ash02, Asp82, Bar96, Bey79a, Boo62, Bow15, Bro86, CK89a, CK01b, Can66, Car69, Cro04a, Cro04b, Dis91, Enr78, Enr80a, Enr80b, Enr88, Fog01, For75, Fra63a, GG79b, Gri78, Hil79, Hil85, Hun96, Kem79, L.90, Lav14, Lea20, Mad86, O’B93, O’D85a, O’H82, Ogb65, Opp71, OS65, Pan63, Pay66, Ric83, Rob79, Ros90, Seg08, Sib95, Smi83, Sus00a, Sus00b, Sus11, Swa91a, Twe89a, Twe89b, Twe91, Twe92, Twe93, Vam85, W.62, W.86, Wil64, Wil95, Wil00a, Wil86, Wor87, Ano93, OLCJ91, She54, She60, Sin99, AAG^{+}00]. Books [Dis91, Hun96, Mad86, Ano11]. Books/The [Mad86]. Boole [Coo05, GG05]. Boracic [Bab43b]. Boston [WSC^{+}88]. Bowditch [BWV^{+}88, BHW^{+}90, Wil87, Wil88]. Bowers [Kem79]. box [Ros94b]. Brain [Ano00, Cro04a, Cro04b, Hor09, Swa00a, CK01b, Hoo12, Swa01a, Wil00a]. Branches [Bab17c, Bab30c]. breakfast [Sny11]. Breaking [Bab57]. Brewster [Bab32a, Bab33a, Her31b, Kou14]. Brian [Kem79, JL11]. bride [Woo99]. Bridgewater [Bab67a, Bab09, Car69, Bab37a, Bab38a, CK89t, Bro66]. Brief [Ano96, Bow53a, Fri92, O’R12]. Briggs [Byr38]. bringing [Kou14]. Britain [BB36, FRW11, Fre03, Mus06, Rei68]. British [Bab39, SBET85, Bab97, Mac19, MT81, WSC^{+}88]. Brittle [Ano91a, JC90]. Bromley [Wil00c]. Brown [CK01b, Cro04b]. Bruce [Dis91, Spe99, Sus11, Ash02]. Build [Ano00, Cro04a, Cro04b, Swa00a, Swa01b, CK01b, GC10, Swa00b, Swa01a, Swa02, SB02, Wil00a]. Builders [VBCC83]. Building [Swa91b]. built [Ano91e]. Bureau [Ano48]. Burroughs [All91, BRG^{+}87]. Bush [Ano00, Ran82a]. Business [CW91, Ano11, Ash94]. Busts [Arc48]. Butt [O’B93]. Buxton [Enr88]. Byron [FFT^{+}03, FF03b, Ess14a, Ess14b, Woo99, Mad86]. Byronic [Mad86].

C [CW91, Mer82, Sev13, Twe91, MÈ78]. C. [Bab35a, BCKS^{+}83, VBCC83]. CACM [HMB22]. Cairnes [O’B93]. calcolatrici [Tri72]. calcolo [Los72]. calcul [Mos83]. Calculated [SS57]. Calculating [Aga18, Air56, Ano48, Bab22b, Bab23b, Bab37c, Bab49, Bab10c, Boo62, Bro87, Edi92, Fra63a, Fra63b, Har49a, Har12, Her31a, Jon16, Lar34, Mad86, Mar25, Mar86, MKW92, Nic43a, Nic43b, O’H82, O’N91, Sto55, Swa04, Tom02, Twe93, W.62, Wil64, Wro56, Ash94, Bab34b, Bab35e, Bab55f, Bab89b, Bab82c, Bab89g, Bai89, Bar72, Bro85, Col71, Col90, Dot96, Jon91, Lar89, MM61a, OLCJ91, Pan63, Swa91c, Bab82a, Mar25, Mar86]. Calculation [Bab22c, CW91, Swa03, Jer89, Mos83]. Calculations [DeF22, Byr38]. Calculator [McC97, Mer77, MM61b, Par96, Com47]. Calculators [Thi00]. calculer [Bab35e, Bab55f]. Calculus [Bab15, Bab16a, Bab17c, Bab17d, BHP20, Bab22e, Bab30a, Bab30c, LaC16, PHB20]. Call [Ano21a]. Callett [Byr38]. calumnies [She54, She60]. Cambridge [Ano91c, CK89a, Enr80a, Enr80b, GG92a, Twe91, Twe92, Twe93, BBKC32, Coo05, Enr83, HM02, Wil90]. came [Wil75]. Campbell [Hun96, Twe89a, Twe89b, Sib95]. Campbell-Kelly [Hun96, Twe89a, Twe89b, Sib95]. Capsule [BCKS^{+}83, SBET85, VBCC83]. Card [CW91, Eck40, Eck84]. Carroll [Ait16]. carry [Mar25, Mar86]. cartoonist [Pad17]. Cass [Car69]. casting [Bar71]. Catalogue [Ano90]. catcher [Jon91]. caught [BW00, Wil00b]. Causes [Bab30i, Bab34a, Bab34c, Bab57, Bab47a, BM70, BR89, CK89q, CK89r]. Cayley [Mer82]. CBI [Gul81]. CBQ [Gri06]. Celebrate [Ano82b]. Celebration [OLCJ91]. census [Sco05]. Center [WSC^{+}88, OLCJ91]. Centuries [Car69]. Century [Ano00, AFB^{+}89, CW91, MN88, Cla03, Fre03, Gre06, Gre01, GST01, Her88, Mac19, Hun96, Swa89, Swa03, Syd09].

Certain [Bab34a, Bab34c, Bab47a, Ben16, Rit16]. Certainty [Swa89]. Ch [Bab35e]. Ch. [BdL80]. Chadwick [O’B93]. chair [Ric97]. Chance [Bab21, Bab53]. Chandrasekar [Ano91a]. Change [De 06]. changed [Cha12, Sny11]. changes [Bar96]. channels [Schxx]. chapter [Bab64a, Wel89]. character [Wil91c]. Characters [Bab30j]. Charles [Ano82b, Ano91a, Ano91c, Ano00, Ano16, Ano17, Ano18a, Ano18b, Ano18c, Ano21a, Ano22, AB86, Bey79a, Bey79b, Bey79c, Bla91, Boo62, BHW^{+}90, Bül88, Cal68, CK89a, CK92, CK01b, Can66, Car69, CSW90, CW91, Cro04a, Cro04b, Enr78, Enr80a, Enr80b, Enr88, Fah91, Fra63a, Fra63b, GG79b, Gri78, Gul81, Hil79, Hor09, HG79, Kem79, Kirxx, MN88, Men43, O’B93, O’H82, Ogb65, Opp71, OS65, Rob79, Ros90, Swa91a, Swa09, Twe89a, Twe89b, Twe91, Twe92, Twe93, VBCC83, All91, Ano71b, Ano71c, Ano71d, Ano71a, Ano33, Ano74, Ano80, Ano82a, Ano90, Ano91e, Ano91d, Ano08, Ano21b, Arc35, Ash96, Ash02, Asp07, BBKC32, Bab56c, Bab89b, Bab48, BR89, BH89, BCKR89, Bab91b, Bab91a, Bab97, Bab08, Bab10c, Bar71, Bar72]. Charles [Bow93, Bro82, Bro85, Bro98, Bro92, Bul68, Bul71, Bul08, BH88, CK88, CK89b, CK89e, CK89h, CK89i, CK89j, CK89k, CK89l, CK89m, CK89n, CK89o, CK89p, CK89q, CK89r, CK89s, CK89t, CK89f, CK89g, CK94, CK00, CK01a, CK01c, Cha01, Cla03, Col71, Col90, CM98, De 06, Dij86, Dod74, Dod00, Dol98, Don92, Dub78, DR92a, DR92b, EEM^{+}05, Edi92, Fin52, EB19, Fra63b, Gar69, GP08, Goo02, GG92b, Gre01, Gri06, Gri10, Gul80, Hal70, Har49b, HH80, Hym82a, Hym82b, Hym83, Hym87, Hym88, Hym02, JC90, Jon91, KT99, Lea20, Lew07, Lin87, Lin90, Los72, Lyo40, Mar22, Men21b, Mis07, MM52, MM61a, MM61b, Mos83, Mos64a, Mos64b, Mos65, Mul48]. Charles [Nim08, Nor01, Nud69, O’D85b, O’N91, Ort07b, Ozg10, Par96, Por07, Pur93, Rat52, Ric83, Ric39, Roj21, Ros88, Ros94a, Seg08, Sei95, Sen71, She06, Sti91, SHH81, Sus94, Swa91c, Swa91d, Swa93, Swa96b, Swa00a, Swa00b, Swa01a, Swa01b, Swa02, SB02, Swa03, Swa05, Swa10, Swa13, Syd09, Tay92, Tee82, Tee83, Tee84, Tee86, Tee94, Tom02, Tri72, Tur91, Van80, Wil92a, Wil00a, Wil02, Wil80, Wil81b, WSC^{+}88, Wil98, WB76, Ynt77, Yos21, Asp82, Boo62, Fra63a, Hil85, O’D85a, Pan63, Sib95, Smi83, Tro84, W.62, W.86, Wil64]. Charlz [AMÈ81]. Chas. [Bab69]. Chatto [Twe89b]. Chess [Bab17a]. Christie [Ano26, Ste95]. Chronology [BEG^{+}87]. Chuck [Col98]. cipher [Fra93]. Ciphers [AAG^{+}00, Pin06]. Circle [Bab16b, Jar03]. Circuit [CW91]. Circumstances [Bab29c, Nic43a, Nic43b]. Civil [Bab56c]. Class [Bab26e]. Classes [Bab19b, Bab30h]. Classic [Ano30, Lew21]. Classical [C^{+}09]. Clearing [Bab56a]. Clement [Rob93, WWCP92, Wil92b]. Cliffe [O’B93]. Clifford [Mer82]. clippings [Bab51d]. club [Sny11]. Co [CK01b, Ano91c]. co-published [Ano91c]. code [HP16, Sin99, AAG^{+}00]. Codebreaker [Pin06]. Codes [SBET85, Pin06]. coefficient [And99a, And99b]. cognitive [Gre01]. Cogwheel [Ano00, Cro04a, Cro04b, Swa00a, CK01b, Swa01a, Wil00a]. collaboration [KT99]. Collection [AB86, BM23, Bab32a, Bab33a, Bab10c, All91, Ano91d, Bab32b, Bab35b, Bab56e, Bab89b, Bab82c, PHB20]. Collections [BEG^{+}87]. Collier [Ash02]. COLOSSUS [Ran76, BWV^{+}88]. Coloured [Bab31c]. Column [Sei95]. Come [Com46]. Comes [Mar08, HP16]. coming [Jen85]. commemorate [Ano90]. Comment [Ber92b]. commentary [Joh95]. Comments [Ano81, Ber92a, BRG^{+}87, BEG^{+}87, BWV^{+}88, BHB^{+}89, BHW^{+}90, BDH90, Hym92, Rob87a, Rob87b, Rob93, Wil00b]. commissioners [Ano06, Gre06]. Committee [Mer82, Wro56, Her89, Bab27a, Her57, Sto55, WSC^{+}88]. Common [Bab27b]. Commons [Gre06, Bab27a]. Company [Cro04b, Bab24b]. Comparative [Bab26a, Bab67b, Bab89b, CK89o, CK89p]. compared [Byr38]. competence [Schxx]. competitiveness [OLCJ91]. Complete [Bar00]. complex [Ros88]. complexity [CM11]. Comput. [Ber92b]. computable [Zen13]. Computation [Ano48, Bab22d, Bow53a, Swa10, Eck40, Eck84, MM11, Rat52, Zen13]. Computational [Swa10]. Computations [DeF22]. Computer [AFI72, Ald91, Ano96, Ano00, AFB^{+}89, Atk00, BCKS^{+}83, BRG^{+}87, BEG^{+}87, CW91, Cro04b, DeF22, EB19, Fah91, Fra81, Gig01, Gol72, Gol93, Gri06, Hym87, KT99, Lea20, Lew21, LR67, Mar08, Mat91, Nof15, RR93, Ran82a, Roj21, Sla87, Smi71, Ste95, Swa93, Swa00a, Swa01b, WWCP92, Wil91a, WSC^{+}88, Aga01, Ano74, Ano93, BV07, Bow15, Cer83, Das14, DH90, DH94, DH98, DH04, Ess14b, Hal70, Hun96, Hym82a, Hym82b, Mar22, OLCJ91, Par96, Pat60, SGM07, She06, Smi10, Sta11, Swa00b, Swa01a, Swa02, SB02, Tay92, Wil77, Wil00a, Wil92b, Ano08, BWV^{+}88, CW91, O’D85b, CK01b, Seg08, Cro04a, Hil85, O’D85a, O’H82, Ric83, Smi83, Tro84, W.86, Hun96]. Computerpionier [Hym87]. Computers [Ano91c, Aug84, BRG^{+}87, DB05, Dis91, Plo12, Pur93, Ran73, Ran75, Ran79, Ran82b, ABCK^{+}90, Ber64, Ber81, Fri05, OLCJ91, BWV^{+}88, L.90]. Computing [Ano91c, ABCK^{+}90, Asp07, Bow53b, BEG^{+}87, BHB^{+}89, CK01c, JL11, Mat91, MW80, Mis07, Sev13, Swa11b, Swa13, Twe92, Twe93, Wat12, Wil92a, WSC^{+}88, Wil97, Yos02, Yos21, Ano10a, CK92, HR16, HNW^{+}48, Jon91, MW72, O’R12, Tuk72, Wil85, Ano90, Ano91c, CW91, GT80, MN88, WSC^{+}88, L.90]. concept [Cer83, CW91]. Concise [Cle09]. Conditionals [BCKS^{+}83, Fra81]. conditions [HB13, Mus06]. Conference [AFI72, Ano90, GG92a, OLCJ91]. Conferences [BEG^{+}87]. Congress [Bab29a, OLCJ91]. Congresses [Bab60b]. congruence [FZZ08]. congruences [And99a, And99b]. conjecture [Cat81]. Conn [Mad86]. Connected [Bab21, Bab53]. Connection [Wil87, Wil88]. connections [Tee84, Tee86]. Consciousness [Wat12]. Conservatoire [All91]. consider [Mer82]. consisting [Mer82]. Constable [Wil64]. Constants [Bab32b, Bab32a, Bab33a, Bab35b, Bab56e]. constructing [Mer82]. Construction [Bab10c, Nic43a, Swa91d, Swa05]. Contact [Tom02, Cat81]. containing [Byr38]. contemporary [Jen85]. Contents [Joh95]. contest [Bel84]. Context [Ste84]. Continued [Bab13]. Contributions [Ran82a]. Control [Wil81a, HP16]. Controlled [Com47, OLCJ91]. Copies [Gre06, Ano06]. core [Schxx]. corporation [Bla02, All91]. Correcting [SBET85]. correspondants [Bul71]. correspondence [Ano06, BR89, Bab08, CK89q, CK89r, Gre06, Bab97]. correspondents [Bul71]. could’ve [Col71, Col90]. Council [Wro56, Ano06, Gre06, Sto55]. Countess [HG79, KT99, Men21a]. course [Cha12]. Courtenay [Bab29c]. cow [Jon91]. Cracked [AAG^{+}00]. Cracks [Bar00]. craft [Ada10]. Craig [Ano91c, Twe91]. Crash [Smi06]. Crawford [Ano90]. Created [Isa14, Lew21]. Creation [FF03a, FF15]. Creatively [Fis17]. cross [Bro91]. cross-referenced [Bro91]. Crosthwait [Bab35a]. cruncher [Mat90]. Cryptographic [Bar00]. cryptography [Fra93, Pin06, Sin99, Sta11, BCKS^{+}83]. CS [Ano16, Ano17, Ano18a, Ano18b, Ano18c, Ano21a, Ano22]. Cultural [SU96, Ste84, Ugl96, Bar96]. Culture [Sus00a, Sus00b, Biz03, Cla03, Fri05, Yos21]. cultures [HR16]. Currency [Bab33b]. Currents [Bab56d]. curriculum [Wil90]. Curve [Bar00]. Cyber [Spi02]. Cyberpunk [Sus94]. cyberspace [Cla03]. Cycles [Bab34a, Bab34c, Bab47a]. Cypher [Bab54a, Bab54b, Bab54c, Bab54d, Bab55a, Bro86, Wil86, Wor87, Fra84, Fra85, Vam85].

D. [HG79]. dalla [BV07, Seg08]. Dan [Ack95]. Danish [BRG^{+}87]. dans [Bab31b]. d’aprés [BdL80]. Dasgupta [Bow15, Nof15]. DASK [BRG^{+}87]. Data [Cor87, WSC^{+}88, DB05, HMB22]. datahistoriebok [Ano93]. dates [Ste92]. daughter [Ess14a, Ess14b, Woo99]. David [Bab32a, Bab33a, Her31b, O’B93]. Davy [Bab22b, Bab22a, Bab30j, Kou14]. Dawn [Wat12]. Day [Ano82b]. Days [Hun96]. DC [WSC^{+}88]. death [Sen71]. Debate [Ano81, Ber92a, BRG^{+}87, BEG^{+}87, BWV^{+}88, BHB^{+}89, BHW^{+}90, BDH90, Hym92, Rob87a, Rob87b, Rob93]. Debates [Bre75, Wil00b]. Debuts [Ano08]. December [Fah91]. Decision [BHB^{+}89]. Decline [Bab30i, Bab98, Bab06, BM70, BR89, Bre75, CK89q, CK89r, Her31b]. decomposability [Rei07]. Decyphered [Bab35a]. Dedicated [JL11]. del [BV07, Seg08]. dell’inventore [BV07, Seg08]. Demonstration [Bab19a]. Demonstrations [Bab16b]. Denmark [Bro86, Wor87]. Department [Ano48]. Dependable [JL11]. described [KT99, O’N91]. Describes [Edi92]. describing [Rat52]. Description [Bab38c, Bab43b, Hor09, Ric39, Swa96b]. descriptive [WB76]. Design [Atk00, Wil81a, DS21, MB89]. designed [Ano91e]. desperation [Ric97]. Despite [Mad86]. determinantal [Cat81]. Determination [Bab26e, HB13]. Deutsches [OLCJ91, WSC^{+}88]. Development [MKW92, Rim09, Twe93, Mar25, Mar86, MM61b, Nim08, Ozg10, Wil90]. Developments [Joh92]. Devouring [CW91]. Dickens [Cla03]. Dictionary [Cor87, DG99]. did [Col98, Jon91]. Diecasting [Bar72]. Dies [Bar00]. Difference [Ano30, Ano91c, AFB^{+}89, Bab56c, BCKS^{+}83, Bül88, CW91, Cro85, Dis91, Spe99, Sus94, Sus11, TW87, Twe91, Bab89c, Fah91, Lin87, Lin90, MÈ78, MB89, Par96, Swa91d, Tee94, WB76, Ano91e, Ano08, Atk00, Bab56g, Bro90, BR93, BRG^{+}87, CK89k, Coh88, Cro85, Don92, Fra81, GS91, GS11, Gra71, Her99, Kil10, LL82, Mar22, Par96, Rob87a, Rob87b, Roe09, Swa96b, Swa00b, Swa01b, Swa02, SB02, Swa05, Syd09, Tay92, Tur91, Wil76]. differences [Byr38, HB13]. Different [Bab29c, Bab31c, C.88, Bab31b]. différentes [Bab31b]. Differential [BHP20, LaC16, PHB20]. Differenzmaschine [MB89]. Diffusion [CW91]. Digital [Ano48, Bar00, Bow53b, Gig01, Hun96, Isa14, LR67, Ran73, Ran75, Ran79, Ran82a, Ran82b, Swa13, Wat12, Cer83, Ess14a, HR16, Smi10, All91]. Dionysius [Bla91, O’B93]. Director [Sei95]. Disappearance [Atk00]. Disc [Bab52]. Discourse [NJ05]. discoveries [Kou14]. Discovery [Bab23a, Bab60a]. discussion [HNW^{+}48]. Display [Ano08]. dissertation [Bel84]. Dits [BCKS^{+}83]. Diversity [Atk00]. Diving [Bab26f]. division [Poi82, Rei07]. Dkr [Bro86]. doctoral [Bel84]. Document [Rei68]. Does [BHW^{+}90, LR67]. domino [Mul02]. donates [All91]. Donovan [Sus11]. Dorfman [AFB^{+}89]. Doron [Cro04a, CK01b, Cro04b, O’N91]. Dorothy [Ano00]. Dover [Wil64]. Dr [Mer82]. Dr. [Bab16b, Bab30j, Bab60b, Bab69]. dramatic [Wil91c]. drawing [Mat90]. Drawings [Bab60c]. drawn [Bab56c]. Dream [Com46, Pur93, Swa91b, Wil64, FFT^{+}03, FF03b, Wil75]. Dreams [Sus00a, Sus00b, Fri05]. Drive [Bow93]. Dubbey [Enr80a, Enr80b, HG79, Kem79, Rob79, Bey79a, Bey79b, Bey79c, Enr78, Hil79]. Dudley [THWV88]. Duke [Bab89g]. During [Bab56a, BH25, BH30, Bab52]. Dutch [BWV^{+}88].

each [Tuk72]. Eames [EEM^{+}05]. Early [Ano91c, BEG^{+}87, BWV^{+}88, CW91, WSC^{+}88, Biz03, MT81, Mus06, BWV^{+}88, Dol98]. Earth [Bab56d]. Easily [Bab60c]. Eckert [Smi71]. Eclipse [Bab52]. école [DR92a]. Ecological [BHB^{+}89]. economia [Bab34e, Bab35d]. economic [CM11, Coo05, Rim09]. Economics [O’B93, Bab35d, BdL80, GP08, HR03, Ros94b]. économie [Bab33c, Bab34f, Kan94, BdL80]. économique [Bab34d]. economist [Ozg10, Ros94a]. Economy [Bab35c, Bab71, Bab03a, Bab10a, Ber82, CK89c, Sus00b, Ano32a, Ano32b, BBKC32, Bab32c, Bab32d, Bab33c, Bab33d, Bab34d, Bab34e, Bab34f, Bab46b, Bab63, Bab86, Bab10b, CK89s, Kan94, Lew03]. econophysics [C^{+}09]. ed [CSW90, L.90, Seg08, Twe89b]. Eden [CW91]. Edinburgh [Ano90, Her31b]. edited [Bar96, Twe93, Wil64]. edition [Car69]. Editor [CK89a, Ano91c, Her31b]. educator [Joh92]. Edvard [Ano91c, AFB^{+}89, Bül88, CW91, Lin87, Lin90, Twe91, Arc47]. Edward [O’B93]. Edwin [O’B93]. Effect [Bab33b]. effects [Bab34g]. effets [Bab34g]. efficiency [Ash96]. efficient [Ano10a]. Eighteen [Bar72]. Eighteen-sixties [Bar72]. Eighteenth [Car69]. Eisler [CW91]. Electric [Fri05, Pat60]. Electrical [Bab26d, Bab30e, WSC^{+}88, OLCJ91]. Electronic [Bab08, LR67, Ran82a, Smi71, Swa11b]. Elementary [LaC16]. eleventh [Wel89]. Elimination [Bab17b]. Elliptic [Bar00].

Eloge [WWCP92]. emerged [Par96]. Emergence [Bul08]. Emerson [Ano91c]. Emily [Boo62, Fra63a, Fra63b, W.62, Wil64]. Emmerson [Kem79]. emploi [Bab31b, Bab34g]. Employed [Bab22e]. Enchantress [Swa01c]. Encyclopedia [BCKS^{+}83, Cle09, RR93]. End [CW91]. Endless [Ano00]. Enemy [Bab55b, Bab55c]. Energy [Cle09, Ano10a, CH02]. Engine [Ano08, Atk00, Bab37c, BCKS^{+}83, Bro00, BRG^{+}87, C.88, Don92, For75, Her57, Her99, Nic43b, Rob87a, Rob87b, Spe99, Sus94, Swa96b, Swa05, Tur91, Wil00c, Bab34b, Bab10, Bab89a, Bab89c, Bab89g, CK89l, Gre01, Her89, Lar89, MÈ78, MB89, Swa91d, Tee94, Ald91, Ano30, Ano77, Ano91e, Ano10b, Bab43a, Bab49, Bab56c, Bab56g, Bab89a, Bab82b, Bab82a, Ber64, Ber81, Bro82, BR93, Bro98, CK89k, Coh88, Cro85, DH90, DH94, DH98, DH04, Don92, DR92a, DR92b, Fra81, GS91, GS11, GC10, Gra71, Har49b, Her31a, Kil10, LL82, Lov43, Mar22, Men21a, Men21b, Pad17, Par96, Ran82a, Roe09, Seg03, Sus11, Swa00b, Swa01b, Swa02, SB02, Syd09, Tay92, TW87, Dis91, Pay66]. Engineer [Ano00, Arc47, WWCP92, Wil92b]. Engineering [Bar96, RR93, Hal09, Joh92, OLCJ91, OLCJ91, WSC^{+}88]. Engineers [Bab56c]. Engines [Ano91c, Ano21c, Ash02, AFB^{+}89, Bab10c, Boo62, Bro87, Bül88, CW91, Edi92, Fra63a, Fra63b, Lar34, Nic43a, Swa04, Tom02, Twe91, Ano91b, Bab89b, Bab82c, Bar72, Bro85, CK01a, Col71, Col90, CM98, Lin87, Lin90, MM61a, Swa91c, WB76, Bro90, O’N91, Wil91b, Wil76, W.62, Wil64, Pan63]. England [GG92a, Bab30i, Bab51a, BM70, BR89, Bab98, Bab06, CK89q, CK89r, Her31b]. English [BdL80, DR92a, DR92b, Far64]. Engraving [Bab38c]. ENIAC [Ano00, VBCC83, BEG^{+}87]. Enigma [BEG^{+}87]. Enlightenment [Bab08]. entirely [Byr38]. Entitled [Bab32b, Bab34a, Bab35b, Bab56e, She54, She60]. entre [Kan94]. Entwicklungsgeschichte [Mar25, Mar86]. Entwurf [MB89]. equation [Ben16, Cha33, Jar03, Mac07, NZ06, Rit16]. equations [BK11, BHK12, HB13, Mac07, Ort07a]. Equipment [All91]. Ernst [Twe93]. Erratum [DR92b]. Error [Mar22, SBET85]. Error-Correcting [SBET85]. Error-riddled [Mar22]. Errors [Bab27b, BDH90, MW80, MW72]. Erwin [OLCJ91]. Esq [Bab69, BH88, Enr88]. Esq. [Bab35a, Fin52]. Esqre [BBKC32]. Essay [Bab15, Bab16a, Bab30a, Gri06, Joh95]. Essays [JL11]. estimate [Mer82]. Etas [For75]. etc [Bab89b]. Euler [Bab17a]. eulogy [CK00, Dod74, Dod00]. Europe [SCD03]. European [BWV^{+}88]. event [Ano90]. ever [Ano11]. Everything [Mad86]. Evidence [Bab27a]. Evolution [Bro87, McC97, Bro85, Jen00, Sin99]. evolutionary [HR03]. Examination [Bab21, Bab53]. Examine [Sto55, Wro56]. Examples [BHP20, PHB20]. Excellence [BCKS^{+}83]. Exceptions [Bab48, Fin52]. excerpts [Rat52]. exhibited [Ano26]. Exhibition [BRG^{+}87, Fah91, OLCJ91]. Expectations [TW87, Wil91b]. expense [Mer82]. Experiments [BH25, BH30, Dol98, Ano26]. expert [Pil95]. Explaining [Atk00]. Exploring [Ros94b, Zen13]. Exposing [Bab55b, Bab55c]. Exposition [She54, She60, Bab51a, CK89d, CK89i]. Expressing [Bab26b, Bab30d]. extended [Dij86]. Extension [Bab65]. Extensions [FZZ08]. Extent [Bab34c, Bab47a, Bab34a]. Extinct [Bab60a]. Extract [Bab69]. Extracts [BL53]. Extraordinary [Gib19, Sus11]. eye [Ash94].

F [Ben16, Car69, Twe92]. F. [SBET85].

F.R.S. [Ano33, Bab69, BH88, Hor09, Enr88]. Faber [Bar96]. Fabrikenwesen [Bab33d]. Factory [Fre03]. Failure [Ano91c, AFB^{+}89, Bül88, CW91, Twe91, Lin87, Lin90]. Fall [HB22]. Faraday [GG92a, Sch02]. Farr [Mer82, Bab60b]. Faster [Bow53b]. father [Ano74, Coh88, Hal70]. Feb [Bab35a]. Federico [Lov43]. Fellows [Bab56b]. Fellowship [OLCJ91]. female [Ess14b]. Fiction [Kil10]. Fifty [Cha12, OLCJ91]. films [EEM^{+}05]. Finally [Syd09, Mat90]. finite [BHK12]. firm [Nim08, Rei07]. First [Ano00, Ano08, Bar00, Cro04a, Cro04b, DeF22, KT99, LL82, Roe09, Swa00a, Swa01b, WWCP92, Ano91e, BRG^{+}87, CK01b, HMB22, Mar22, Mer77, Pat60, She06, Swa00b, Swa01a, Swa02, SB02, Wil00a, Wil92b, AFI72]. flashing [Jon91]. Flight [BRG^{+}87]. flood [Gle11]. force [Fis10, Schxx]. Forgotten [Aus82]. Forlag [Bro86]. form [Wil91c]. Formation [Bab56d]. Forming [JC90, Joh92, Ano91a]. formulae [Byr38]. forthcoming [Ano90]. FORTRAN [BDH90]. Fortunato [Bul68]. Found [Bar00]. Foundation [Nor01]. foundations [Sta11]. founded [GG05]. Founder [Bar00]. four [Sny11]. Fourier [BHW^{+}90]. Fra [Ano93]. Fragment [Bab37a, Bab38a, Car69, CK89t]. Fragments [Roe09]. France [CW91, WSC^{+}88]. Franchise [Bab65]. Francis [Bar96]. Francisco [Twe92, Twe93]. Franksen [BCKS^{+}83, Bro86, SBET85, Vam85, Wil86]. Frederick [Sch81]. French [Bab31b, Bab33c, Bab34d, Bab34f, Bab34g, Bab35e, Bab55f, BdL80, Bul71, DR92a, Kan94, Mos83, Ort07a, Ort07b, Poi82]. Frequency [Bab57]. frequent [Bab31b, Bab31b]. Friend [BDH90]. friends [Sny11]. Frontier [Ano00]. fruits [Kou14]. Fuller [Mer82]. Functional [BK11, BHK12, Ben16, Cha33, Mac07, Ort07a, Rit16]. Functions [Bab15, Bab16a, Bab17c, Bab17d, Bab22e, Bab30a, Bab30c, HB13]. Fungus [CW91]. Funktionalgleichung [Cha33]. Futrelle [BCKS^{+}83]. Future [Lew21, Ber64, Ber81].

G [Ano91c, Arc47, For75]. G. [BCKS^{+}83]. Gadfly [Mul48]. Gallery [WSC^{+}88]. Game [Bab17a]. Games [Bab21, Bab53]. Gardiner [Byr38]. Geeks [Isa14]. General [Bab16b, Bab26e, Bab29d, Bab49, Bro85]. Genese [CW91]. Genesis [Nof15, Bow15, Das14, Mou15]. Genius [Cal68, Can66, Ogb65, Opp71, OS65, Phi65, Ess14b, Mos64a, Mos64b, Mos65]. Geniuses [Isa14]. Gentlemen [MT81, Sus11]. geography [SCD03]. Geological [Bab34a, Bab34c, Bab47a]. Geology [Dol98]. Georg [AFB^{+}89, Bül88, Twe91, Lin87, Lin90, Mer77]. George [Kem79, O’B93, GG05, Por07, Swa03]. Gerat [Bab49]. German [Bab33d, Cha33, Dot96, Dot97, Hym87, Mar25, Mar86, MB89, Schxx]. Germany [Bla02]. Gesetz [Schxx]. gets [Mat90]. Ghost [B.88]. Giant [Aus82]. Gibson [Dis91, Spe99, Sus11]. gift [Coh88]. given [Ano91c, HB13]. Glaisher [Mer82, And99a, And99b]. Glass [Bab57]. Glen [Bab68]. Glory [AFB^{+}89, CW91, Lin87, Lin90, Twe91, Ano91c, Bül88]. Goldberg [CSW90]. Golem [Ack95]. Gorn [WWCP92]. Government [WSC^{+}88, Bab51a, CW91]. Great [Bab34a, Hal09, Bab34c, Bab47a, O’N91, Bab29a, BB36, Rei68, Wil92a]. Greenwich [She54, She60]. Group [Isa14, BK11]. groups [BHK12, SBET85]. growth [Mou15]. Guard [THWV88]. Guide [AB86, Mat91, Yos02, Ano91d, Bab97, CBSW17]. guidelines [DS21]. gunnery [Joh95]. Guns [Bab55b, Bab55c].

H [Ano91c, BCKS^{+}83, Enr88, SBET85]. H. [BCKS^{+}83]. Hackers [Isa14]. Hairdressers [GG90]. Half [AFB^{+}89, Her88]. Half-Century [AFB^{+}89, Her88]. Hall [CW91, WSC^{+}88]. Hamden [Mad86]. Hand [Ess04, Par96]. Hand-Loom [Ess04]. Hans [Ano91d]. Happenings [MN88, OLCJ91]. Happy [VW91]. hardback [CK01b, Cro04b]. Harvard [Ano48, Ano82b, Coh88]. Hashing [GT80]. Heights [Bab24a]. Henry [Bab56c, O’B93, Coh88]. Heppenheimer [CW91]. Heritage [Tee83]. Herken [AFB^{+}89]. Herman [Aus82]. Herschel [Ano26, Ash94, Ash96, But74, Sch81, Tur91, Wil90]. Hexadecimal [BRG^{+}87, BEG^{+}87]. High [CW91]. High-tech [CW91]. Himself [Syd09]. Hinsley [SBET85]. Hist. [Ber92b]. histoire [Poi82]. Historic [JL11]. Historical [Cor87, HP16, All91, WB76]. Histories [Tom02]. History [Ano91c, Ano91d, Ano96, Ano08, AB86, Asp07, Aug84, Bab89a, Bab10c, Bow53a, BRG^{+}87, BHW^{+}90, CK01c, CKCFR03, CSW90, Cle09, Gul80, Gul81, Hun96, Kil10, MKW92, MW80, Mis07, Sev13, Sus94, Sus11, Swa96a, Twe92, Twe93, WSC^{+}88, Wil97, Ano93, Ano18d, Bab97, CK92, Cha12, Fis17, Fri92, Gle11, HHW08, Mar25, Mar86, MW72, Nor01, OLCJ91, O’R12, Pin06, Poi82, Ros94b, Wel89, Wil85, All91, WSC^{+}88]. hobby [Fra86]. Hodgkin [Ano00]. Holes [Ano91a, JC90]. Hollar [Sev13]. Hollerith [Aus82]. Holocaust [Bla02]. Home [Ano80, THWV88]. Hon [Bab29c]. Hoofs [Bab37b, Bab38b]. Hord [BCKS^{+}83]. Horizon [Tom02]. Horse [Bab37b, Bab38b]. House [Bab27a, Gre06, Bab56a]. Hoxton [Bab35a]. Human [Bab33b, Bab60a, Rat52]. Humphry [Bab22a, Bab22b, Bab30j]. Hut [BCKS^{+}83]. Hutton [Byr38]. Hyman [CK89a, O’H82, Ric83, Ros90, VBCC83, CSW90, Enr88, Hil85, O’D85a, Smi83, Tro84, W.86]. Hymanm [Asp82]. HyperCard [DH90, DH94].

I. [BCKS^{+}83]. i.e [Bab89b]. IBM [Ano91c, Bla02, BEG^{+}87]. Idea [Wol16]. Idealistic [Bow93]. Ideas [CW91, Lew21, Wol16]. idéologie [Ort07a]. ideology [Ort07b]. IEEE [Ano16, Ano17, Ano18a, Ano18b, Ano18c, Ano21a, Ano22, WSC^{+}88]. IEEE-CS [Ano16, Ano17, Ano18a, Ano18b, Ano18c]. ihre [Mar25, Mar86]. II [Bab16a, CK89n, O’B93]. III [Bab43a]. Illiac [BCKS^{+}83]. illus [L.90, Twe93]. Illustrated [Aug84, Hun96, Mad86, O’H82]. Illustration [Bab38c]. Imaginary [Pur93]. Imagination [Tom02]. Immanuel [Vam85, Wil86]. Impact [Gig01]. implementations [Sta11]. Impossible [SHH81]. Impressions [Bab37b, Bab38b]. improved [Byr38]. inadvertent [Ozg10]. including [Jon91]. incomplete [Bab89a]. Inconsistent [Gri10]. industrial [Ash96, Mus06]. Industries [AFB^{+}89]. Industry [Bab51a, Biz03]. Infinite [Bab19b, Bab26e, Bab30h]. Influence [Bab27c, Nim08]. influences [GST01]. influential [Ano11, Cur10]. Information [Aus82, CW91, Cur12, Ess04, Gul80, Gul81, Hun96, WSC^{+}88, Yos21, CH02, GP08, Gle11, OLCJ91, Hun96]. informatique [CW91]. Innocence [CW91]. Innovation [Aga18, AFB^{+}89, De 06, Jon16, CW91]. Innovators [Isa14]. Inquiry [Gib19]. Inspiration [Ric97]. inspired [Mar22]. Institute [CK92, Gul80, Gul81, MN88, Sei95, Twe92, Twe93, WSC^{+}88, Yos21, All91, Ano80, AB86, Asp07, CK01c, Mis07, Nor01]. Institution [WSC^{+}88, All91, Bab56c]. Institutions [Bab26a, Bab67b, Bab89b, CK89o, CK89p, Coh88]. Instruments [Har49a, Har12]. Integral [BHP20, LaC16, PHB20]. Intelligence [SBET85]. Interdisciplinary [Yos21]. Interface [Wil00c, Bro00]. international [Bel84, Bab60b, OLCJ91]. Internet [Ano00, DH98, DH04]. Interpolation [Com36]. introducing [NJ05, Pil95]. Introduction [Sta11, Wil00c, WSC^{+}88, DH90, DH94, DH98, DH04, Twe92]. introductory [Bab29b]. Invent [Bab99]. Invented [Men21b, Sla87, Jon91, Smi10]. Inventing [Ano00]. Invention [Nic43a, Ano06, Gre06, Hal09, Mou15, SU96, Ugl96]. inventor [BV07, Mos64a, Mos65, Cal68, Can66, Ogb65]. Inventors [Smi71, Cur10]. Inverse [Com36]. Investigating [Bab19b, Bab30h]. investigation [GG05]. Invitation [SGM07]. Involvement [WSC^{+}88]. involving [Mac07]. Iowa [L.90]. Irascible [Mos64a, Mos64b, Mos65, Phi65, Cal68, Can66, Ogb65, Opp71, OS65]. Irish [GP08]. iron [Ano26]. ISBN [Ano91c, Bar96, Bro86, CK01b, Cro04b, Seg08, Twe89b, Twe91, Twe92, Twe93]. Issac [O’B93]. Issue [BDH90]. issued [OLCJ91]. Italian [For75, Bab34e, BV07, Bul71]. italiens [Bul71]. IV [BCKS^{+}83].

J

[Bab35a, Ben16, Bey79a, Bey79b, BCKS^{+}83, CW91, Enr78, Hil79, Kem79, Mer82]. J. [Bey79c, BCKS^{+}83, HG79, Rob79]. Jacquard [Ess04]. Jahr [MB89]. James [Ash02, O’B93]. Jan [Ano91a]. January [Ano48]. Japan [AFI72, AFI72]. Java [SGM07]. Jean [O’B93]. Jean-Baptiste [O’B93]. Jennings [CW91]. Jenny [Bar96]. Jeremy [O’B93, Pay66]. Joan [Mad86]. Johann [Ano91c, AFB^{+}89, Bül88, CW91, Twe91, Lin87, Lin90]. John [Ano91c, Enr80a, Enr80b, O’B93, Sch81, Tur91, Ano91c, Ash96, But74, Sev13, Smi10, Twe92, Kem79]. Johnson [Ano91a, Ano91c]. Jointly [Ano48]. Jones [Aga18, O’B93]. Joplin [O’B93]. Joseph [Ano91d, BHW^{+}90, WWCP92, Wil92b]. Journal [Her31b, HBC^{+}55, Ano91a]. July [Ano90, GG92a]. June [She54, She60].

Karl [O’B93, Wen09]. Katz [Bar00]. Kelly [Hun96, Twe89a, Twe89b, Sib95]. Kennedy [Mer82]. Kernkompetenzen [Schxx]. Kevin [Sus11]. Key [Bar00]. Kidwell [Ano00, Twe93]. Kings [Lee95]. Knight [Bab17a]. Kompas [For75]. Kraftvereinigung [Schxx]. Kuhn [CW91]. Kyrupaedi [Schxx]. Kyrupaedie [Schxx].

L [Aga18, HG79, Mer82]. lab [Fis10]. labor [Poi82, Rei07]. Laboratory [Ano48]. Labour [Bab33b]. Labours [Enr88, BH88, WB76]. Lady [SHH81, HH80, Ste84]. landmark [KT99]. language [Ort07a]. Languages [Ort07b, Bab31b, Fri92]. langues [Bab31b]. Lardner [Bla91, O’B93]. Large [Ano48]. Large-Scale [Ano48]. Last [Bab56b, Wil98]. Late [Enr88, BH88]. Lauderdale [O’B93]. launched [Ess14a]. Law [Schxx, THWV88]. Laws [Bab51b, GG05]. Laying [Bab55b, Bab55c]. Leadership [Asp07]. League [Sus11]. Learned [Wro56]. learning [SCD03]. Lectures [Mac19]. Led [Ess04]. Lee [WWCP92]. legacy [HR16, AFB^{+}89]. Lego [Col06, Fis10]. Lehre [Schxx]. Leno [Ack95]. Leslie [O’B93]. less [Bab31b]. Lessons [AFB^{+}89]. Let [GC10]. Letter [Bab22a, Bab22b, Bab29c, Bab32a, Bab33a, Bab35a, Bab39, Bab60b, Bab69, Ber92a, Van83, Bab35e, Ber92b, Fin52, Her31b, She54, She60]. letters [Bab31b]. lettre [Bab35e]. lettres [Bab31b]. Library [Ano90, Bab97, Wil80, Wil81b, Bro91]. Licklider [CW91]. Life [Ano00, AFB^{+}89, Bab24b, Bab56f, Bab64b, Bab69, Bab89e, Bab94, BV07, Bab11, BHW^{+}90, Cal68, Can66, CW91, Enr88, Far64, MM52, Opp71, Wil95, Ano82a, Bab68, Bab91a, BH88, CK89j, Goo02, Jon91, Kou14, Mos64a, Mos64b, Mos65, Ogb65, Sch81, WB76]. Lifetimes [Far64]. Lighthouses [Bab51c, Bab53, Jon91]. Lights [Bab51c].

Limehouse [Ack95]. Limits [WSC^{+}88]. Lindgren [Ano91c, Twe91, AFB^{+}89, Bül88, CW91]. Line [Tom02]. Linear [BHW^{+}90, NZ06]. list [Bro91]. Listing [Bab97, Ano91d]. Lists [Schxx]. Literature [Mat91, CH02]. Little [CK01b, Cro04b, Col71, Col90]. Lives [Bab26a, Bab67b, Bab89b, CK89o, CK89p, Fis14, Wol16, Sib95, CK94]. Local [Bab23a]. Localities [Bab60a]. Locating [Ber92a, Ber92b]. Logarithm [Ano03, Bab31a]. Logarithmic [GG90]. Logarithms [Bab27b, Bab27d, Bab31c, Bab15, CK88, Bab29e, Byr38, CK89b, Mou15]. logic [GG05]. London [Ano90, CK01b, Car69, Cro04b, Enr80a, Enr80b, Twe89b, Twe91, Twe92, Twe93, Wil64, Bab97, Ric97]. Longfield [O’B93]. Look [DeF22]. Loom [Ess04]. Lord [Sus11, Ess14a, Ess14b]. lords [Ano06, Gre06]. Losano [For75]. Lovelace [Sus11, Ess14a, Ess14b, FFT^{+}03, FF03b, FF03a, FF15, Hoo12, HH80, HG79, KT99, KM12, Lov12, Men21a, Men21b, Pad15, Pad17, Ste84, SHH81, Swa11a, Swa13, Swa17, BL53]. Ltd [Twe89b]. Lucas [FZZ08]. Lucasian [BBKC32, Gri78]. Luddites [Spe99]. Ludgate [Ran82a]. Luigi [Lov43]. Lukoff [BCKS^{+}83]. Lyell [Dol98]. Lyle [Ano91c].

M

[BBKC32, BR89, Bey79a, Bey79b, Bey79c, BCKS^{+}83, CK89q, CK89r, Enr78, Hil79, HG79, Kem79, Rob79, SBET85, Wro56]. M. [BH25, BH30, Bab35e, Bab43a, Men43, Sto55]. mêmes [Bab31b]. Maboth [Cal68, Can66, Ogb65, Opp71]. macchina [For75]. macchine [Bab34e, Los72, Tri72]. Machine [Air56, AFB^{+}89, B.88, Men43, Pur93, Sto55, Sus00a, Sus00b, Swa91b, THWV88, Wat12, Wro56, Aga01, Ano91e, Ano10a, Bab35e, Bab55f, DR92a, Her88, Jen85, PP79, BCKS^{+}83, BHB^{+}89, Com36, Mer82, Hun96]. Machinery [Ano91c, Bab22b, Bab22c, Bab22d, Bab23b, Bab26b, Bab29d, Bab30d, Bab33b, Bab35c, Bab38c, Bab99, Bab71, Bab03a, Bab10a, Ber82, CK89c, CW91, SS57, Sus00b, Ano32a, Ano32b, BBKC32, Bab32c, Bab32d, Bab33c, Bab33d, Bab34d, Bab34e, Bab34f, Bab46b, Bab63, Bab86, BB36, CK89s, Jer89, Lew03, Ano48]. Machines [Aga18, Bow53b, Com31, Har49a, Har12, Jon16, Mar25, Mar86, MKW92, Swa89, Twe93, Bab33c, Bab34f, Bab34g, BdL80, Bab10b, Bai89, Bro92, Cha12, Col71, Col90, Coo05, Dot96, HNW^{+}48, Jon91, OLCJ91]. Macht [Dot97]. MacLachlan [Ash02]. Made [HB22, Byr38, Lee95, Par96]. Magazine [BR89, CK89q, CK89r]. magnesium [Ano26]. Magnetic [Bab26d, Bab30e]. Magnetism [BH25, BH30]. magnetological [Pat60]. Majesty [Ano06, Gre06]. Major [Yos21, Bab56c]. Makers [AG02, Arc48, Wol16]. Making [Ber82, Kou14, Swa89, Aga01, Bab89c, CK89k, Fah91]. Malthus [O’B93]. Man [O’H82, Bro92, Smi10]. management [Ano11, Lew07]. Managing [CW91]. manifatture [Bab34e]. Manifested [BH25, BH30]. Manual [Com47]. Manufactories [Bab03b]. Manufacture [Bab29d]. Manufactures [Bab35c, Bab71, Bab03a, Bab10a, CK89c, Ano32a, Ano32b, Bab29b, BBKC32, Bab32c, Bab32d, Bab33c, Bab33d, Bab34d, Bab34e, Bab34f, Bab46b, BdL80, Bab63, Bab86, Bab10b, BB36, CK89s, Lew03, Sus00b]. Manufacturing [Bab33b, Biz03, BdL80]. manuscripts [Bab97]. many [Bab27b]. Marian [BEG^{+}87]. Mario [For75]. Mark [O’B93, Smi71]. Market [AFB^{+}89, Bar00]. Marshall [CM11, Nim08]. Martin [Hun96, Sib95, Twe89a, Twe93, Twe89b]. Marx [O’B93, Wen09]. Mary [Sin99].

Maschinen [Bab33d]. Maschinen- [Bab33d]. Mass [Twe91, Twe92, Twe93]. Materials [Ano91a, JC90, Tee82]. Mathematical [Arc48, Bab22b, Bab22d, Bab27c, Bab30b, Bab37c, Bab82a, Bab89d, Bey79a, Bey79b, Bey79c, CK89h, CKCFR03, Enr78, GG79b, Gri78, Hil79, HG79, Kem79, Rob79, Bab55f, Bai89, Dub78, DS21, Enr80a, Enr80b, GG05, Jer89, Ort07b, Pad17, Sta11, GT80]. Mathematician [Hym87, Syd09]. Mathematicis [Arc35]. Mathematics [BBKC32, FRW11, GG79a, GG79b, BS05, Enr83, Plo12, Ric97]. Mathematiker [Hym87]. mathématiques [Bab55f]. Matter [Aga18, Jon16]. Matters [Kil10]. Matthew [Bab16b, Aga18]. Matyas [BCKS^{+}83]. Mauchly [Smi71]. May [Bab34a, Bab69, Ber92b, Bab34c, Bab38c, Bab47a, Van83]. McCulloch [O’B93]. McKay [Ano91c, Twe91]. Meaning [Ano00]. Means [Bab17d, Mer82]. Measurement [Bab24a]. mecánica [Bab35d]. mécanique [Mos83]. Mechanical [Bab29d, Bab51b, Bab56c, Bab56g, Swa89, Swa93, Bab89a, CK89l, HHW08, Mos83, Rat52]. Mechanician [Arc47]. mechanics [Bab35d]. mechanism [Bab22f]. mechanization [BS05]. Medal [Bab47c]. mediated [NJ05]. Meet [Col06]. meeting [She54, She60, Ano82b, BHB^{+}89]. Meets [Sus94]. member [BBKC32]. Members [Bab39]. Memoir [BH88, WB76, Bab43a, Enr88]. memoirs [Bab43a]. memorial [Kno15]. Memoriam [CW91]. Memory [Ash96]. Men [Bab55b, Bab55c, GP08]. Menabrea [Bab43a]. Menebrae [Lov43]. mental [Poi82]. merchants [GP08]. merely [Hoo12]. Merrifield [Mer82]. Messrs. [Ano26]. Metal [Bab47b]. metals [Ano26]. Method [Bab17a, Bab26b, Bab30d, Bab55b, Bab55c, Bab03b, Swa10, Bab38c, Ort07a]. Methods [Bab19b, Bab30h, Eck40, Eck84].

Métiers [All91]. Meyer [BCKS^{+}83]. Michael [Ano91c, Bül88, Enr80a, Enr80b, Sus11, Twe93, Twe91]. Michigan [BHB^{+}89]. microfilm [Ano91d]. Micrometer [Bab26c]. Microprogramming [BHW^{+}90]. Microsystems [CW91]. middle [Pil95]. Milan [For75]. Mill [O’B93]. mind [Ash96, HHW08]. Minds [Coo05, HM02]. Mini [Ano10a, Tay92]. mini-computer [Tay92]. miscellaneous [Bab89f, CK89m, CK89n]. Mixed [Bab60a]. MM. [Bab55f]. Modern [Bec80, CW91, Com31, Gib19, Aga01, Ano91c, Don92, Dub77, Joh91]. moderne [Tri72]. moins [Bab31b]. Moll [Rei68]. Monuments [Arc48]. Moorcock [Sus11]. Moore [HG79, Sus11, Joh95]. Morgan [Ric97]. Morrison [Boo62, Fra63a, Fra63b, Pan63, W.62, Wil64, Boo62, Fra63a, Fra63b, W.62]. mortal [Hoo12]. Moseley [Cal68, Can66, Ogb65, Opp71]. Most [EB19, Ano11, Bla02, Cur10, Lea20]. motion [Joh95]. Mounier [CW91]. Mounier-Kuhn [CW91]. Mountifort [O’B93]. Move [Bab17a]. Mr [BCKS^{+}83, Gre06, Mer82, SBET85, Wil95]. Mr. [Ano06, Bab35e, Bab35a, Bab39, Bab49, Bab54a, Bab54b, Bab54c, Bab54d, Bai89, Col06, Fra81, Fra84, Fra85, Fra86, Her31a, Hor09, Mer82, Nic43a, Nic43b, She54, She60, VW91, Wil91c, Bro86, Vam85, Wil86, Wor87]. MR0714019 [Ber92b]. MR1503574 [Ben16]. MTAC [GT80]. Muller [AFB^{+}89, Bül88, Ano91c, CW91, Twe91, Lin87, Lin90]. Munich [WSC^{+}88]. Murphy [THWV88]. Musée [All91]. Museum [All91, Bro91, Fah91, O’N91, Sus11, WSC^{+}88, OLCJ91, Ano08, OLCJ91]. Museums [All91, WSC^{+}88]. My [CW91, Hoo12]. Mysterious [Ran83]. mystery [Fra93].

N. [VBCC83]. Napier [Twe92, Kno15, Nag80a, Nag80b]. Naples [Bab34c, Bab46a, Bab47a]. Nash [Ano91c]. Nassau [O’B93]. National [De 06, WSC^{+}88, All91, Kou14, All91, Com36]. Natural [Bab27d, Bab31c, Bab15, Bab29e, Byr38, GST01]. Nature [Bab32b, Bab32a, Bab33a, Bab35b, Bab56e, Byr38, MM11, Zen13]. naval [Joh95]. Navigation [Bab55d]. Navy [Ano48]. Nazi [Bla02]. NCC [Ano82b]. Near [Bab33b, Bab46a, Bab47a, Bab34c, Rei07]. near-decomposability [Rei07]. needlessly [Thi00]. Neptune [Bab47c]. Nervous [Mus06]. Neumann [Ano91c, CW91, Gol72, Gol93]. Neuromancer [Sus11]. never [Ano11]. News [Ano82b, Bar00]. newspaper [Bab51d]. Nineteenth [Car69, Cla03, Fre03, Mac19, Swa89]. nineteenth-century [Fre03]. Ninth [Bab37a, Bab38a, CK89t, Car69, Bab09]. No [Ano08, Bab49, BRG^{+}87, Don92, Rob87a, Rob87b, Swa96b, Swa05, Tur91, Tee94, Tay92]. no. [Ber92b]. Nominations [Ano21a]. Norberg [Asp07, CW91, Mis07]. Norrington [Sus11]. North [Ano08]. Norwegian [Ano93]. notable [Wol16]. Notation [Bab22e, Bab30f, Bab51b, Bab56c, Bab56g, BCKS^{+}83, BRG^{+}87, BEG^{+}87, Fra81, Bab89a, Byr38, CK89l]. notations [DS21]. Note [Ano61, Bab17b, Bab22c, Bab51c, Bab52, Ben16, Rei68]. Notebooks [Wil81a]. ’Notes [FF15, Arc48, KM12, Lov43, Men21b, Ste84, Coh88, Men21a, BHW^{+}90, FF03a]. Nothing [Mor95]. Notice [Bab27b]. Notices [Ano82b]. Notions [Men43]. Novelty [Dol98]. nuisances [Bab64a]. Number [Bab29c, Swa91a, Swa01c, Bab22f, Mat90]. Numbers [Bab19a, Bab27d, Bab31c, Bab32b, Bab35b, Bab56e, Bab15, Bab29e, Byr38].

O [Bro86]. O. [BCKS^{+}83, SBET85].

Obituary [Ano71c, Ano71d]. Object [Par96]. Observations [Bab17c, HB22, Bab22d, Bab22e, Bab24a, Bab30c, Bab34c, Bab46a, Bab47a, Bab56b, Bab56c, Bab56f, Bab60a, Bab68, Her31b, Bab34a]. Observatory [Ano90, She54, She60, THWV88]. Observed [Bab52]. observers [Jen85]. Observing [Bab03b]. Occasion [JL11, Sen71]. Occulting [Bab51c, Bab55e]. Occurrence [Bab57]. Ocean [Bab56d]. October [AFI72]. Oh [Wei88]. Old [Ano91e]. Ole [Vam85, Wil86]. OM [Lew03]. One [Gib19]. O’Neill [Sus11]. ongoing [Kou14]. opens [OLCJ91]. Operation [Com47]. operations [Lew07]. operators [NZ06]. Ophthalmoscope [Lyo40, Ric02]. Opthalmoscope [Bab54e, Jon54]. Optici [BM23]. Optics [BM23]. Oral [BRG^{+}87, AB86]. Ordnance [Ano48]. Origin [Ano61, Bab60b, BCKS^{+}83, Fra81, Byr38]. Original [Bab35a]. Origins [Ano91c, CW91, Ran73, Ran75, Ran79, Ran82b, Swa13, PP79]. Other [Bab17c, Kil10, Ano26, Bab30c, Coh88, MM61b, Tuk72]. Our [Ano18d]. outils [Bab34g]. ouvrage [BdL80]. Overcoming [THWV88]. Oxford [Sus11].

P [Bab29c, Pan63, Bab43a, For75]. P. [Arc47]. Packings [SBET85]. Padua [Sus11]. pages [Wor87]. Paintings [Arc48]. Palmer [Ano91c]. Pandæmonium [Jen85]. Paper [Bab34a, Ben16, Lov43]. Papers [Ano91d, Bab31c, Bab10c, Lew21, Ran73, Ran75, Ran82b, Van80, Bab89b, Bab82c, Bab89d, Bab89f, Bro91, CK89h, CK89m, CK89n, Gre06, Bab97, BL53]. Parallel [Bab68]. Parliamentary [Gre06]. Part [Ste95, Bab16a, Bab43a, Bab97, Mar25, Mar86]. Partnership [EB19, Lea20]. Parts [Sha10]. Pascal [Aga18, Gol72, Gol93, Jon16]. Passages [Bab64b, Bab68, Bab69, Bab89e, Bab91a, Bab94, BV07, Bab11, CK89j, Gro22, Gro23, Wil95]. Passaggi [BV07, Seg08]. Passion [Mad86]. Passionate [EB19, Lea20]. past [Ber64, Ber81, Mat90]. Peacock [Bec80, Dub77, Wil90]. Peerage [Bab56f]. Peggy [Twe93, Ano00]. Pentium [Ano93]. Penzias [CW91]. people [Wol16, OLCJ91]. perfection [CK01a, CM98, Ash02]. Permanent [Ano80]. Personal [Atk00, CSW90, Wol16]. Perspective [BEG^{+}87, CM11, Nor01]. Perspectives [Wil92a, Wol16]. Peter [BR89, CK89q, CK89r]. Peters [BCKS^{+}83]. pharaohs [Pin06]. Phil [Bar00]. Philip [Boo62, Fra63a, W.62, Wil64, Fra63b]. Philosoph [Hym87]. Philosopher [Bab64b, Bab69, Bab89e, Bab94, BV07, Bab11, Hym87, Bab68, Bab91a, CK89j, Rat52, Wil95]. Philosophers [Bab29a, GP08]. Philosophical [Bab30j, Sny11, BR89, CK89q, CK89r]. Philosophy [Hal09, Fis17, GST01]. physicists [Mac19]. Pickering [Twe89b]. Picturing [Pad17]. Piece [Swa96a, OLCJ91]. Pink [Bab52]. Pioneer [Gri06, Hil85, Hym87, O’D85a, O’H82, Ric83, Smi83, Swa11a, Tro84, W.86, Wil91a, Hym82a, Hym82b, Jon91, KT99, Lew07, Ros94a, Smi10, Wil77, Ano82b, BHW^{+}90, O’D85b]. Pioneers [Mat91, Wei88, O’B93, O’B93]. piracy [Col98]. PKWare [Bar00]. Plan [Bab49]. Planet [Bab47c]. Planing [Bab47b]. Plans [Bro00, Wil00c]. Plate [Bab57]. Plate-Glass [Bab57]. plus [Bab31b]. pocket [Tee95]. Pole [Mer82]. politic [Mus06]. politica [Bab35d]. Political [Ber82, Bab35d]. popular [Ano26]. Porisms [Bab23a, Bab30g]. portion [Coh88]. portrait [FF03b]. Portraits [Arc48]. possibility [Gre01]. Possible [Bab55e]. postmodern [Cla03]. pour [Bab55f]. Power [Dot97]. powerful [Bla02]. Powers [Bab37c, Bab82a]. Pozzuoli [Bab34a, Bab34c, Bab46a, Bab47a]. Pp [CK01b, Cro04b, Twe91, Twe92, Twe93, Ano91a, Ano91c, Bro86, Dis91, Enr80a, Enr80b, Hun96, L.90, Mad86, O’H82, O’N91, Seg08]. pp. [CK89a]. práctica [Bab35d]. practical [Bab35d, GP08]. Prandi [Bul68]. Pray [Wil91c]. Pre [Ste95, Swa11b]. Pre-Computer [Ste95]. Pre-electronic [Swa11b]. preceding [Ben16]. Precursor [Enr83]. precursore [Tri72]. predict [Col98]. Preface [BH13]. prehistory [Cer83]. Premiums [Far64]. prepared [Bab56c]. present [Ber64, Ber81]. president [Ano06, Gre06, Bab56b]. Press [Ano91c, CK89a, Enr80a, Enr80b, L.90, Mad86, O’H82, Twe91, Twe92, Twe93, Asp82]. Press/Los [Twe92]. Pressure [Bar72]. Prevost [Bab56c]. price [Ano91c]. Prime [Bab19a]. Princeton [Asp82, O’H82]. principle [Schxx, HR03, Rei07]. Principles [Bab23b, Bab29d, Bab47b, Bab48, Bab29b, Fin52, MM61b]. Printed [CW91, SS57, Van80]. Printer [Bar00]. Printing [Bab22b, Bai89, Jer89, Mer82, Mer77]. Prinzip [Schxx]. Private [Hym92]. probabilities [GG05]. Problem [Bab17a, BCKS^{+}83, Fra81, CM11]. Problems [Bab17d]. Proceedings [AFI72, Ano48, Wro56]. Processing [Ano91a, Aus82, Cor87, Gul81, WSC^{+}88, DB05, OLCJ91, Gul80, WSC^{+}88]. Produce [Bab34a, Bab34c, Bab47a]. Production [BRG^{+}87, GG90, Rob87a, Rob87b, Fre03]. Products [Bab13]. Professor [BBKC32, Mer82, Ano32a, Ano32b, Gri78, Mer82]. Profile [Swa11a]. Progenitor [Schxx]. Program [CW91, SHH81, All91, Cer83, KT99, OLCJ91]. program-controlled [OLCJ91]. programmazione [Los72]. Programmer [Wil00c, Bro00, HMB22]. Programming [Fra81, Ran91, Bel84, Fri92, PP79, BCKS^{+}83, BHW^{+}90]. Programs [Roj21]. progress [OLCJ91]. Projections [Bab52]. promises [Ano10a]. Promotion [Bab39]. Prony [GG90, Poi82]. Properties [Bab16b]. Property [Bab48, Fin52]. Proportionate [Bab29c]. Protector [Bab24b]. Protocybernetic [Tom02]. Prototype [Roe09]. Proves [Syd09]. PRS [Bab22b]. psychology [GST01]. publication [KT99]. Publications [All91, Fra63b]. Publicist [Arc47]. published [Ano91c]. Publishers [Twe89b, Twe92, Twe93]. Pugh [Ano91c]. Punched [CW91, Eck40, Eck84]. Purpose [Bab22b].

quantum [Pin06, Sin99]. Queen

[Sin99, Dis91]. Queries

[Ano81, Ber92a, BRG^{+}87, BEG^{+}87, BWV^{+}88, BHB^{+}89, BHW^{+}90, BDH90, Hym92, Rob87a, Rob87b, Rob93, Wil00b]. Query [Ano81, BRG^{+}87, BEG^{+}87]. Quest [Ano00, Cro04a, Cro04b, Swa00a, Swa01b, CK01b, Swa00b, Swa01a, Swa02, SB02, Wil00a]. Question [Ber82]. Questions [Bab21, Bab30b, Bab53]. Quetelet [Ber92b, Bab35e, Ber92a, Ber92b, Van83]. qui [Bab34g].

R

[Ano91c, BCKS^{+}83, CW91, Twe92, Twe93]. R. [BCKS^{+}83]. Rabdology [Twe92]. Races [Bab60a]. Racial [Kil10]. radicalism [Por07]. Radio [BCKS^{+}83]. Rae [O’B93]. rally [Mul02]. Ramsay [O’B93]. Randell [JL11]. Random [Ano10b]. Rates [Bab24b]. read [Ano11]. reader [Ada10, Spi02]. Reading [Wil81a]. ready [Ano91e]. real [Ben16, Rit16]. Realities [WSC^{+}88]. reappraising [Lew03]. Reason [Bul08, Woo99]. Reasoning [Bab27c]. recently [Byr38]. receptions [Coo05]. Rechen [Dot96]. Rechen-Automate [Dot96]. Rechenmaschinen [Mar25, Mar86, Twe93, MKW92].

Reckoners [Cer83]. Reckoning [Jon16, Aga18]. Reclaiming [Lew07]. Recognisable [Bab60c]. records [All91, WSC^{+}88]. Recursiveness [BCKS^{+}83, Fra81]. Redeeming [Swa93]. Rediscovered [LL82]. Reduced [Swa91a, Bab22f]. Reference [Bab48, Fin52]. referenced [Bro91]. Reflections [Bab30i, BM70, BR89, Bab06, Wil81a, HP16, Rat52, CK89q, CK89r]. Reform [Bow93, CK89a, CSW90, Ros90, BH89, Hym88, Hym02, Por07]. Reformer [Asp82]. Regulate [Bab29d]. Rejewski [BEG^{+}87]. Relating [Bab19a, BM23, Bab10c, WSC^{+}88, Bab89b]. Relative [Bab17a, Bab57, Ano06, Bab35e, Gre06]. relativement [Bab35e]. relays [Cer83]. Remains [Bab60a]. remarkable [Sny11]. Remarks [Bab34a, Bab34c, Bab47a]. rendered [Bab38c]. renewal [Enr83]. Repetition [BH25, BH30]. reply [Fin52, She54, She60]. Report [Ano91a, Bab54e, Dod00, Her31a, Her57, Her89, Jon54, Mer82, Sto55, Wro56, Bab91b, JC90]. representation [CH02]. Representing [Dol98]. Reprint [CK92, CK01c, Twe93]. Research [Ald91]. Resembling [Bab37b, Bab38b]. Residence [Ric39]. Resources [Yos02]. Respecting [Ano61, Bab17b, Bab22c, Bab52, Nic43b, Bab27b, Bab51c, Bab89g]. Response [BEG^{+}87, BHW^{+}90]. rest [Lee95]. result [Bab34g]. resulted [KT99]. résultent [Bab34g]. Rethinking [Spe99]. Review [Aga18, Ano91a, Ash02, Asp82, Bar96, Bey79a, Boo62, Bow15, Bro86, CK89a, CK01b, Can66, Car69, Cro04a, Cro04b, Dis91, Enr78, Enr80a, Enr80b, Enr88, Fah91, For75, Fra63a, GG79b, Gri78, Gri06, Hil79, Hil85, Hun96, Kem79, L.90, Lav14, Lea20, Mad86, O’B93, O’D85a, O’H82, O’N91, Ogb65, Opp71, OS65, Pan63, Pay66, Ric83, Rob79, Ros90, Seg08, Sib95, Smi83, Sus00a, Sus00b, Sus11, Swa91a, Twe89a, Twe89b, Twe91, Twe92, Twe93, Vam85, W.62, W.86, Wil95, Wil00a, Wil86, Wor87, Bül88, Cal68, Nof15, Tro84]. Reviews [Ano91c, Ano00, AFB^{+}89, Bey79b, Bey79c, BCKS^{+}83, CK01a, CSW90, CW91, HG79, O’D85b, SBET85, VBCC83, Wil64, Fog01]. revisited [CM11]. Revolution [Isa14]. Revolutionaries [Mat91]. Ricardo [O’B93]. Richard [O’B93]. Richardson [Twe92]. riddled [Mar22]. Rider [Twe92]. Right [Bab29c]. Rights [Bar00]. Ritt [Ben16]. Roads [Bab68]. Robert [O’B93]. Robots [Kil10]. rockets [Joh95]. Roget [BR89, CK89q, CK89r]. romance [Woo99]. Rose [CW91]. Rosen [CW91]. Rotation [BH25, BH30, Ano26]. Rotations [Bab26d, Bab30e]. Roy [Bab68]. Royal [Ano90, Bab56b, Gre06, She54, She60, Ano06, Bab47c, Her57, Wel89, Wro56]. run [Ano91e]. Russell [Kem79].

S [Ano91a, BCKS^{+}83, Kem79]. Samples [Ano10b]. Samuel [WWCP92]. San [Twe92, Twe93]. Sandstone [Bab37b, Bab38b]. Saul [WWCP92]. sausage [Lee95]. Savings [Bab27a]. Say [O’B93]. Scale [Ano48]. Scene [BEG^{+}87]. Scheutz [THWV88, Twe91, Wro56, Air56, Ano91c, Arc47, AFB^{+}89, Bab55f, Bab56c, Bab56g, CW91, LL82, Lin87, Lin90, Mer77, Sto55, Bül88]. scholastic [Bel84]. school [DR92a, DR92b, Pil95]. Schriften [Dot96]. Science [AG02, Ano30, Bab30i, Bab34d, Bab51a, BH89, Bab98, Bab06, BCKS^{+}83, BHW^{+}90, CSW90, Fah91, Gib19, Hym88, Kil10, Lew21, O’N91, RR93, Rei68, WSC^{+}88, Ano91d, Bab22f, BM70, BR89, Bab97, Bre75, CK89q, CK89r, CH02, Das14, DH90, DH94, DH98, DH04, Gre01, GST01, HMB22, Her31b, Hym02, Kou14, MT81, Mus06, OLCJ91, Plo12, Por07, SGM07, SCD03, Sny11, Woo99, Bab39, Bro91, CK89a, MT81, Sus11, Nof15, Bow15, Her31b, Ros90].

Sciences [Gig01, Seg03]. Scientific [Ano91c, Arc47, Bab89f, CK89m, CK89n, Com36, Mul48, Wil80, Wil81b, Yos02, Bab97, Eck40, Eck84, Fis17, HM02, Bab43a]. Scientist [Asp82, Rat52]. Scientists [DG99, Hal09]. scienziato [BV07, Seg08]. Scots [Sin99]. Scott [Kem79]. Scrapbook [Bab51d]. Scriptores [BM23]. Scrope [O’B93]. sculpture [OLCJ91]. Search [BCKS^{+}83]. Sebastapol [Bab55e]. Second [Car69, SBET85]. secrecy [Sin99]. Secret [Wil86, Bro86, Fra84, Fra85, Fra86, Vam85, Wor87, SBET85]. See [Fah91]. Seeking [Ano91e]. seen [Jen85]. Selected [Boo62, CK89a, CSW90, Fra63a, Fra63b, Ran73, Ran75, Ran82b, Ros90, BH89, Dot96, Hym88, MM61a]. selection [Bro66]. Selections [Ano32a, Ano32b]. Selects [Ano80]. Sells [Ste95]. Semiconductor [AFB^{+}89]. seminal [MM61b]. Senior [O’B93]. September [Bab29a]. Sequence [Com47]. Serapis [Bab34a, Bab34c, Bab46a, Bab47a]. Series [Bab19b, Bab26e, Bab30h, CK01c, Twe92, Twe93, Bab97, CK92]. sessional [Gre06]. Several [Bab19b, Bab30h, BBKC32, Byr38]. Sexes [Bab29c]. Shadow [But74]. Shadowy [BS05]. Shaped [Gib19]. Sharp [Bab35a]. Shocking [Swa04]. Shoe [Mad86]. Shooting [Byr04]. Short [Bab38c]. Shorthand [Bab35a]. Shot [Bab55b, Bab55c]. Showed [CW91]. Signs [Bab26b, Bab27c, Bab30d, Bab60c, Dot97, Ort07b]. Silent [Ait16]. Simonde [O’B93]. Simple [SBET85, Byr38]. Simulation [Ald91]. simulators [Don92]. Sinderen [Ber92b]. Sine [BRG^{+}87, Rob87a, Rob87b]. Sir [Ano91d, Bab22a, Bab22b, Bab30j, Bab32a, Bab33a, Kem79, Sch81, Tur91]. Sismondi [O’B93]. Six [BCKS^{+}83]. sixties [Bar72]. Sketch [Bab30j, Men21a, Men21b]. Sloane [Ano91d]. Smith [O’B93]. Smithsonian [WSC^{+}88, All91, BRG^{+}87, Dod00]. Social [Gig01]. Societies [Wro56]. Society [Bab47c, Enr83, Gre06, Gib19, Her57, HBC^{+}55, Ano91d, Joh91, Ano61, Ano06, Bab56b, Wel89, Wro56]. software [Col98, Don92]. Solar [Bab52]. Sold [Swa96a]. Solutions [Bab17d, Ben16, Rit16]. Solving [Bab17a]. Some [Bab16b, Bab17d, Bab19b, Bab21, Bab30h, Bab27b, Bab30i, Bab53, BM70, BR89, CK89q, CK89r, Mac07, Wol16]. Son [Bab56c]. sorts [Jon91]. Soul [Ano91e, Kou14]. Space [WSC^{+}88]. Spanish [Bab35d]. spanners [BS05]. Specimen [Bab31a]. Specimens [SS57]. speedometers [Jon91]. Sphere [SBET85]. Spirit [Gib19]. Sponsored [Ano48]. spread [Jon91]. Spreadsheets [CKCFR03]. Spufford [Bar96]. stamp [OLCJ91]. Stand [BHW^{+}90]. Standard [Bar00]. Start [Ano10b]. Started [LR67, Ess14b]. State [L.90, Rei68]. Statement [Bab89g, Nic43a, Nic43b]. Statistical [Bab60b, Ano61]. Statistics [Bab53, Bab56a, Tuk72]. Staubbach [HB22]. Steampunk [Sus11, Fis10, Sus11, Sus11]. Stein [AFB^{+}89]. Stephen [Ano91c, Sus11]. Stereomoulded [SS57]. Sterling [Dis91, Spe99, Sus11]. Stern [VBCC83]. Stewart [Bab16b]. Stimulus [Bow93]. Stock [Bar00]. Stokes [BCKS^{+}83]. Stored [CW91, Cer83]. Story [BHW^{+}90, Hun96, Lee95, She06, BCKS^{+}83]. Strandberg [Wor87]. Strandbergs [Bro86]. Strange [MM52]. Strata [Bab56d]. strategic [Bla02]. street [Bab64a]. String [Mad86]. Structure [AFB^{+}89]. struggle [Kou14]. Stuart [O’B93]. students [Pil95]. Studies [GT80, BHB^{+}89]. Study [BHW^{+}90, Wil91c]. Style [Mar08]. subjects [Byr38]. Submarine [Bab55d]. Subrata [Bow15, Nof15]. Subsists [Bab17c, Bab30c]. Substances [BH25, BH30]. substitutions [BK11, BHK12]. suèdoise [Bab55f]. Sulla [Bab34e]. Sumer [CKCFR03]. summary [Ano26]. Sumptuous [OLCJ91]. Sums [Bab19b, Bab30h]. Sun [CW91, Byr04, Bab52]. Sunburst [CW91]. Suppressed [Her31b]. surfaces [Cat81]. survey [Her88, AFB^{+}89]. Swade [CK01b, Cro04b, O’N91, Cro04a]. Swedish [Bab55f]. Sydney [Sus11]. symbolic [Ash96]. symmetric [Cat81]. Symposium [Bow53b, Ano48, MN88]. System [Bab33b, De 06, VBCC83, Byr38]. systems [Pil95, Ano91c].

T [Bab29c, O’B93]. T. [BCKS^{+}83, SBET85]. Table [Arc48, Bab27d, Bab29e, Bab31c, Bab57, Bab15, CK88, Far64, Swa89, Bab89c, Byr38, CK89b, CK89k]. Tables [Ano03, Bab22b, Bab22c, Bab22d, Bab23b, Bab27b, Bab31a, BRG^{+}87, CKCFR03, Far64, GG90, Rob87a, Rob87b, SS57, Bab55f, Bai89, Byr38, Jer89, Mar22, Mer82]. Tabulation [Swa03]. Tabulative [GT80]. TAC [Pil95]. Tale [Wil86, Bro86, Fra84, Fra85, Vam85, Wor87]. talents [Jon91]. Tales [CW91]. Tax [Bab48, Fin52]. Taxation [Bab48, Fin52]. Taylor [Byr38, CW91]. tech [CW91]. Technical [Ste84, Swa96b, OLCJ91]. Technique [Ano91a, Bar72, JC90]. techniques [All91]. Technological [De 06, OLCJ91]. technologie [Kan94, CW91]. Technology [Bow93, CW91, Sus00a, Sus00b, Wil97, Ano91d, Bab97, CH02, GST01, Joh92, Kan94, OLCJ91, Ros94b, SU96, Ugl96, Wen09, Wil85, Ano91a, Ano00, BCKS^{+}83]. Technopolitics [Spe99]. Technotation [Dot97]. Teil [Mar25]. telecommunications [Jen00]. Telegraph [Bab55e]. Telescope [But74]. Temple [Bab34a, Bab34c, Bab46a, Bab47a]. ten [Mac19]. tens [Mar25, Mar86]. tercentenary [Kno15]. Term [Bab26e]. Text [Ste84]. Their [MKW92, Twe93, Bab10c, Byr38, HB13, Mar25, Mar86, She54, She60]. them [Bab89b, Bab10c]. Theorem [Bab19a, Dij86]. Theorems [Bab16b, Bab23a]. Theoretical [Bab23b]. théorie [Mos83]. theory [Gle11, GG05, Mos83, Nim08, Ort07b, Schxx]. things [Jon91]. Thinker [GG92b]. Thinking [Aga18, BCKS^{+}83, Jon16]. thinly [Jon91]. Thomas [O’B93]. Thompson [SBET85]. Thornton [O’B93]. Those [Bab37b, Bab38b]. Thought [Bow53b, KM12, GG05, Sch81]. Thoughts [Bab48, Bab65, Fin52]. Thrilling [Pad15, Sus11]. Thwaites [Bab54a, Bab54b, Bab54c, Bab54d]. TIC [Pil95]. TIC-TAC-TOE [Pil95]. til [Ano93]. Time [Ano08, GG79a, GG79b, Ano91e, Ano11, Cur10, SU96, Ugl96]. times [Jon91]. today [Par96]. TOE [Pil95]. Tokyo [AFI72]. Tomash [OLCJ91, Twe92, Twe93]. tomorrow [FFT^{+}03, FF03b]. too [Jon91, Mor95]. Tooke [O’B93]. Tools [Bab47b, Bab34g]. Torino [Seg08]. Torrens [O’B93]. Torres [Ran82a]. Total [Bab52]. Toy [Fis10]. Toy-de-force [Fis10]. Tracts [BM23]. Tragedies [Ano00]. Traité [Bab33c, Bab34f, Bab34g]. Transatlantic [Wil87, Wil88]. transdisciplinary [CM11]. transfert [CW91]. transformation [GST01]. transformed [Sny11]. translated [Ano91c, Twe91, Twe92, Twe93]. Translation [Men21b, Lov43]. translator [Men21a]. Tratado [Bab35d]. travail [Poi82]. Travels [SCD03]. Treasury [Gre06, Kou14, Ano06]. Treatise [Bab37a, Bab38a, Car69, LaC16, BB36, CK89t, Joh95, Bab34g, Bab35d, Bab67a, Bab09]. Treatises [Bro66]. Trigonometric [GG90]. trigonometrical [Byr38]. Trilogy [MW80, MW72]. Triumphs [Ano00]. True [Com46, Mor95, Wil75]. Truth [Swa04]. Tubes [BCKS^{+}83]. Turin [Seg03]. Turing [THWV88, Aga01, CBSW17, Her88, Swa13, Swa17]. Turn [Yos21]. Turning [Bab47b].

Turns [Mar22]. Tuscany [Bab43b]. Two [Bab29c, Fis14, BWV^{+}88].

U.S [WSC^{+}88]. U.S. [Mar08]. Ubiquitous [Wil95]. Uchronic [Tom02]. Ueber [Bab33d]. Uglow [Bar96]. ultimate [Mul02]. uncle [O’N91, Wil92a]. undecided [Fis17]. understanding [Zen13]. Unerring [Swa89]. Unexpectedly [Bar00]. unification [Schxx]. Unisys [All91]. Unit [Wil81a]. UNIVAC [VBCC83]. Universal [AFB^{+}89, Wat12, Aga01, Her88]. universe [Zen13]. University [Ano48, Asp82, BBKC32, BHB^{+}89, CK89a, Enr80a, Enr80b, GG92a, L.90, O’H82, Ric97]. Unveiled [Bar00]. upon [Bab65]. USA [AFI72]. Use [Bab55e, Bab31b, Bab34g, Bar72, Byr38]. useful [Bab38c]. using [DH90, DH94, DH98, DH04]. Utet [Seg08].

Valves [BCKS^{+}83]. Vannevar [Ano00]. vaporware [Mar22]. varieties [Cat81]. Various [BH25, Bab26a, BH30, Bab60a, Bab67b, Bab89b, CK89o, CK89p]. Vega [Byr38]. Ventilation [Ric39]. vers [Mos83]. version [SGM07]. versus [Swa03]. Victoria [Dis91]. Victorian [Biz03, Dol98, FRW11, HR16, Mat90, Sus94]. videorecording [EEM^{+}05, FFT^{+}03]. View [Bab26a, Bab67b, Bar00, Bab29b, Bab89b, CK89o, CK89p]. Viewpoint [Col98]. Views [Bab51a, Bar00, Poi82]. viii [Enr80a, Enr80b]. Villa [Seg08]. vindications [Biz03]. vision [BS05]. Visions [WSC^{+}88]. Visitors [She54, She60]. vita [BV07, Seg08]. Vol [CK89h, CK89i, CK89j, CK89k, CK89l, CK89m, CK89n, CK89s, Sus11, Bab43a, CK89o, CK89p, CK89q, CK89r, CK89t]. vols [Twe89b]. Volume [EEM^{+}05, O’B93, Twe93, Kno15]. Vorläufer [Schxx]. vs [Kou14].

W [Ano91a, Ano91c, Ber92b, BCKS^{+}83, Enr88, Mer82, Twe92]. War [Sus11, SBET85]. Warming [Ric39]. Was [BW00, Wil00b, Ano91e, Lee95]. Washington [WSC^{+}88, Bab69]. Waterman [BCKS^{+}83]. Way [CW91]. Web [Ess04, Jen00]. Welchman [BCKS^{+}83]. Wellington [Bab89g]. Wesley [Ano91c]. West [CW91, O’B93]. Wheatstone [Kem79]. Wheels [CK01b]. Where [HP16]. Whewell [Bla91, O’B93]. Which [Bab16b, Bab17c, Bab30c, Bab29d, Bab34a, Bab34c, Bab34g, Bab38c, Bab47a, Byr38, GG05]. Whiggism [BHW^{+}90]. Whirlwind [BHW^{+}90, BHW^{+}90]. Who [Sla87, Smi10, Sny11]. William [Dis91, Hun96, L.90, O’B93, Sus11, Ano91c, Bla91, Joh95, Spe99]. Williams [Twe93]. Wilson [O’B93]. Windows [Bab57]. wireless [Jen00]. Wise [Bab33e, Bab56f]. Without [Bab55b, Bab55c]. Wollaston [Bab30j]. Wolstenholme [And99a, And99b]. wonderful [Mor95]. Wood [Bab38c]. Woodhouse [Bec80]. Word [Bab33e, Bab56f, Wil98]. Work [Bab48, Bey79a, Bey79b, Bey79c, BHW^{+}90, Enr78, GG79b, GG90, Gri78, Hil79, HG79, Kem79, Rob79, Ano32a, Ano32b, Ano82a, BdL80, Dub78, Enr80a, Enr80b, Rat52]. Works [Bab43b, CK89a, CK89q, CK89r, CSW90, LR67, Ros90, Swa91a, Twe89a, Twe89b, Ano91e, BH89, BCKR89, CK89e, CK89h, CK89i, CK89j, CK89k, CK89l, CK89m, CK89n, CK89o, CK89p, CK89s, CK89t, CK89f, CK89g, Dot96, Goo02, Hym88, Ort07b]. workshop [OLCJ91, Ait16]. Workstations [CSW90]. World [Ano00, SBET85, Swa91a, Joh91, KT99, Ros88, Sny11, Wil02, CW91]. Writing [Bab55a]. Writings [Boo62, Fra63a, Fra63b, Gri06, MM61a, MM61b].

x [CK01b, Cro04b, L.90, Arc35]. Xenophon [Schxx]. Xenophons [Schxx]. XII [Bab43a]. xiii [Seg08]. XrML [Bar00]. xvii [Twe93]. xxxvii [Twe92].

Year [Bab56a, MB89]. Years [BCKS^{+}83, Ano91e, MT81, OLCJ91]. York [Ano91c, Dis91, Enr80a, Enr80b, Hun96, Pan63, Wil64]. York/London [Enr80a, Enr80b]. you’ll [Ano11]. Youth [Gri10].

Zealand [Tee82, Tee84, Tee86]. Zehnerübertragung [Mar25, Mar86]. Zeichen [Dot97]. Zenith [Bab26c].

References

Almgren:2000:HWC

[AAG^{+}00] Fredrik Almgren, Gunnar Andersson, Torbjörn Granlund, Lars Ivansson, and Staffan Ulfberg. How we cracked the Code Book ciphers. Technical report, ????, ????, October 11, 2000.

40 pp. URL http://frode.
home.cern.ch/frode/crypto/
codebook_solution.pdf;   http:
//www.simonsingh.com/cipher.

htm. See [Sin99].

Aspray:1986:GOH

[AB86] William Aspray and Bruce Bruemmer, editors. Guide to the Oral History Collection of the Charles Babbage Institute. Charles Babbage Institute, the Center for the History of Information Processing, University of Minnesota, Minneapolis, MN, USA, 1986. 110 pp. LCCN ZD16.14 .C53 1986; Z5640 .C46 1986. with the assistance of Hassan Melehy and Thomas Traub.

Aspray:1990:CBC

[ABCK^{+}90] William Aspray, Allan G. Bromley, Martin Campbell-Kelly, Paul E. Ceruzzi, and Michael R. Williams, editors. Computing before computers. Iowa State University Press, Ames, IA, USA, 1990. ISBN 0-8138-0047-1. ix + 266 pp. LCCN QA76 .C5834 1990. URL http:

//www.computerhistory.org/
collections/DocumentArchive/
Documents/Books/Computing\
%20Before\%20Computers/CBC.
html.

Ackroyd:1995:DLL

[Ack95] Peter Ackroyd. Dan Leno and the Limehouse Golem. Minerva, London, UK, 1995. ISBN 0-7493-9659-8. 281 pp. LCCN PR6051.C64 D36 1995.

Adamson:2010:CR

[Ada10] Glenn Adamson, editor. The craft reader. Berg Publishers, Oxford, UK, 2010. ISBN 1-84788-304-4 (hardcover), 1-84788-303-6 (paperback). xii + 641 pp. LCCN https://www.math.utah.edu/pub/tex/bib/adabooks.bib; TT149 .C733 2010.

Aspray:1989:RDI

[AFB^{+}89] William Aspray, Kenneth Flamm, Ralf Bülow, Charles Hall, and Doron Swade. Reviews: Dorfman: Innovation and Market Structure: Lessons from the Computer and Semiconductor Industries; Herken: The Universal Machine: A Half-Century Survey; Lindgren: Glory and Failure: The Difference Engines of Johann Muller, Charles Babbage and Georg and Edvard Scheutz; Stein: Ada: Life and Legacy. Annals of the History of Computing, 11(1): 54–60, January/March 1989. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1989/pdf/a1054.pdf;   http:
//www.computer.org/annals/
an1989/a1054abs.htm.

AFIPS:1972:FUJ

[AFI72] AFIPS, editor. First USA–Japan Computer Conference, Proceedings: October 3–5, 1972, Tokyo, Japan. AFIPS Press, Montvale, NJ, USA, 1972. LCCN QA76 .U2 1972.

Allaby:2002:MS

[AG02] Michael Allaby and Derek Gjertsen. Makers of Science. Makers of science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-521680-6 (set). 96 (vol. 1) pp. LCCN Q141 .A44 2002.

URL http://www.gbv.de/dms/
goettingen/360423353.pdf;
http://www.loc.gov/catdir/
enhancements/fy0611/2001048396-
d.html.

Agar:2001:TUM

[Aga01] Jon Agar. Turing and the universal machine: the making of the modern computer. Revolutions in science. Icon, Cambridge, UK, 2001. ISBN 1-84046-250-7. iv + 153 pp. LCCN QA76.2.T87 A43

Agar:2018:BRM

[Aga18]  Jon Agar.  Book review:

Matthew L. Jones. Reckoning with Matter: Calculating Machines, Innovation, and Thinking about Thinking from Pascal to Babbage. Isis, 109(3):634– 635, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Airy:1856:SCM

[Air56]  George Biddell Airy.  On

Scheutz’s calculating machine. Philosophical Magazine, 12 (series 4)(78):225–226, July 1856. CODEN PHMAA4. ISSN 0031-8086. URL http:

                                        //www.tandfonline.com/doi/
                                        full/10.1080/14786445608642168.

Comments on a report [Wro56] on Scheutz’s engine.

Aitken:2016:BCS

[Ait16] Neil Aitken. Babbage and Carroll in the Silent Workshop, 1867. College Mathematics Journal, 47 (3):215, May 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

                                        //www.tandfonline.com/doi/
                                        abs/10.4169/college.math.j.
                                        47.3.215.

Aldenderfer:1991:AEC

[Ald91] Mark Aldenderfer. The Analytical Engine: Computer simulation and archaeological research. Archaeological Method and Theory, 3(2):195–247, ????

                                        1991. URL http://www.jstor.
                                        org/stable/20170216.

Allison:1991:MAM

[All91] David Allison. Museums and archives. Musée national des techniques. Conservatoire national des arts et Métiers. National Museum of American History, Smithsonian Institution publications and artifacts. Unisys donates Burroughs historical records to Charles Babbage Institute. Digital Equipment Corporation’s historical collection program. Annals of the History of Computing, 13 (4):351–353, October/December 1991. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1991/pdf/a4351.pdf;   http:
//www.computer.org/annals/
an1991/a4351abs.htm.

Apokin:1981:CB

[AMÈ81] I. A. Apokin, L. E. Maĭstrov, and I. S. Èdlin. Charlz Bebidzh. “Nauka”, Moscow, USSR, 1981. 128 pp. 1791– 1871, Nauchno-Biograficheskaya Seriya. [Scientific-Biographic Series].

Andrews:1999:ABCa

[And99a] George E. Andrews. q-analogs of the binomial coefficient congruences of Babbage, Wolstenholme and Glaisher. Discrete Mathematics, 204(1–3):15–25, 1999. CODEN DSMHA4. ISSN 0012-365X (print), 1872-681X (electronic).

Andrews:1999:ABCb

[And99b] George E. Andrews. q-analogs of the binomial coefficient congruences of Babbage, Wolstenholme and Glaisher. Lithos, 151(??):15– 25, June 6, 1999. CODEN ????

                                        ISSN ???? URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0012365X98003641.

Anonymous:1826:PSE

[Ano26] Anonymous. A popular summary of the experiments of Messrs. Barlow, Christie, Babbage and Herschel: on the magnesium of iron and other metals, as exhibited by rotation. Edinburgh, Scotland, 1826. 13–19 pp. LCCN Micro-fiche 960; Q111 .H35TN799.M2.

Anonymous:1832:SPBa

[Ano32a] Anonymous. Selections from Professor Babbage’s work on the economy of machinery and manufactures. Journal of The Franklin Institute, 14(5):355– 356, November 1832. CO-DEN JFINAB. ISSN 0016-0032 (print), 1879-2693 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0016003232914690.

Anonymous:1832:SPBb

[Ano32b] Anonymous. Selections from Professor Babbage’s work “On the economy of machinery and manufactures”. Journal of The Franklin Institute, 14(6):414– 415, December 1832. CO-DEN JFINAB. ISSN 0016-0032 (print), 1879-2693 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0016003232903111.

Anonymous:1871:CBc

[Ano71a] Anonymous. Charles Babbage.

Scientific American, 25(26):402– 403, December 23, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/
scientificamerican/journal/
v25/n26/pdf/scientificamerican12231871-
402a.pdf.

Anonymous:1871:CBH

[Ano71b] Anonymous. Charles Babbage, 1792–1871, 1871. URL http:

//hdl.loc.gov/loc.pnp/cph.

3b13540. 1 print, head and shoulders, facing left, from The Illustrated London News, 1871 Nov. 4, v. 59, p. 424.

Anonymous:1871:CBOa

[Ano71c] Anonymous. Charles Babbage [obituary]. Nature, 5(106):28– 29, November 9, 1871. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/
nature/journal/v5/n106/pdf/
005028a0.pdf.

Anonymous:1871:CBOb

[Ano71d] Anonymous. Charles Babbage [obituary]. Scientific American, 25(26):402–403, December 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/
scientificamerican/journal/

                                        v25/n26/pdf/scientificamerican12231871-
                                        402a.pdf.

Anonymous:1890:CCL

[Ano90] Anonymous. Catalogue of the Crawford Library of the Royal Observatory, Edinburgh. Published by authority of Her Majesty’s government, Edinburgh, Scotland, 1890. viii + 497 pp. LCCN Z5156 .E23.

Anonymous:1930:CSD

[Ano30] Anonymous. “A Classic of Science”: The Difference Engine. The Science News-Letter, 18 (495):218–220, October 4, 1930. CODEN SNLEAI. ISSN 0096-4018 (print), 2326-1285 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/3905897.

Anonymous:1933:CBF

[Ano33] Anonymous. Charles Babbage, F.R.S. Nature, 132(3347):959, December 23, 1933. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

                                        URL http://www.nature.com/
                                        nature/journal/v132/n3347/
                                        pdf/132959a0.pdf.

Anonymous:1948:PSL

[Ano48] Anonymous, editor. Proceedings of a Symposium on Large-Scale Digital Calculating Machinery: Jointly Sponsored by The Navy Department Bureau of Ordnance and Harvard University at The Computation Laboratory 7–10 January 1947, volume 16 of The Annals of the Computation Laboratory of Harvard University. Harvard University Press, Cambridge, MA, USA, 1948. LCCN QA75 .S96 1947.

Anonymous:1961:BNR

[Ano61] Anonymous. Babbage’s note respecting the origin of the Statistical Society. Journal of the Royal Statistical Society. Series A (General), 124(4):546, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342933.

Anonymous:1974:CBF

[Ano74] Anonymous. Charles Babbage:

father of the computer, 1974. Videocassette (28 min.). Narrator, Bill Woodward; director Warren F. Miller.

Anonymous:1977:BAE

[Ano77] Anonymous. Babbage Analytical Engine, 1977. URL http://

hdl.loc.gov/loc.pnp/ppmsca.

24133. 1 photographic print [between 1971 and 1977].

Anonymous:1980:CBI

[Ano80]  Anonymous.  Charles Bab-

bage Institute selects permanent home. Science, Technology and Human Values, 6(33):58, Autumn 1980. ISSN 0162-2439 (print),1552-8251 (electronic).

URL http://www.jstor.org/
stable/689292.

Anonymous:1981:CQD

[Ano81] Anonymous. Comments, queries, and debate: Query on Babbage.

Annals of the History of Computing, 3(3):288, July/September 1981. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1981/pdf/a3288.pdf;   http:
//www.computer.org/annals/
an1981/a3288abs.htm.

Anonymous:1982:LWC

[Ano82a] Anonymous. The life and work of Charles Babbage. Historia Mathematica, 9(4):485, November 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0315086082901082.

Anonymous:1982:NNPb

[Ano82b] Anonymous. News and notices: Pioneer Day to Celebrate Harvard at 1983 NCC; Meeting on Charles Babbage. Annals of the History of Computing, 4(4):373, October/December 1982. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1982/pdf/a4373.pdf;   http:
//www.computer.org/annals/
an1982/a4373abs.htm.

Anonymous:1990:BCC

[Ano90] Anonymous. The Bicentennial Conference on Computing: a forthcoming event to commemorate the 200th anniversary of the birth of Charles Babbage London, 1–3 July 1991. Historia Mathematica, 17(3):259, August 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/
article/pii/031508609090004W.

Anonymous:1991:ARB

[Ano91a] Anonymous. Article review:

Forming Holes in Brittle Materials. An Antique Technique and its Report by Charles Babbage, 1792–1871: W. Johnson, S. Chandrasekar, Journal of Materials Processing Technology, 21(1), pp. 43–50. (Jan 1990). Precision Engineering, 13(1):70, January 1991. CODEN ????

ISSN ???? URL http://www.
sciencedirect.com/science/
article/pii/014163599190272K.

Anonymous:1991:BE

[Ano91b] Anonymous. Babbage’s engines.

The Lancet (London, England), 338(8758):24, July 6, 1991. CO-DEN LANCAO. ISSN 0140-6736 (print), 1474-547X (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/014067369190012E.

Anonymous:1991:HCC

[Ano91c] Anonymous. History of computers and computing: Reviews of William Aspray, John von Neumann and the Origins of modern Computing, (Cambridge: MIT Press, 1990 — $35.00, ISBN 0-262-01121-2, 376 pp.); Stephen G. Nash, editor, A History of Scientific Computing (New York: Association for Computing Machinery [ACM] Press; co-published by Addison-Wesley, 1990 — price not given, ISBN 0-201-50814-1, 359 pp.); Michael Lindgren, Glory and Failure: The Difference Engines of Johann Müller, Charles Babbage and Edvard Scheutz [translated by Craig G. McKay] (Cambridge: MIT Press, 1990 — $45.00, ISBN 0-262-12146-8, 414 pp.); Emerson W. Pugh, Lyle R. Johnson, and John H. Palmer, IBM’s 360 and Early 370 Systems (Cambridge: MIT Press, 1991 — $37.50, ISBN 0-262-16123-0, 819 pp.) . Communication Booknotes Quarterly, 22 (4):78, ???? 1991. ISSN 1094-8007 (print), 1532-6896 (elec-

tronic).  URL http://www.
tandfonline.com/doi/abs/10.
1080/10948009109487995.

Anonymous:1991:HST

[Ano91d] Anonymous. The History of science and technology: a listing and guide to the microfilm collection: Papers of Sir Hans Sloane, 1660–1754: Papers of Sir Joseph Banks, 1743–1820: Papers of Charles Babbage, 1791– 1871: Science and society, 1660– 1773.: Science and society, 1660–1773. Adam Matthew Publications, Wiltshire, UK, 1991. ISBN 1-85711-016-1 (Ser.1, pt.1), 1-85711-052-8 (Ser.2, pt.1-3), 1-85711-128-1 (Ser.2, pt.4), 1-85711-101-X (Ser.3, pt.1). ???? pp. LCCN Z7401 .H57 1991.

Anonymous:1991:HMW

[Ano91e] Anonymous. How the machine works: Seeking the soul of an old machine: Charles Babbage’s Difference Engine is ready to run — built for the first time 150 years after it was designed. Science, 252(5011):1370–1371, June 7, 1991. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/
252/5011/1370.2.full.

Anonymous:1993:FBT

[Ano93] Anonymous. Fra Babbage til Pentium: en datahistoriebok. (Norwegian) [From Babbage to Pentium: a computer history book]. Computerworld Norge, 11 (21):130, 1993.

Anonymous:1996:BHC

[Ano96]  Anonymous.  A brief his-

tory of the computer. New York Times, ??(??):D3, February 19, 1996. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL

http://search.proquest.com/
docview/109624296.

Anonymous:2000:RPA

[Ano00] Anonymous. Reviews: Peggy Aldrich Kidwell: The Cogwheel Brain: Charles Babbage and the Quest to Build the First Computer: Endless Frontier: Vannevar Bush, Engineer of the American Century: Dorothy Hodgkin: a Life: ENIAC: The Triumphs and Tragedies of the World’s First Computer: Inventing the Internet: Meaning in Technology. IEEE Annals of the History of Computing, 22(4):62–63, October–December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        dlib.computer.org/an/books/
                                        an2000/pdf/a4062.pdf.

Anonymous:2003:BLT

[Ano03] Anonymous. Babbage’s logarithm tables, 2003. URL http:

                                        //www.archive.org/details/
                                        BabbagesLogarithmTables.

Anonymous:2006:MBI

[Ano06] Anonymous. Mr. Babbage’s invention. Copies of the correspondence between the lords commissioners of His Majesty’s Treasury and the president and council of the Royal Society, relative to an invention of Mr. Babbage. Pro-Quest LLC, ????, 2006. LCCN ????

Anonymous:2008:CHM

[Ano08] Anonymous. The Computer History Museum debuts Charles Babbage’s Difference Engine No. 2, on display for the first time in North America. Web press release, May 5, 2008. URL https:/

                                        /computerhistory.org/press-
                                        releases/babbage-engine-exhibit/

.

Anonymous:2010:MBM

[Ano10a] Anonymous. Mini Babbage machine promises energy efficient computing. New Scientist, 205 (2753):19, March 24, 2010. CO-DEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0262407910607014.

Anonymous:2010:RSS

[Ano10b] Anonymous. Random samples: Start the Engine. Science, 330(6002):301, October 15, 2010. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:/

/plan28.org/;   http://www.
sciencemag.org/content/330/
6002/301.1.full.

Anonymous:2011:BBB

[Ano11] Anonymous, editor. The best business books ever: the most influential management books you’ll never have time to read. Basic Books, New York, NY, USA, revised edition, 2011. ISBN 0-465-02236-7 (paperback), 0-465-02634-6 (e-book). ix + 269 pp. LCCN Z7164.O7 B323 2011.

Anonymous:2016:ICCb

[Ano16]  Anonymous.  2016 IEEE-CS

Charles Babbage Award. Computing in Science and Engineering, 18(6):c3, November/December 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/cs/
2016/06/mcs20160600c3.pdf.

Anonymous:2017:ICC

[Ano17]  Anonymous.  2017 IEEE-CS

Charles Babbage Award. IEEE Annals of the History of Computing, 39(3):8, July/September 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL https://www.

                                        computer.org/csdl/mags/an/
                                        2017/03/man2017030008.pdf.

Anonymous:2018:ICCa

[Ano18a] Anonymous. 2019 IEEE-CS Charles Babbage Award. IEEE Software, 35(4):43, July/August 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Anonymous:2018:ICCb

[Ano18b] Anonymous. 2019 IEEE-CS Charles Babbage Award. Computer, 51(5):c4, May 2018. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/co/
                                        2018/05/mco20180500c4.pdf.

Anonymous:2018:ICCc

[Ano18c] Anonymous. 2019 IEEE-CS Charles Babbage Award. Computer, 51(6):29, June 2018. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/co/
                                        2018/06/mco2018060029.pdf.

Anonymous:2018:OH

[Ano18d] Anonymous. Our history. British Science Association Web site.,

                                        2018.  URL https://www.
                                        britishscienceassociation.
                                        org/history.

Anonymous:2021:ICC

[Ano21a] Anonymous. 2022 IEEE CS: Charles Babbage Award: Call for award nominations. IEEE Annals of the History of Computing, 43(3):5, July/September 2021. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Anonymous:2021:CBA

[Ano21b] Anonymous. Charles Babbage Award. IEEE Micro, 41(5):14, September/October 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2021:E

[Ano21c] Anonymous. The engines.

Computer History Museum, Mountain View, CA, USA,

2021.  URL https://www.
computerhistory.org/babbage/
engines/.

Anonymous:2022:ICC

[Ano22] Anonymous. 2023 IEEE CS Charles Babbage Award. Computing in Science and Engineering, 24(1):87, January/February 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Archibald:1935:BMX

[Arc35] R. C. Archibald. Bibliographia de mathematicis X: Charles Babbage. Scripta Mathematica, 3 (??):266–267, ???? 1935. ISSN 0036-9713.

Archibald:1947:PGS

[Arc47] Raymond C. Archibald. P. G.

Scheutz, publicist, author, scientific mechanician, and Edvard Scheutz, engineer—biography and bibliography. Mathematical Tables and Other Aids to Computation, 2(18):238–245, April 1947. CODEN MTTCAS. ISSN 0891-6837. Original journal has only ‘R. C. A.’ as author: Goldstine [Gol72, page 16–17] confirms that the author of this paper is Raymond C. Archibald.

Archibald:1948:MTM

[Arc48] Raymond Clare Archibald. Mathematical Table Makers: Portraits, Paintings, Busts, Monuments: Bio-Bibliographical Notes, volume 3 of The Scripta Mathematica Studies. Scripta Mathematica, New York, NY, USA, 1948. vii + 82 pp.

URL https://ia802706.us.
archive.org/6/items/mathematicaltabl00arch/
mathematicaltabl00arch.pdf.

Ashworth:1994:CEB

[Ash94] William J. Ashworth. The calculating eye: Baily, Herschel, Babbage and the business of astronomy. British Journal for the History of Science, 27(4): 409–441, December 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4027624.

Ashworth:1996:MES

[Ash96] William J. Ashworth. Memory, efficiency, and symbolic analysis: Charles Babbage, John Herschel, and the industrial mind. Isis, 87(4):629–653, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic).  URL http://www.
jstor.org/stable/235196.

Ashworth:2002:BRB

[Ash02] William J. Ashworth. Book review: Bruce Collier; James MacLachlan: Charles Babbage and the Engines of Perfection. Isis, 93(1):127–128, March 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/10.
1086/343300;   http://www.
jstor.org/stable/pdfplus/10.
1086/343300.pdf.

Aspray:1982:BRB

[Asp82] William Aspray. Book review:

Charles Babbage, A Scientist and Reformer, by Anthony Hymanm, Princeton: Princeton University Press. Science, 218(4573):672– 673, November 12, 1982. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic).  URL http://www.
jstor.org/stable/1688618.

Aspray:2007:LCH

[Asp07] William Aspray. Leadership in computing history: Arthur Norberg and the Charles Babbage Institute. IEEE Annals of the History of Computing, 29(4): 16–26, October/December 2007. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Atkinson:2000:DEE

[Atk00] Paul Atkinson. The (in)Difference engine: Explaining the disappearance of diversity in the design of the personal computer. Journal of Design History, 13(1): 59–72, ???? 2000. ISSN 0952-4649 (print), 1741-7279 (electronic).

Augarten:1984:BBI

[Aug84] Stan Augarten. Bit by Bit: an Illustrated History of Computers. Ticknor and Fields, New York, NY, USA, 1984. ISBN 0-89919-268-8 (hardcover), 0-89919-302-1 (paperback). 324 pp. LCCN QA76.17 .A94 1984. US$29.85 (hardcover), US$17.95 (paperback).

Austrian:1982:HHF

[Aus82] Geoffrey D. Austrian. Herman Hollerith — Forgotten Giant of Information Processing. Columbia University Press, New York, NY, USA, 1982. ISBN 0-231-05146-8. xvi + 418 pp. LCCN QA76.2.H64 .A97.

B:1988:GM

[B.88] T. M. B. Ghost in the machine. Scientific American, 259 (6):31, December 1988. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

                                        URL http://www.nature.com/
                                        scientificamerican/journal/
                                        v259/n6/pdf/scientificamerican1288-
                                        31a.pdf.

Babbage:1813:CP

[Bab13] Charles Babbage. On continued products. Memoirs of the Analytical Society [Cambridge], page ??, 1813. Reprinted in [Sch81].

Babbage:1815:ETC

[Bab15] C. Babbage. An essay towards the calculus of functions. Philosophical Transactions of the Royal Society of London, 105 (??):389–423, ???? 1815. CO-DEN PTRSAV. ISSN 0370-2316 (print), 2053-9207 (elec-

tronic).  URL http://www.
jstor.org/stable/107377.

Babbage:1816:ETC

[Bab16a] C. Babbage. An essay towards the calculus of functions. Part II. Philosophical Transactions of the Royal Society of London, 106 (??):179–256, ???? 1816. CO-DEN PTRSAV. ISSN 0370-2316 (print), 2053-9207 (elec-

tronic).  URL http://www.
jstor.org/stable/107524.

Babbage:1816:DSD

[Bab16b] Charles Babbage. Demonstrations of some of Dr. Matthew Stewart’s general theorems; to which is added, an account of some new properties of the circle. Journal of Science, 1(??): 6–24, ???? 1816. CODEN ???? ISSN ????

Babbage:1817:AEM

[Bab17a] Charles Babbage. An account of Euler’s method of solving a problem relative to the move of the knight at the game of chess. Journal of Science, 3(??):72–77, ???? 1817. CODEN ???? ISSN ????

Babbage:1817:NRE

[Bab17b] Charles Babbage. Note respecting elimination. Journal of Science, 3(??):355–357, ???? 1817. CODEN ???? ISSN ????

Babbage:1817:OAW

[Bab17c] Charles Babbage. Observations on the analogy which subsists between the calculus of functions and other branches of analysis. Philosophical Transactions of the Royal Society of London, 107(1/3):197–216, ???? 1817. CODEN PTRSAV. ISSN 0370-2316 (print), 2053-9207 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/107581.

Babbage:1817:SSP

[Bab17d] Charles Babbage. Solutions of some problems by means of the calculus of functions. Journal of Science, 2(??):371–379, ???? 1817. CODEN ???? ISSN ????

Babbage:1819:DTR

[Bab19a] Charles Babbage. Demonstration of a theorem relating to prime numbers. Edin. Phil. Jrl, 1(??): 46–9, ???? 1819. CODEN ???? ISSN ????

Babbage:1819:SNM

[Bab19b] Charles Babbage. On some new methods of investigating the sums of several classes of infinite series. Philosophical Transactions of the Royal Society of London, 109(1/3):249–282, ???? 1819. CODEN PTRSAV. ISSN 0370-2316 (print), 2053-9207 (electronic). URL http://www.

jstor.org/stable/107505.

Babbage:1821:ESQ

[Bab21] Charles Babbage. An examination of some questions connected with games of chance. Trans. Roy Soc., 9(??):153–177, ???? 1821. CODEN ???? ISSN ????

Babbage:1822:LSHa

[Bab22a] Charles Babbage. A letter to Sir Humphry Davy. Brewster’s Edinburgh Journal of Science, 8 (??):122–??, 1822.

Babbage:1822:LSHb

[Bab22b] Charles Babbage. A letter to Sir Humphry Davy, Bart. PRS, on the application of machinery to the purpose of calculating and printing mathematical tables. London, UK, 1822.

Babbage:1822:NRA

[Bab22c] Charles Babbage. A note respecting the application of machinery to the calculation of astronomical tables. Memoirs of the Royal Astronomical Society, 1(2):309, June 2, 1822. CODEN ???? ISSN ???? URL https:

//babel.hathitrust.org/cgi/
pt?id=mdp.39015077097874&view=
1up&seq=89.

Babbage:1822:OAM

[Bab22d] Charles Babbage. Observations on the application of machinery to the computation of mathematical tables. Mem. Astron. Soc., 1(??):311–314, ???? 1822. CO-DEN ???? ISSN ????

Babbage:1822:ONE

[Bab22e] Charles Babbage. Observations on the notation employed in the calculus of functions. Transactions of the Cambridge Philosophical Society, 1(??):63– 76, ???? 1822. CODEN ???? ISSN ????

Babbage:1822:SNR

[Bab22f] Charles Babbage. The science of number reduced to mechanism. Reprinted in [CK89k]., 1822.

Babbage:1823:AAD

[Bab23a] Charles Babbage. On the application of analysis to the discovery of local theorems and porisms. Trans. Roy. Son. Edin., 9(??): 337–52, ???? 1823. CODEN ???? ISSN ????

Babbage:1823:TPM

[Bab23b] Charles Babbage. On the theoretical principles of the machinery for calculating tables. Edin. Phtl. Jrl., 8(??):122–128, ???? 1823. CODEN ???? ISSN ????

Babbage:1824:OMH

[Bab24a] Charles Babbage. Observations on the measurement of heights by the barometer. Edin. Jrl. Sci., ?? (??):??, ???? 1824. CODEN ???? ISSN ????

Babbage:1824:RPL

[Bab24b] Charles Babbage. Rates of the Protector Life Assurance Company, 1824. Privately printed, London, UK.

Babbage:1826:CVV

[Bab26a] Charles Babbage. A Comparative View of the Various Institutions for the Assurance of Lives. J. Mawman, London, UK, 1826. xxxi + 170 + 27 + 1 pp. LCCN ???? German translation, Weimar, 1827.

Babbage:1826:MES

[Bab26b] Charles Babbage. On a method of expressing by signs the action of machinery. Philosophical Transactions of the Royal Society of London, 116(1/3): 250–265, ???? 1826. CO-DEN PTRSAV. ISSN 0370-2316 (print), 2053-9207 (elec-

tronic).  URL http://www.
jstor.org/stable/107813.

Babbage:1826:NZM

[Bab26c] Charles Babbage. On a new zenith micrometer. Mem. Astron. Soc., 2(??):101–103, ???? 1826. CODEN ???? ISSN ????

Babbage:1826:EMR

[Bab26d] Charles Babbage. On electrical and magnetic rotations. Philosophical Transactions of the Royal Society of London, 116 (1/3):494–528, ???? 1826. CO-DEN PTRSAV. ISSN 0370-2316 (print), 2053-9207 (elec-

tronic).  URL http://www.
jstor.org/stable/107825.

Babbage:1826:DGT

[Bab26e] Charles Babbage. On the determination of the general term of a new class of infinite series.

Transactions of the Cambridge Philosophical Society, 2(??):217– 25, ???? 1826. CODEN ???? ISSN ????

Babbage:1826:DB

[Bab26f] Charles Babbage. On the diving bell. In E. Smedley et al., editors, Encyclopaedia Metropolitana, volume 18, pages 157–167. ????, ????, 1826. LCCN ????

Babbage:1826:ESB

[Bab27a] Charles Babbage. Evidence on savings banks, before a Committee of the House of Commons. Reports from the Committees of the House of Commons, III(??): ??, ???? 1826–1827. CODEN ???? ISSN ????

Babbage:1827:NRS

[Bab27b] Charles Babbage. Notice respecting some errors common to many tables of logarithms. Mem. Astron. Soc., 3(??):65–67, ???? 1827. CODEN ???? ISSN ????

Babbage:1827:ISM

[Bab27c] Charles Babbage. On the influence of signs in mathematical reasoning. Transactions of the Cambridge Philosophical Society, 2(??):325–377, ???? 1827. CO-DEN ???? ISSN ????

Babbage:1827:TLN

[Bab27d] Charles Babbage. Table of logarithms of the natural numbers from 1 to 108000. London, UK, 1827. ?? pp.

Babbage:1829:AGC

[Bab29a] Charles Babbage. Account of the Great Congress of Philosophers at Berlin on the 18th September 1828. Edin. Jrl. Sci., 10(??):225– 34, ???? 1829. CODEN ???? ISSN ????

Babbage:1829:IVP

[Bab29b] Charles Babbage. An introductory view of the principles of manufactures. In E. Smedley et al., editors, Encyclopaedia Metropolitana, volume VIII, page ???? ????, ????, 1829. LCCN ????

Babbage:1829:LRH

[Bab29c] Charles Babbage. A letter to the Right Hon. T. P. Courtenay, on the proportionate number of births of the two sexes under different circumstances. Edin. Jrl. Sci., NS, 1(??):85–104, ???? 1829. CODEN ???? ISSN ????

Babbage:1829:GPW

[Bab29d] Charles Babbage. On the general principles which regulate the application of machinery to manufacture and the mechanical arts. In E. Smedley et al., editors, Encyclopaedia Metropolitana, volume 8, pages 1–84. ????, ????, 1829.

Babbage:1829:TLN

[Bab29e] Charles Babbage. Table of logarithms of the natural numbers, from 1 to 108 000. ????, ????, 1829. 20 + 201 pp. LCCN ????

Babbage:1830:ETC

[Bab30a] C. Babbage. An essay towards the calculus of functions [abstract]. Abstracts of the Papers Printed in the Philosophical Transactions of the Royal Society of London, 2(??):29, ???? 1830.

Babbage:1830:MQ

[Bab30b] Charles Babbage. Mathematical questions. Mathematical Repository, 5 (new series)(??):51, 63–4, 66–7, 71–2, 178–80, ???? 1830. CODEN ???? ISSN ????

Babbage:1830:OAW

[Bab30c] Charles Babbage. Observations on the analogy which subsists between the calculus of functions and other branches of analysis. [abstract]. Abstracts of the Papers Printed in the Philosophical Transactions of the Royal Society of London, 2(1/3):69, ???? 1830.

Babbage:1830:MES

[Bab30d] Charles Babbage. On a method of expressing by signs the action of machinery. [abstract]. Abstracts of the Papers Printed in the Philosophical Transactions of the Royal Society of London, 2 (1/3):273–274, ???? 1830.

Babbage:1830:EMR

[Bab30e] Charles Babbage. On electrical and magnetic rotations. [abstract]. Abstracts of the Papers Printed in the Philosophical Transactions of the Royal Society of London, 2(1/3):287–288, ???? 1830.

Babbage:1830:N

[Bab30f] Charles Babbage. On notation.

In ????, editor, Edinburgh Encyclopaedia, volume 15, pages 394– 399. ????, ????, 1830.

Babbage:1830:P

[Bab30g] Charles Babbage. On porisms.

In ????, editor, Edinburgh Encyclopaedia, volume 17, pages 106– 14. ????, ????, 1830.

Babbage:1830:SNM

[Bab30h] Charles Babbage. On some new methods of investigating the sums of several classes of infinite series. [abstract]. Abstracts of the Papers Printed in the Philosophical Transactions of the Royal Society of London, 2(1/3):115, ???? 1830.

Babbage:1830:RDS

[Bab30i] Charles Babbage. Reflections on the decline of science in England, and some of its causes. London, UK, 1830. ?? pp.

Babbage:1830:SPC

[Bab30j] Charles Babbage. Sketch of the philosophical characters of Dr. Wollaston and Sir Humphry Davy. London, UK, 1830.

Babbage:1831:SLT

[Bab31a] Charles Babbage. Specimen of Logarithm Tables. B. Fellowes, London, UK, 1831. ?? pp. 21 volumes, but on a single copy was printed; it is now in the Crawford Library.

Babbage:1831:EPO

[Bab31b] Charles Babbage. Sur l’emploi plus ou moins fréquent des mêmes lettres dans les différentes langues. (French) [On the more or less frequent use of letters in different languages]. Corr. mathématique et physique, 7(??): 135–137, ???? 1831. CODEN ???? ISSN ???? Letter to L. A. J. Quetelet.

Babbage:1831:TLN

[Bab31c] Charles Babbage. A Table of Logarithms of the Natural Numbers from 1–108 000, on Different Coloured Papers. B. Fellowes, London, UK, 1831. ?? pp. Single copy printed; 28 volumes are in the Crawford Library, but Babbage said that 35 were printed.

Babbage:1832:LSD

[Bab32a] Charles Babbage. Letter to Sir David Brewster, on the advantage of a collection of the constants of nature and art. Brewster’s Edinburgh Journal of Science, 6(??):334–??, ???? 1832.

Babbage:1832:ACN

[Bab32b] Charles Babbage. On the advantage of a collection of numbers, to be entitled the constants of nature and art. Edin. Jrl. Sci., 6 (new series)(??):334–340, ???? 1832. CODEN ???? ISSN ????

Babbage:1832:EMMb

[Bab32c] Charles Babbage. On the economy of machinery and manufactures. C. Knight, London, UK, second edition, 1832. xxiv + 387 pp. LCCN ????

Babbage:1832:EMMc

[Bab32d] Charles Babbage. On the economy of machinery and manufactures. C. Knight, London, UK, third edition, 1832. xxiv + 392 pp. LCCN ????

Babbage:1833:LSD

[Bab33a] Charles Babbage. Letter to Sir David Brewster, on the advantage of a collection of the constants of nature and art. Cambridge, UK, 1833. ?? pp. LCCN ???? Reprint of [Bab32a].

Babbage:1833:CNS

[Bab33b] Charles Babbage. On currency, on a near system of manufacturing, and on the effect of machinery on human labour. Pall Mall East, London, UK, 1833. ?? pp. Three chapters from [Bab32d].

Babbage:1833:TCM

[Bab33c] Charles Babbage. Traité sur l’économie des machines et des manufactures. (French) [On the economy of machinery and manufactures]. Bachelier, Paris, France, 1833. xvi + 515 pp. LCCN ????

Babbage:1833:UMF

[Bab33d] Charles Babbage. Ueber Maschinen- und Fabrikenwesen. (German) [On the economy of machinery and manufactures]. Im Verlage der Stuhrschen Buchhandlung, Berlin, Germany, 1833. iii + 462 pp. LCCN ????

Babbage:1833:WW

[Bab33e] Charles Babbage. A word to the wise. London, UK, July 1833. 16 pp. Reprinted in [Bab56f].

Babbage:1834:APE

[Bab34a] Charles Babbage. Abstract of a paper entitled: Observations on the Temple of Serapis at Pozzuoli; with Remarks on Certain Causes which May Produce Geological Cycles of Great Extent. London, UK, 1834. ?? pp. See also [Bab34c], which appears to be the same as this, with another title. Full paper published in [Bab47a].

Babbage:1834:ABC

[Bab34b] Charles Babbage. Articles on Babbage’s calculating engine by Babbage and others. ????, Edinburgh, Scotland, 1834. 300 pp. LCCN ????

Babbage:1834:OTS

[Bab34c] Charles Babbage. Observations on the Temple of Serapis at Pozzuoli, near Naples; with remarks on certain causes which may produce geological cycles of great extent. Proceedings of the Geological Society of London, 2:72–76, 1834. URL http:

//www.archive.org/download/
proceedings218331838geol/proceedings218331838geol.

pdf. Read [in abbreviated form] 12 March 1834 by W. H. Fitton. The full article did not appear until 13 years later; see [Bab47a].

Babbage:1834:SCM

[Bab34d] Charles Babbage. Science économique des manufactures. (French) [On the economy of machinery and manufactures]. Dondey-Dupré, Paris, France, 1834. xxiii + 392 pp. LCCN ????

Babbage:1834:SED

[Bab34e] Charles Babbage. Sulla economia delle macchine e delle manifatture. (Italian) [On the economy of machinery and manufactures]. G. Piatti, Firenze[, Italy], 1834. 311 pp. LCCN ????

Babbage:1834:TCM

[Bab34f] Charles Babbage. Traité sur l’économie des machines et des manufactures. (French) [On the economy of machinery and manufactures]. H. Dumont, Bruxelles, Belgium, third edition, 1834. viii + 468 pp. LCCN ????

Babbage:1834:TEA

[Bab34g] Charles Babbage. Traité sur les effets et les avantages qui résultent de l’emploi des outils et des machines. (French) [Treatise on the effects and advantages which result from the use of tools and machines]. Société d’encouragement pour l’instruction élémentaire, ????, France, 1834. 270 pp. LCCN ????

Babbage:1835:LMS

[Bab35a] Charles Babbage. Letter from Mr. Abraham Sharp to Mr. J. Crosthwait, Hoxton, Feb. 2, 1721–2, decyphered by C. Babbage, Esq. from the original letter in shorthand. In Francis Baily, editor, An Account of the Revd John Flamsteed, the First Astronomer-Royal, pages 348, 390. Admiralty, London, UK, 1835.

Babbage:1835:ACN

[Bab35b] Charles Babbage. On the advantage of a collection of numbers, to be entitled the constants of nature and art. Compte Rendus des Travaux du Congrès Général de Statistique, ??(??):222–230, ???? 1835. CODEN ???? ISSN ???? Expanded version of [Bab32b].

Babbage:1835:EMM

[Bab35c] Charles Babbage. On the Economy of Machinery and Manufactures. Charles Knight, Pall Mall East, London, UK, fourth edition, 1835. xxiv + 408 pp. LCCN ????

Babbage:1835:TMP

[Bab35d] Charles Babbage. Tratado de mecánica práctica y economia politica. (Spanish) [Treatise on practical mechanics and political economics]. Martinez, Madrid, Spain, third edition, 1835. xx + 356 + 12 + 1 pp. LCCN ???? Translation to Spanish by José Diez Imbrechts.

Babbage:1835:LMQ

[Bab35e] Charles Babbage. Une lettre à M. Quetelet de M. Ch. Babbage relativement à la machine à calculer. (French) [A letter to Mr. Quetelet from Mr. Ch. Babbage relative to the calculating machine]. Acad. Roy. Bruxelles, Bulletins, 2(??):123–126, ???? 1835. CODEN ???? ISSN ????

Babbage:1837:NBT

[Bab37a] Charles Babbage. The Ninth Bridgewater Treatise: a Fragment. John Murray, London, UK, 1837. ?? pp.

Babbage:1837:ISR

[Bab37b] Charles Babbage. On impressions in sandstone resembling those of horses’ hoofs. Philosophical Magazine, 10 (series 3)(??): 474, ???? 1837. CODEN PH-MAA4. ISSN 0031-8086.

Babbage:1837:MPC

[Bab37c] Charles Babbage. On the mathematical powers of the calculating engine. Unpublished manuscript Buxton MS7 held by Museum of the History of Science, Oxford, UK. Reprinted in [Ran82b, §2.1, pages 19–54], December 26, 1837.

Babbage:1838:NBT

[Bab38a] Charles Babbage. The Ninth Bridgewater Treatise: a Fragment. John Murray, London, UK, second edition, 1838. ?? pp.

Babbage:1838:ISR

[Bab38b] Charles Babbage. On impressions in sandstone resembling those of horses’ hoofs. Proceedings of the Geological Society of London, 2:439, 1838. CODEN ???? ISSN ???? URL http:

//www.biodiversitylibrary.
org/item/105665#page/355/mode/
1up.

Babbage:1838:SAM

[Bab38c] Charles Babbage. Short account of a method by which engraving on wood may be rendered more useful for the illustration and description of machinery. Report of Meeting of British Association of Newcastle, ??(??):154–??, ???? 1838.

Babbage:1839:LMB

[Bab39] Charles Babbage. Letter from Mr. Babbage to the Members of the British Association for the Promotion of Science. London, UK, ???? 1839. ?? pp.

Babbage:1843:AMM

[Bab43a] Charles Babbage. Addition to the memoir of M. Menabrea on the Analytical Engine. Scientific memoirs, vol. III. Part XII. p. 666. Philosophical Magazine, 23(151):235–239, ???? 1843. CODEN PHMAA4. ISSN 0031-8086. URL http://www.

tandfonline.com/doi/abs/10.
1080/14786444308644724.

Babbage:1843:DBA

[Bab43b] Charles Babbage. Description of the boracic acid works of Tuscany. In Handbook for Travellers in Central Italy, pages 178–179. John Murray, London, UK, ???? 1843.

Babbage:1846:OTS

[Bab46a] Charles Babbage. Observations on the Temple of Serapis at Pozzuoli, near Naples. Proceedings of the Geological Society of London, ??(??):??, ???? 1846. URL http:

//www.biodiversitylibrary.
org/item/109652.

Babbage:1846:EMM

[Bab46b] Charles Babbage. On the economy of machinery and manufactures. J. Murray, London, UK, fourth edition, 1846. xxiv + 408 pp. LCCN ????

Babbage:1847:OTS

[Bab47a] Charles Babbage. Observations on the Temple of Serapis at Pozzuoli, near Naples, with remarks on certain causes which may produce geological cycles of great extent. The Quarterly Journal of the Geological Society of London, 3:186–217, 1847. URL http:

//www.archive.org/download/
quarterlyjournal31847geol/
quarterlyjournal31847geol.

pdf. Read 12 March 1834. First publication of an older paper given only in abbreviated form [Bab34a].

Babbage:1847:PTT

[Bab47b] Charles Babbage. On the principles of tools for turning and planing metals’. In Charles Holtzapffel, editor, Turning and Mechanical Manipulation, volume 2, pages 984–987. Holtzapffel, London, UK, 1847.

Babbage:1847:PNR

[Bab47c] Charles Babbage. The planet Neptune and the Royal Astronomical Society’s Medal. The Times [London, UK], ??(??):??, March 15, 1847. CODEN ???? ISSN ????

Babbage:1848:TPT

[Bab48] Charles Babbage. Thoughts on the principles of taxation with reference to a property tax, and its exceptions. London, UK, 1848. 24 pp.

Babbage:1849:AMM

[Bab49] Charles Babbage. General plan, no. 25, of Mr. Babbage’s gerat calculating or Analytical Engine, 1849. Lithographed at Paris, 24 by 36 inches.

Babbage:1851:EVI

[Bab51a] Charles Babbage. The Exposition of 1851; or Views of the Industry, the Science, and the Government of England. John Murray, London, UK, second edition, 1851. ???? pp. LCCN ????

Babbage:1851:LMN

[Bab51b] Charles Babbage. Laws of mechanical notation, 1851. ?? pp. Privately printed, London, UK.

Babbage:1851:NRL

[Bab51c] Charles Babbage. Note respecting lighthouses (occulting lights). ????, November 1851. ?? pp. Privately printed, London, UK. Communicated to the Trinity House, 30 November 1851.

Reprinted in the Appendix to the Report on Lighthouses presented to the Senate of the United States, February 1852. Reprinted in the Mechanics’ Magazine, and in various other periodicals and newspapers, 1852–1853. Reprinted in various parts of the Report of Commissioners appointed to examine into the state of Lighthouses, Parliamentary Paper, 1861.

Babbage:1851:SNC

[Bab51d] Charles Babbage. [Scrapbook of newspaper clippings from 1847? to 1851?]. ????, London, England [??], 1851. LCCN ???? 3646 items.

Babbage:1852:NRP

[Bab52] Charles Babbage. Note respecting the pink projections from the Sun’s disc observed during the total solar eclipse in 1851. Monthly Notices Astron. Soc., 12 (??):209–10, ???? 1852. CODEN ???? ISSN ????

Babbage:1853:SL

[Bab53] Charles Babbage. On the statistics of lighthouses. Compte Rendus des Travaux du Congrès Général de Statistique, Bruxelles, ??(??):237, September 1853. CO-DEN ???? ISSN ????

Babbage:1854:MTCa

[Bab54a] Charles Babbage. Mr. Thwaites’s cypher. Jrl. Soc. Arts, 90(??): 663–664, ???? 1854. CODEN ???? ISSN ????

Babbage:1854:MTCb

[Bab54b] Charles Babbage. Mr. Thwaites’s cypher. Jrl. Soc. Arts, 93(??): 707–708, July 1854. CODEN ???? ISSN ????

Babbage:1854:MTCc

[Bab54c] Charles Babbage. Mr. Thwaites’s cypher. Jrl. Soc. Arts, 95(??): 732–733, ???? 1854. CODEN ???? ISSN ????

Babbage:1854:MTCd

[Bab54d] Charles Babbage. Mr. Thwaites’s cypher. Jrl. Soc. Arts, 98(??): 776–777, August 1854. CODEN ???? ISSN ????

Babbage:1854:RO

[Bab54e] Charles Babbage. Report on the opthalmoscope. British and Foreign Medical Review, 14(??):425– 432, ???? 1854. CODEN ???? ISSN ????

Babbage:1855:CW

[Bab55a] Charles Babbage. Cypher writing. Jrl. Soc. Arts, 159(??):40– 41, ???? 1855. CODEN ???? ISSN ????

Babbage:1855:MLGa

[Bab55b] Charles Babbage. A method of laying the guns of a battery without exposing the men to the shot of the enemy. Illustrated London News, ??(??):??, ???? 1855. CO-DEN ???? ISSN ????

Babbage:1855:MLGb

[Bab55c] Charles Babbage. On a method of laying guns in a battery without exposing the men to the shot of the enemy. The Times [London, UK], ??(??):??, August 8, 1855. ISSN 0140-0460, 0956-1382.

Babbage:1855:SN

[Bab55d] Charles Babbage. On submarine navigation. Illustrated London News, 749(??):623–624, June 23, 1855. CODEN ???? ISSN ????

Babbage:1855:PUO

[Bab55e] Charles Babbage. On the possible use of the occulting telegraph at Sebastapol. The Times [London, UK], ??(??):??, July 16, 1855. CODEN ???? ISSN ????

Babbage:1855:MSM

[Bab55f] Charles Babbage. Sur la machine suèdoise de MM. Scheutz pour calculer les tables mathématiques. (French) [On the Swedish machine of MM. Scheutz for calculating mathematical tables]. Comptes rendus hebdomadaires, Académie des Sciences, Paris, 41(??):557– 560, ???? 1855. CODEN ???? ISSN ????

Babbage:1856:ASC

[Bab56a] Charles Babbage. Analysis of the statistics of the Clearing House during the year 1839. Journal of the Statistical Society of London, 19(1):28–48, March 1856. CO-DEN ???? ISSN ???? Also published as a separate pamphlet.

Babbage:1856:OAL

[Bab56b] Charles Babbage. Observations addressed at the last anniversary to the President and Fellows of the Royal Society. London, UK, 1856. ?? pp.

Babbage:1856:OCB

[Bab56c] Charles Babbage. Observations by Charles Babbage, on the mechanical notation of Scheutz’s Difference Engine, prepared and drawn up by his son, Major Henry Prevost Babbage, addressed to the Institution of Civil Engineers. Minutes of Proceedings [of the Institution of Civil Engineers], 15(??):??, ???? 1856.

Babbage:1856:AOC

[Bab56d] Charles Babbage. On the action of ocean currents in the formation of the strata of the Earth. Quart. Jrl. Geol. Soc., 12(??): 366–368, ???? 1856. CODEN ???? ISSN ????

Babbage:1856:ACN

[Bab56e] Charles Babbage. On the advantage of a collection of numbers, to be entitled the constants of nature and art. The Annual Report of the Board of Regents of the Smithsonian Institution, ??(??): 289–302, ???? 1856. CODEN ???? ISSN ???? Expanded version of [Bab32b].

Babbage:1856:WWO

[Bab56f] Charles Babbage. A word to the wise: Observations on peerage for life. London, UK, 1856. ?? pp. Reprint of [Bab33e].

Babbage:1856:SDE

[Bab56g] Henry P. Babbage. Scheutz’s Difference Engine and Babbage’s mechanical notation. Proc. Inst. Civil Eng., 15(??):497–514, ???? 1856. CODEN ???? ISSN ???? Also published as separate pamphlet.

Babbage:1857:TRF

[Bab57] Charles Babbage. Table of the relative frequency of occurrence of the causes of breaking of plate-glass windows. Mechanics’ Magazine, 66(??):82, January 24, 1857. CODEN ???? ISSN ????

Babbage:1859:ODV

[Bab60a] Charles Babbage. Observations on the discovery in various localities of the remains of human art mixed with the bones of extinct races of animals. Proceedings of the Royal Society of London, 10 (??):59–72, May 26, 1859–1860. CODEN PRSLAZ. ISSN 0370-1662 (print), 2053-9126 (elec-

tronic).  URL http://www.
jstor.org/stable/111607.

Babbage:1860:LDF

[Bab60b] Charles Babbage. Letter to Dr.

Farr, on the origin of the International Statistical Congresses. In Proceedings of the Fourth International Statistical Congress, page ?? ????, ????, ???? 1860.

Babbage:1860:ERS

[Bab60c] Charles Babbage. On easily recognisable signs in drawings. In Proceedings of the Fourth International Statistical Congress, page ?? ????, ????, ???? 1860.

Babbage:1864:CSN

[Bab64a] Charles Babbage. A chapter on street nuisances. ????, London, UK, third edition, 1864. 300 pp. LCCN ????

Babbage:1864:PLP

[Bab64b] Charles Babbage. Passages from the Life of a Philosopher. Longman, Green, Longman, Roberts, and Green, London, UK, 1864. xii + 496 pp. LCCN QA29.B2 A2. A Chapter on Street Nuisances was published earlier by John Murray.

Babbage:1865:TUE

[Bab65] Charles Babbage. Thoughts upon an Extension of the Franchise. Longman, Green, London, UK, 1865. ?? pp.

Babbage:1868:OPR

[Bab68] Charles Babbage. Observations on the parallel roads of Glen Roy. Quart. Jrl. Geol. Soc., 24(??): 273–277, ???? 1868. CODEN ???? ISSN ????

Babbage:1869:ELA

[Bab69] Chas Babbage. Extract from a letter addressed by Chas. Babbage, Esq., F.R.S., to Dr. Bache, of Washington, May 10, 1852. Proceedings of the Royal Society of London, 17(??):133, ???? 1869. CODEN PRSLAZ. ISSN 0370-1662 (print), 2053-9126 (electronic).

Babbage:1889:HAE

[Bab89a] Charles Babbage. History of the Analytical Engine [incomplete].

In Babbage [Bab89b], page ?? LCCN ???? Incomplete at Babbage’s death. Published posthumously.

Babbage:1889:BCB

[Bab89b] Henry Prevost Babbage, editor. Babbage’s [i.e. Charles Babbage’s] calculating engines. Being a collection of papers relating to them, etc. E. & F. N. Spon, London, UK, 1889. 342 pp. LCCN ???? Reprinted in The Charles Babbage Institute Reprint Series, volume 2.

Babbage:1899:HIM

[Bab99] Charles Babbage. How to invent machinery. ????, 1899. ???? pp. LCCN ???? Extract of [Bab32d].

Babbage:1910:BBA

[Bab10] H. P. Babbage. Babbage: Babbage’s analytical engine. Monthly Notices of the Royal Astronomical Society, 70(??):517–526, 645, 1910. CODEN MNRAA4. ISSN 0035-8711 (print), 1365-2966 (electronic). Reprinted in [Ran82b, §2.3].

Babbage:1915:TLN

[Bab15] Charles Babbage. A Table of Logarithms of the Natural Numbers from 1 to 108 000. F. & F. N. Spon, London, UK, 1915. xx + 222 pp. Reproduced in [Ano03].

Babbage:1948:WCB

[Bab48] Richard H. Babbage. The work of Charles Babbage. In Anonymous [Ano48], pages 13–22. LCCN QA75 .S96 1947.

Babbage:1953:ESQ

[Bab53] Charles Babbage. An examination of some questions connected with games of chance, volume 3 of Natural science papers. Academic Reprints, Stanford, CA, USA, 1953. 154–177 pp. LCCN QA273 .B23.

Babbage:1963:EMM

[Bab63] Charles Babbage. On the economy of machinery and manufactures. Reprints of economic classics. A. M. Kelley, New York, NY, USA, fourth (1835) edition, 1963. xxiv + 408 pp. LCCN TJ153 .B112 1835.

Babbage:1967:BT

[Bab67a] Charles Babbage. The 9. Bridgewater Treatise, volume 6 of Cass library of science classics. ????, London, UK, 1967. 273 pp. LCCN ????

Babbage:1967:CVV

[Bab67b] Charles Babbage. A Comparative View of the Various Institutions for the Assurance of Lives. Augustus M. Kelley, New York, NY, USA, 1967. xxxi + 170 pp. LCCN ???? Reprint of [Bab26a].

Babbage:1968:PLP

[Bab68] Charles Babbage. Passages from the life of a philosopher. Dawsons, 1968. ISBN 0-7129-0312-7. xii + 496 pp. LCCN QA29.B2 A2 1864a. Facsimile reprint of first edition, London, Longmans, 1864.

Babbage:1969:PLP

[Bab69]  Charles Babbage.  Passages

from the Life of a Philosopher. Reprints of Economic Classics. Augustus M. Kelley, New York, NY, USA, 1969. ISBN 0-678-00479-X. xii + 496 pp. LCCN QA29.B2 A2. Reprint of [Bab64b].

Babbage:1971:EMM

[Bab71] Charles Babbage. On the Economy of Machinery and Manufactures. Augustus M. Kelley, New York, NY, USA, 1971. xxiv + 408 + 28 + 15 pp. LCCN TJ153 .B13 1971. Reprint of 1835 edition, with the addition of Thoughts on the principles of taxation, third edition (1852).

Babbage:1982:MPC

[Bab82a] Charles Babbage. On the mathematical powers of the Calculating Engine (1837). In Randell [Ran82b], pages 19– 54. ISBN 0-387-11319-3, 3-540-11319-3. LCCN TK7885.A5 O741 1982.

Babbage:1982:BAE

[Bab82b] H. P. Babbage. Babbage’s Analytical Engine (1910). In Randell [Ran82b], pages 67– 71. ISBN 0-387-11319-3, 3-540-11319-3. LCCN TK7885.A5 O741 1982.

Babbage:1982:BCE

[Bab82c] Henry Prevost Babbage, editor.

Babbage’s calculating engines: a collection of papers, volume 2 of Charles Babbage Institute reprint series for the history of computing. Tomash Publishers, Los Angeles, CA, USA, 1982. ISBN 0-938228-04-8. xix + 342 + 22 pp. LCCN QA75 B323 1982. With a new introduction by Allan G. Bromley.

Babbage:1986:EMM

[Bab86] Charles Babbage. On the economy of machinery and manufactures. Reprints of economic classics. Augustus M. Kelley, Fairfield, NJ, USA, fourth enlarged edition, 1986. ISBN 0-678-00001-8. xxiv + 408 pp. LCCN TJ153 .B13 1986.

Babbage:1989:AEM

[Bab89a] Charles Babbage. The analytical engine and mechanical notation, volume 3 of The Works of Charles Babbage. New York University Press, New York, NY, USA, 1989. ISBN 0-8147-1113-8 (set). 6 + 253 + 2 pp. LCCN QA3 .B18 1989 vol. 3QA75. US$995.00. URL

http://www.loc.gov/catdir/
enhancements/fy0808/88032928-
b.html;   http://www.loc.gov/
catdir/enhancements/fy0808/
88032928-d.html.

Babbage:1989:CVV

[Bab89b] Charles Babbage. A comparative view of the various institutions for the assurance of lives, volume 6 of The Works of Charles Babbage. New York University Press, New York, NY, USA, 1989. 6 + xx + 129 pp. LCCN ????

Babbage:1989:DET

[Bab89c] Charles Babbage. The difference engine and table making, volume 2 of The Works of Charles Babbage. New York University Press, New York, NY, USA, 1989. ISBN 0-8147-1113-8 (set). 6 + 233 pp. LCCN QA3 .B18 1989 vol. 2QA75. US$995.00.

Babbage:1989:MP

[Bab89d] Charles Babbage. Mathematical papers, volume 1 of The Works of Charles Babbage. New York University Press, New York, NY, USA, 1989. ISBN 0-8147-1113-8 (set). 51 + 456 pp. LCCN QA3 .B18 1989 vol. 1. US$995.00.

Babbage:1989:PLP

[Bab89e] Charles Babbage, editor. Passages from the Life of a Philosopher, volume 11 of The Works of Charles Babbage. New York University Press, New York, NY, USA, 1989. ISBN 0-8147-1113-8 (set). 6 + xi + 425 pp. LCCN QA3 .B32 1989 V.11; QA29.B2A3. US$995.00.

Babbage:1989:SMP

[Bab89f] Charles Babbage. Scientific and miscellaneous papers, volume 4– 5 of The Works of Charles Babbage. New York University Press, New York, NY, USA, 1989. ISBN 0-8147-1113-8 (set). ???? pp. LCCN QA3 .B18 1989 vol. 4– 5Q113. US$995.00.

Babbage:1989:SAD

[Bab89g] Charles Babbage. Statement addressed to the Duke of Wellington respecting the calculating engine (1834). In Campbell-Kelly [CK89l], pages 2–8. ISBN 1-85196-503-3, 1-85196-005-8 (set). LCCN ????

Babbage:1991:PLP

[Bab91a] Charles Babbage. Passages from the life of a philosopher: the autobiography of Charles Babbage. Pickering and Chatto, London, UK, 1991. ISBN 1-85196-040-6 (paperback). viii + 383 pp. LCCN ???? £11.95: CIP entry ( Oct.). Edited, and with a new introduction, by Martin Campbell-Kelly.

Babbage:1991:ARB

[Bab91b] Neville F. Babbage. Autopsy report on the body of Charles Babbage. The Medical Journal of Australia, 154(??):758–759, ???? 1991. CODEN MJAUAJ. ISSN 0025-729X (print), 1326-5377 (electronic).

Babbage:1994:PLP

[Bab94] Charles Babbage, editor. Passages from the Life of a Philosopher. Rutgers University Press, New Brunswick, NJ, USA, 1994. ISBN 0-8135-2066-5 (paperback). xliv + 383 pp. LCCN QA29.B2 A3 1994. Edited, and with a new introduction, by Martin Campbell-Kelly.

Babbage:1997:HST

[Bab97]  Charles Babbage.  The his-

tory of science and technology. Series 3. Papers of Charles Babbage, (1791–1871). Part 1.

Correspondence and scientific manuscripts from the British Library, London. Listing and guide to Part 1. Adam Matthew Publications, Marlborough, Wiltshire, England, 1997. ISBN 1-85711-101-X. 89 pp. LCCN QA3 .B23 1997. URL http:

//webdoc.sub.gwdg.de/zdmdm/
mifoguide/matthew/HISTORY_
OF_SCIENCE_and_TECHNOLOGY_
S3_p1.pdf;   http://www.ampltd.
co.uk/digital_guides/history_
of_science_series_three_part_
1/Contents.aspx.

Babbage:1998:DSE

[Bab98]  Charles Babbage.  Decline

of Science in England, volume 1216 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://

uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/dosie10.
zip.

Babbage:2003:EMM

[Bab03a] Charles Babbage. From on the economy of machinery and manufactures (1832). In Freedgood [Fre03], page ?? ISBN 0-19-514872-X (paperback), 0-19-516101-7. LCCN TS155 .F4855 2003. URL

http://www.loc.gov/catdir/
enhancements/fy0613/2002025781-
d.html;   http://www.loc.gov/
catdir/enhancements/fy0613/
2002025781-t.html.

Babbage:2003:MOM

[Bab03b] Charles Babbage. On the method of observing manufactories (1832). In Freedgood [Fre03], page ?? ISBN 0-19-514872-X (paperback), 0-19-516101-7. LCCN TS155 .F4855 2003. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0613/2002025781-
                                        d.html;   http://www.loc.gov/
                                        catdir/enhancements/fy0613/
                                        2002025781-t.html.

Babbage:2006:RDS

[Bab06] Charles Babbage. Reflections on the Decline of Science in England. Project Gutenberg Literary Archive Foundation, ????, 2006. LCCN ????

Babbage:2008:CBC

[Bab08] Charles Babbage. [Charles Babbage correspondence]: Electronic Enlightenment. Electronic Enlightenment Project, Oxford, UK, 2008. ISSN 1754-1603. ???? pp. LCCN ???? URL http://

                                        hdl.library.upenn.edu/1017.
                                        12/1220513.

Babbage:2009:NBT

[Bab09] Charles Babbage. The Ninth Bridgewater Treatise. Cambridge Library Collection — Religion. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-511-70071-7. 252 (est.) pp. LCCN ????

Babbage:2010:EMMa

[Bab10a] Charles Babbage. On the Economy of Machinery and Manufactures. Cambridge Library Collection — History of Printing, Publishing and Libraries. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-511-69637-X. 344 (est.) pp. LCCN ????

Babbage:2010:EMMb

[Bab10b] Charles Babbage. On the economy of machines and manufactures. In Adamson [Ada10], page ?? ISBN 1-84788-304-4 (hardcover), 1-84788-303-6 (paperback). LCCN https://www.math.utah.edu/pub/tex/bib/adabooks.bib; TT149 .C733 2010.

Babbage:2010:BCE

[Bab10c] Henry P. Babbage, editor. Babbage’s Calculating Engines: Being a Collection of Papers Relating to them; their History and Construction, by Charles Babbage. Cambridge Library Collection — Mathematics. Cambridge University Press, Cambridge, UK, 2010. ISBN 1-108-00096-7 (paperback), 0-511-69472-5 (e-book). LCCN ????

Babbage:2011:PLP

[Bab11]  Charles Babbage.  Passages

from the Life of a Philosopher. Cambridge Library Collection - Technology. Cambridge University Press, Cambridge, UK, 2011. ISBN 1-139-10367-9. 516 pp. LCCN ????

Baily:1989:MBN

[Bai89] Francis Baily. On Mr. Babbage’s new machines for calculating and printing mathematical and astronomical tables (1823). In Campbell-Kelly [CK89k], pages 44–56. ISBN 1-85196-502-5, 1-85196-005-8 (set). LCCN ????

Barton:1971:CBB

[Bar71] H. Barton. Charles Babbage and the beginning of die casting. Machinery and Production Engineering, ??(??):624–631, ???? 1971.

Barton:1972:PDE

[Bar72] H. Barton. Pressure diecasting in the eighteen-sixties: Charles Babbage’s use of the technique on his “calculating engines”. Diecasting and Metal Moulding, ??(??):14–15, May/June 1972.

Barbrook:1996:BRE

[Bar96] Richard Barbrook. Book review: Engineering the changes: Cultural Babbage, edited by Francis Spufford and Jenny Uglow, Faber & Faber, £14,99, ISBN 0-571-17242-3. New Scientist, ?? (2021):??, March 16, 1996. CO-DEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (elec-

tronic).  URL http://www.
newscientist.com/article/mg14920215.
100-engineering-the-changes-
cultural-babbage-edited-by-
francis-spufford-and-jenny-
uglow-faber--faber-1499-isbn-
0-571-17242-3.html.

Baran:2000:NVB

[Bar00]  Nicholas Baran.  News and

views: 108-bit elliptic curve cryptographic key found; new algorithm cracks the stock market; first complete Babbage printer unveiled; XrML view to be digital rights standard; PKWare founder [Phil Katz] dies unexpectedly. Dr. Dobb’s Journal of Software Tools, 25(7):18, July 2000. CODEN DDJOEB. ISSN 1044-789X.

Barlow:1836:TMM

[BB36] Peter Barlow and Charles Babbage. A treatise on the manufactures and machinery of Great Britain. Baldwin and Cradock, London, UK, 1836. viii + 834 + 87 pp. LCCN ????

Babbage:1832:EMMa

[BBKC32] Charles Babbage, J. Bate, Charles Knight, and Richard Clay. On the economy of machinery and manufactures by Charles Babbage, Esqre A. M. Lucasian Professor of Mathematics in the University of Cambridge, and member of several academies.: Economy of manufactures. Charles Knight, Pall Mall East, London, UK, 1832. xvi + 320 + 2 pp. LCCN ???? Handwritten dedication: “To Mr Berzelius from the author”.

Babbage:1989:WCB

[BCKR89] Charles Babbage, Martin Campbell-Kelly, and Peter M. Roget, editors. The works of Charles Babbage. The Pickering masters. W. Pickering, London, UK, 1989. ISBN 1-85196-005-8 (set), 1-85196-501-7 (vol. 1), 1-85196-502-5 (vol. 2), 1-85196-503-3 (vol. 3), 1-85196-504-1 (vol. 4), 1-85196-505-X (vol. 5), 1-85196-506-8 (vol. 6), 1-85196-507-6 (vol. 7), 1-85196-508-4 (vol. 8), 1-85196-509-2 (vol. 9), 1-85196-510-6 (vol. 10), 1-85196-511-4 (vol. 11). LCCN QA37.2. 11 volumes.

Bromley:1983:RFM

[BCKS^{+}83] Allan G. Bromley, Martin Campbell-Kelly, K. W. Smillie, Eric A. Weiss, Saul Rosen, and Cipher A. Deavours. Reviews: O. I. Franksen: Mr. Babbage, the Difference Engine, and the Problem of Notation: An Account of the Origin of Recursiveness and Conditionals in Computer Programming; H. Lukoff: from Dits to Bits; I. Asimov: Asimov’s Biographical Encyclopedia of Science and Technology; J. Futrelle: Thinking Machine; R. M. Hord: The Illiac IV; C. H. Meyer and S. M. Matyas: Cryptography; T. J. Peters and R. H. Waterman: In Search of Excellence; J. W. Stokes: 70 Years of Radio Tubes and Valves; G. Welchman: The Hut Six Story; capsule reviews. Annals of the History of Computing, 5 (4):411–427, October/December 1983. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1983/pdf/a4411.pdf;   http:
//www.computer.org/annals/
an1983/a4411abs.htm.

Buchholz:1990:CQDb

[BDH90] Werner Buchholz, Donald Watts Davies, and Mark Halpern. Comments, queries, and debate: Babbage’s friend; errors in the FORTRAN issue. Annals of the History of Computing, 12(2):147, April/June 1990. CODEN AHCOE5.

ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1990/pdf/a2147a.pdf;   http:
//www.computer.org/annals/
an1990/a2147aabs.htm.

Babbage:1880:EMM

[BdL80] Charles Babbage and Charles Pierre Lefebvre de Laboulaye. Économie des machines et des manufactures d’aprés l’ouvrage anglais de Ch. Babbage. (French) [The economics of machines and manufacturing according to the English work of Ch. Babbage]. Librairie du dictionnaire des arts et manufactures, Paris, France, 1880. vi + 460 pp. LCCN

???? URL http://catalog.
hathitrust.org/api/volumes/
oclc/39709573.html.

Becher:1980:WBP

[Bec80] Harvey W. Becher. Woodhouse, Babbage, Peacock, and modern algebra. Historia Mathematica, 7(4):389–400, November 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0315086080900038.

Buchholz:1987:CQDb

[BEG^{+}87] Werner Buchholz, Ralph Erskine, Herb Grosch, William Aspray, J. A. N. Lee, and Paul M. McConnell. Comments, queries, and debate: Marian Rejewski and the chronology of Enigma; early IBM computer conferences; the computing scene 1945–1955: Another perspective; response; the ENIAC collections; response to query on hexadecimal notation. Annals of the History of Computing, 9 (3/4):369–373, July/September 1987. CODEN AHCOE5. ISSN 0164-1239. URL http://

                                        dlib.computer.org/an/books/
                                        an1987/pdf/a3369.pdf;   http:
                                        //www.computer.org/annals/
                                        an1987/a3369abs.htm. See orig-

inal query on hexadecimal notation [BRG^{+}87].

Bell:1984:ADD

[Bel84] Gwen Bell. ACM doctoral dissertation award: ACM international scholastic programming contest awards. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 16(1): 43, February 1984. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 15th SIGCSE Technical Symposium on Computer Science Education.

Bennett:1916:NPP

[Ben16] Albert A. Bennett. Note on the preceding paper: “On certain real solutions of Babbage’s functional equation” [MR1503574] by J. F. Ritt. Annals of Mathematics (2), 17(3):123, 1916. CO-DEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic).

Bernstein:1964:AEC

[Ber64] Jeremy Bernstein. The Analytical Engine: computers — past, present, and future. Random House, New York, NY, USA, 1964. xi + 113 pp. LCCN QA76 .B45 1964.

Bernstein:1981:AEC

[Ber81] Jeremy Bernstein. The Analytical Engine: computers — past, present, and future. Morrow, New York, NY, USA, revised edition, 1981. ISBN 0-688-00484-9. 131 pp. LCCN QA76 .B45 1981. Revision of [Ber64].

Berg:1982:MQMb

[Ber82] Maxine Berg. The Machinery Question and the Making of Political Economy. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-28759-6. viii + 379 pp. LCCN HB85 .B47 1982.

Berg:1992:CQD

[Ber92a]  Herman Berg.  Comments,

queries, and debate: On locating the Babbage–Quetelet letter. IEEE Annals of the History of Computing, 14(1):7–9, January– March 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

dlib.computer.org/an/books/
an1992/pdf/a1006.pdf.

Berg:1992:LBQ

[Ber92b]  Herman Berg.  On locating

the Babbage–Quetelet letter. comment on: “Babbage’s letter to Quetelet, May 1835” [Ann. Hist. Comput. 5 (1983), no. 3, 263–267; MR0714019 (85a:01041)] by A. W. Van Sinderen. IEEE Annals of the History of Computing, 14(1):7– 9, 1992. CODEN IAHCEX.

ISSN 1058-6180 (print), 1934-1547 (electronic).

Beyer:1979:BRB

[Bey79a] W. A. Beyer. Book review: The Mathematical Work of Charles Babbage by J. M. Dubbey. American Mathematical Monthly, 86 (1):66–67, January 1979. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/2320321.

Beyer:1979:RBM

[Bey79b] W. A. Beyer. Reviews:

booktitleThe Mathematical Work of Charles Babbage, by J. M. Dubbey. American Mathematical Monthly, 86(1):66–67, January 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Beyer:1979:RMW

[Bey79c] W. A. Beyer. Reviews: The Mathematical Work of Charles Babbage, by J. M. Dubbey. American Mathematical Monthly, 86(1):66–67, January 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Babbage:1813:P

[BH13] Charles Babbage and John F. W. Herschel. The preface. Memoirs of the Analytical Society [Cambridge], page ??, 1813. Reprinted in [Sch81].

Babbage:1825:ARM

[BH25] C. Babbage and J. F. W. Herschel. Account of the repetition of M. Arago’s experiments on the magnetism manifested by various substances during the act of rotation. Philosophical Transactions of the Royal Society of London, 115(4):467–496, ???? 1825. CODEN PTRSAV. ISSN 0370-2316 (print), 2053-9207 (elec-

tronic).  URL http://www.
jstor.org/stable/107753.

Babbage:1830:ARM

[BH30] C. Babbage and J. F. W. Herschel. Account of the repetition of M. Arago’s experiments on the magnetism manifested by various substances during the act of rotation. [abstract]. Abstracts of the Papers Printed in the Philosophical Transactions of the Royal Society of London, 2(1):249–250, ???? 1830.

Buxton:1988:MLL

[BH88] H. W. (Harry Wilmot) Buxton and Anthony Hyman, editors. Memoir of the life and labours of the late Charles Babbage Esq., F.R.S., volume 13 of Charles Babbage Institute reprint series for the history of computing. MIT Press, Cambridge, MA, USA, 1988. ISBN 0-585-36067-7 (e-book), 0-262-02269-9. xix + 401 pp. LCCN QA29.B2 B89 1988. Edited and with an introduction by Anthony Hyman; From Computing Reviews: “Charles Babbage was a nineteenth-century polymath who initially studied mathematics at Cambridge and subsequently pursued interests in fields as diverse as philosophy, political economy, and mechanical computing. While attempting to realize a machine for calculating and printing error-free mathematical tables, he visited factories to learn metalworking techniques, invented a system of mechanical notation to represent the states of the machine, and devised a method for correct typesetting. As the editor of this series notes, the scope of Babbage’s work makes him a difficult subject to understand in our modern age of specialized scholarship.”.

Babbage:1989:SRS

[BH89] Charles Babbage and Anthony Hyman. Science and reform: selected works of Charles Babbage. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-34311-9. vii + 356 + 9 pp. LCCN QA3 .B2 1989; QA3 .B33 1989;

QA37.2. URL http://www.
loc.gov/catdir/description/
cam023/88010294.html;   http:
//www.loc.gov/catdir/toc/cam028/
88010294.html.

Buchholz:1989:CQDb

[BHB^{+}89] Werner Buchholz, Anthony Hyman, Ralf Bülow, Bernard A. Galler, and Jack Minker. Comments, queries, and debate: Babbage Studies; An Ecological Computing Machine; The University of Michigan’s B-5000 Decision; ACM 20th Anniversary Meeting. Annals of the History of Computing, 11(3):225–229, July/September 1989. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1989/pdf/a3225.pdf;   http:
//www.computer.org/annals/
an1989/a3225abs.htm.

Bessenyei:2012:FEF

[BHK12] Mihály Bessenyei, Gábor Horváth, and Csaba G. Kézi. Functional equations on finite groups of substitutions. Expositiones Mathematicae, 30(3):283–294, ???? 2012. CODEN ???? ISSN

????  URL http://www.
sciencedirect.com/science/
article/pii/S0723086912000266.

Babbage:1820:EDI

[BHP20] Charles Babbage, John Herschel, and George Peacock. Examples to the Differential and Integral Calculus. J. Deighton, Cambridge, UK, 1820. ???? pp. Babbage wrote Part 111, Examples of the Solutions of Functional Equations.

Buchholz:1990:CQDa

[BHW^{+}90] Werner Buchholz, R. Anthony Hyman, Maurice V. Wilkes, Richard E. Smith, George T. Jacobi, Alice R. Burks, Arthur W. Burks, I. Bernard Cohen, I. O. Grattan-Guinness, S. I. Gass, and Mike Woodger. Comments, queries, and debate: Whiggism in the History of Science and the Study of the Life and Work of Charles Babbage; Whirlwind and Microprogramming; More on Wilkes, Whirlwind, and Microprogramming; Pioneer Biographies; The Atanasoff Story — A Response; Notes on Babbage, Aiken, and Bowditch; Joseph Fourier’s Anticipation of Linear Programming; What Does BNF Stand for? Annals of the History of Computing, 12(1):62–72, January/ March 1990. CODEN AHCOE5. ISSN 0164-1239. URL http://

                                        dlib.computer.org/an/books/
                                        an1990/pdf/a1062.pdf;   http:
                                        //www.computer.org/annals/
                                        an1990/a1062abs.htm.

Bizup:2003:MCV

[Biz03] Joseph Bizup. Manufacturing culture: vindications of early Victorian industry. Victorian literature and culture series. University of Virginia Press, Charlottesville, VA, USA, 2003. ISBN 0-8139-2246-1 (hardcover). xii + 229 pp. LCCN DA533 .B57 2003.

Bessenyei:2011:FEG

[BK11] Mihály Bessenyei and Csaba G. Kézi. Functional equations and group substitutions. Linear Algebra and its Applications, 434 (6):1525–1531, March 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0024379510006051.

Babbage:1953:AEL

[BL53] Charles Babbage and Ada Augusta Lovelace. Appendix 2: Extracts from the Lovelace Papers. In Bowden [Bow53b], page 409. LCCN QA76.5 .B66.

URL https://archive.org/
details/FasterThanThought.

Quotes from four letters between Charles Babbage and the Countess of Lovelace.

Blaug:1991:WWD

[Bla91] Mark Blaug, editor. William Whewell (1794–1866), Dionysius Lardner (1793–1859), Charles Babbage (1792–1871), volume 19 of Pioneers in economics; An Elgar reference collection. E. Elgar, Aldershot, UK, 1991. ISBN 1-85278-481-4. x + 273 pp. LCCN ????

Black:2002:IHS

[Bla02] Edwin Black. IBM and the Holocaust: the strategic alliance between Nazi Germany and America’s most powerful corporation. Three Rivers Press, New York, NY, USA, 2002. ISBN 0-609-80899-0 (paperback). 551 pp. LCCN HD9696.2.U64 I253 2002.

Babbage:1823:SOC

[BM23] Charles Babbage and Francis Maseres, editors. Scriptores Optici; or, a Collection of Tracts Relating to Optics. Baldwin, Craddock & Joy, London, UK, 1823. ?? pp. Preface by C. Babbage.

Babbage:1970:RDS

[BM70] Charles Babbage and Gerard Moll. Reflections on the decline of science in England, and on some of its causes,: On the alleged decline of science in England. Reprints of economic classics. Augustus M. Kelley, New York, NY, USA, 1970. ISBN 0-678-00645-8. xvi + 228 + 33 pp. LCCN Q127.G4 B3 1970.

Booth:1962:BRB

[Boo62] A. D. Booth. Book review: Charles Babbage and his Calculating Engines: Selected Writings by Charles Babbage and Others, by Philip Morrison, Emily Morrison. Journal of the Royal Statistical Society. Series A (General), 125(3):491–493, ???? 1962. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.
                                        org/stable/2982420.

Bowden:1953:BHC

[Bow53a] Baron Bertram Vivian Bowden. A brief history of computation. In Faster Than Thought: a Symposium on Digital Computing Machines [Bow53b], pages 3–31. LCCN QA76.5 .B66.

                                        URL https://archive.org/
                                        details/FasterThanThought.

With a foreword by the Right Honourable the Earl of Halsbury.

Bowden:1953:FTT

[Bow53b] Baron Bertram Vivian Bowden, editor. Faster Than Thought: a Symposium on Digital Computing Machines. Sir Isaac Pitman and Sons, Ltd., London, UK, 1953. LCCN QA76.5 .B66.

                                        URL https://archive.org/
                                        details/FasterThanThought.

With a foreword by the Right Honourable the Earl of Halsbury.

Bowles:1993:IST

[Bow93] Mark D. Bowles. The Idealistic Stimulus of Technology: Charles Babbage and the Drive for Reform. Case Western Reserve University, Cleveland, OH, USA, 1993. ISBN ???? ???? pp. LCCN ???? Cited in [Swa01b, page 323] but with ‘Cape’ instead of ‘Case’; is this a thesis? Not in Dissertation Abstracts, and not among the 50 or so reports by Bowles in the CWU library catalog.

Bowen:2015:BRI

[Bow15] Jonathan P. Bowen. Book review: It began with Babbage: the genesis of computer science, by Subrata Dasgupta. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(3): 263–265, 2015. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1080/17498430.2015.
1036336.

Babbage:1989:RDS

[BR89] Charles Babbage and Peter Mark Roget. Reflections on the decline of science in England and on some of its causes: with a new appendix of correspondence by Charles Babbage and Peter M. Roget from the Philosophical Magazine, volume 7 of The Works of Charles Babbage. New York University Press, New York, NY, USA, 1989. xvi + 133 pp. LCCN ????

Bromley:1993:DE

[BR93] Allan G. Bromley and Anthony Ralston. Difference Engine. In Ralston and Reilly, Jr. [RR93], pages 431–432. ISBN 0-442-27679-6. LCCN QA76.15 .E48 1993.

Bramwell:1991:BB

[Bra91] Bob Bramwell. Back to Babbage. Technical Report 91/422/06, University of Calgary, Calgary, AB, Canada, March 1991.

Brewster:1975:DDS

[Bre75] David Brewster, editor. Debates on the decline of science. History, philosophy and sociology of science. Arno Press, New York, NY, USA, 1975. ISBN 0-405-06632-5. 220 (est.) pp. LCCN Q127.G4 D4.

Buchholz:1987:CQDa

[BRG^{+}87] Werner Buchholz, C. J. D. Roberts, Fred Gruenberger, Bruce H. Bruemmer, Jakob Nielson, and Paul Ceruzzi. Comments, queries, and debate: Babbage’s Difference Engine No. 1 and the production of sine tables; query on hexadecimal notation; Burroughs B 5000 oral history; DASK, the first Danish computer; Smithsonian exhibition on computers and flight. Annals of the History of Computing, 9(2):210–213, April/ June 1987. CODEN AHCOE5. ISSN 0164-1239. URL http://

                                        dlib.computer.org/an/books/
                                        an1987/pdf/a2210.pdf;   http:
                                        //www.computer.org/annals/

an1987/a2210abs.htm. See re-

sponse on hexadecimal query [BEG^{+}87].

Brock:1966:SAB

[Bro66] W. H. Brock. The selection of the authors of the Bridgewater Treatises. Notes and Records of the Royal Society of London, 21(2): 162–179, December 1966. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Bromley:1982:CBA

[Bro82]  Allan G. Bromley.  Charles

Babbage’s Analytical Engine, 1838. Annals of the History of Computing, 4(3):196– 217, July/September 1982. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1982/pdf/a3196.pdf;   http:
//www.computer.org/annals/
an1982/a3196abs.htm.

Bromley:1985:BGE

[Bro85] Allan G. Bromley. Babbage general: the evolution of Charles Babbage’s calculating engines. Technical report 272, Basser Department of Computer Science, University of Sydney, Sydney, NSW, Australia, 1985. ISBN 0-949269-91-3. 30 + 20 pp.

Brooke:1986:BRB

[Bro86] N. Michael Brooke. Book review: Mr. Babbage’s secret: the tale of a Cypher — and APL: O. I. Franksen. Strandbergs Forlag, Denmark (1984). 320 pp. Dkr.

320.00. ISBN: 87-872-0086-4. Information Processing and Management, 22(1):67–68, ???? 1986. CODEN IPMADK. ISSN 0306-4573 (print), 1873-5371 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/030645738690018X.

Bromley:1987:EBC

[Bro87] Allan G. Bromley. The evolution of Babbage’s calculating engines. Annals of the History of Computing, 9(2):113–136, April/ June 1987. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1987/pdf/a2113.pdf;   http:
//www.computer.org/annals/
an1987/a2113abs.htm.

Bromley:1990:DAE

[Bro90] Allan G. Bromley. Difference and Analytical Engines. In Aspray et al. [ABCK^{+}90], page ?? ISBN 0-8138-0047-1. LCCN QA76 .C5834 1990. URL http:

//www.computerhistory.org/
collections/DocumentArchive/
Documents/Books/Computing\
%20Before\%20Computers/CBC.
html.

Bromley:1991:BPS

[Bro91] Allan G. Bromley. The Babbage papers in the Science Museum Library: a cross-referenced list. Science Museum, London, UK, 1991. ISBN 0-901805-44-0 (paperback). 294 (est.) pp. LCCN QA75 .B76 1991; QA75 .S35 1991.

Brown:1992:CBM

[Bro92] Donald H. Brown. Charles Babbage: the man and his machines. Totnes Museum Society, Totnes, UK, 1992. 42 pp. LCCN ????

Bromley:1998:CBA

[Bro98]  Allan G. Bromley.  Charles

Babbage’s Analytical Engine, 1838. IEEE Annals of the History of Computing, 20(4): 29–45, October–December 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:/

/ieeexplore.ieee.org/iel4/
85/15706/00728228.pdf.

Bromley:2000:BAE

[Bro00] Allan G. Bromley. Babbage’s Analytical Engine Plans 28 and 28a. The programmer’s interface. IEEE Annals of the History of Computing, 22(4): 4–19, October–December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

dlib.computer.org/an/books/
an2000/pdf/a4005.pdf;   http:
//ieeexplore.ieee.org/iel5/
85/19196/00887986.pdf;   http:
//www.computer.org/annals/
an2000/a4005abs.htm. With

an introduction by Maurice V. Wilkes.

Belanger:2005:SVS

[BS05] Jay Belanger and Dorothy Stein.

Shadowy vision: spanners in the mechanization of mathematics. Historia Mathematica, 32 (1):76–93, February 2005. CO-DEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0315086004000217.

Bulferetti:1968:ACB

[Bul68]  Luigi Bulferetti.  Un amico

di Charles Babbage, Fortunato Prandi, volume 30:2 of Memorie dell’Istituto lombardo-accademia di scienze e lettere, Classe di lettere, scienze morali e storiche. Istituto lombardo di scienze e lettere, Milano, Italy, 1968. ISBN 99-0258041-6. 86–164 pp. LCCN AS222.I85 A28 v.30.

Bulferetti:1971:CIC

[Bul71] Luigi Bulferetti. Les correspondants italiens de Charles Babbage. (French) [The Italian correspondents of Charles Babbage]. In Actes du XIIe Congrès International d’Histoire des Sciences (Paris, 1968), Tome XI: Sciences et sociétés: relationsinfluences-écoles, pages 29–38. Librairie Sci. Tech. Blanchard, Paris, 1971.

Bulow:1988:GFD

[Bül88] Ralf Bülow. Glory and Failure: The Difference Engines of Johann Muller, Charles Babbage and Georg and Edvard Scheutz by Michael Lindgren (review). Technology and Culture, 29(4): 956–958, October 1988. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL http://www.jstor.org/

stable/3105078;   https://
muse.jhu.edu/pub/1/article/
889191/pdf.

Bullock:2008:CBE

[Bul08] Seth Bullock. Charles Babbage and the emergence of automated reason. In Husbands et al. [HHW08], pages 19–40. ISBN 0-262-08377-9 (hardcover). LCCN Q335 .M3956 2008.

Buttmann:1974:ST

[But74] Günther Buttmann, editor. The Shadow of the Telescope: a Biography of John Herschel. Lutterworth Press, Guildford, UK, 1974. ISBN 0-7188-2087-8. xiv + 219 pp. LCCN QB36.H59 B813 1974. Page 18 contains a letter from John Herschel to Charles Babbage.

Babbage:2007:PDV

[BV07] Charles Babbage and Andrea Villa, editors. Passaggi dalla vita di uno scienziato: autobiografia dell’inventore del computer. (Italian) [Passages from the Life of a Philosopher: autobiography of the inventor of the computer]. Frontiere / [UTET libreria]. UTET libreria, Torino, Italia, 2007. ISBN 88-02-07715-0. xvi + 414 pp. LCCN ???? With an introduction by Vittorio Marchis.

Buchholz:2000:WBC

[BW00] W. Buchholz and M. Wilkes.

Was Babbage caught in the act? IEEE Annals of the History of Computing, 22(4):

69–70, October–December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
85/19196/00887996.pdf.

Buchholz:1988:CQDc

[BWV^{+}88] Werner Buchholz, Maurice V. Wilkes, Alfred W. Van Sinderen, C. J. Fern, Jr., and W. L. van der Poel. Comments, queries, and debate: Babbage and the Colossus; Babbage and Bowditch; Two Early European Computers; Early Dutch Computer. Annals of the History of Computing, 10(3):218–221, July/September 1988. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1988/pdf/a3218.pdf;   http:
//www.computer.org/annals/
an1988/a3218abs.htm.

Byrne:1838:NIS

[Byr38] Oliver Byrne. New and improved system of logarithms: with a table of the logarithms of the natural numbers, from 1 to 100,000, and by differences, from 1 to 10,000,000; the calculations have been recently made, and afterwards compared with the tables of Callett, Hutton, Vega, Briggs, Gardiner, Taylor, and Babbage; to which are added, an account of their origin, nature and use; and their application to several subjects to which they appertain; also, and appendix, containing tables of trigonometrical formulae, (the notation entirely new and simple). W. Day, London, UK, 1838. xii + 106 pp. LCCN ????

Byrd:2004:SS

[Byr04] Max Byrd. Shooting the sun.

Bantam Books, New York, NY, USA, 2004. ISBN 0-553-80208-

9. 306 pp. LCCN PS3552.Y675

S47 2004. URL http://www.
loc.gov/catdir/description/
random0414/2003052403.html.

C:1988:DE

[C.88]  E. C.  A different engine?

Scientific American, 259(6): 111, December 1988. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/
scientificamerican/journal/
v259/n6/pdf/scientificamerican1288-
111.pdf.

Cockshott:2009:CE

[C^{+}09] W. Paul Cockshott et al., editors. Classical econophysics, volume 12 of Routledge advances in experimental and computable economics. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2009. ISBN 0-415-47848-0 (hardcover), 0-203-87754-3 (e-book). xvi + 364 pp. LCCN HB131 .C53 2009; HG131.

Calvert:1968:IGL

[Cal68] H. R. Calvert. Irascible Genius:

a Life of Charles Babbage, Inventor by Maboth Moseley (review). Technology and Culture, 9 (1):122–123, January 1968. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL http://www.jstor.org/
stable/3102064;   https://
muse.jhu.edu/pub/1/article/
894364/pdf.

Cannon:1966:BRB

[Can66] Walter F. Cannon. Book review: Irascible Genius: A Life of Charles Babbage, Inventor, by Maboth Moseley. Victorian Studies, 10(2):225–226, December 1966. ISSN 0042-5222 (print), 1527-2052 (elec-

tronic).  URL http://www.
jstor.org/stable/3825202.

Cardwell:1969:BRE

[Car69] D. S. L. Cardwell. Book review: Eighteenth and Nineteenth Centuries: The Ninth Bridgewater Treatise: A Fragment. By Charles Babbage. Second edition. 1837. London: F. Cass. 1967. 90s. British Journal for the History of Science, 4(4): 421–422, December 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4025254.

Catanese:1981:BCC

[Cat81] F. Catanese. Babbage’s conjecture, contact of surfaces, symmetric determinantal varieties and applications. Inventiones Mathematicae, 63(3):433– 465, 1981. CODEN INVMBH. ISSN 0020-9910 (print), 1432-1297 (electronic).

Copeland:2017:TG

[CBSW17] B. Jack Copeland, Jonathan Bowen, Mark Sprevak, and Robin Wilson, editors. The Turing guide. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2017. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). xv + 546 pp. LCCN QA29.T8 C67 2017.

Ceruzzi:1983:RPD

[Cer83] Paul E. Ceruzzi. Reckoners:

the prehistory of the digital computer, from relays to the stored program concept, 1935–1945, volume 1 of Contributions to the study of computer science. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1983. ISBN 0-313-23382-9 (hardcover). ISSN 0734-757X. xii + 181 pp. LCCN QA76.5 .C4164 1983.

Clarke:2002:EIR

[CH02] Bruce Clarke and Linda Dalrymple Henderson, editors. From energy to information: representation in science and technology, art, and literature. Writing science. Stanford University Press, Stanford, CA, USA, 2002. ISBN 0-8047-4176-X, 0-8047-4210-3 (paperback). xviii + 440 pp. LCCN T15 .F78 2002. URL

http://www.loc.gov/catdir/
enhancements/fy0710/2002070594-
b.html;   http://www.loc.gov/
catdir/enhancements/fy0710/
2002070594-d.html;   http:/
/www.loc.gov/catdir/toc/fy033/
2002070594.html.

Chajoth:1933:FB

[Cha33] Z. Chajoth. Über eine Funktionalgleichung von Babbage. (German) [On a functional equation of Babbage]. Monatshefte für Mathematik und Physik, 40(1): 235–240, 1933. CODEN MM-PHA8. ISSN 0026-9255 (print), 1436-5081 (electronic).

Champion:2001:CB

[Cha01] Neil Champion. Charles Babbage. Groundbreakers. Heinemann Library, Oxford, UK, 2001. ISBN 0-431-10460-3 (paperback), 0-431-10448-4 (cased), 1-57572-367-0. 48 pp. LCCN QA29.B2 C48 2000.

Chaline:2012:FMC

[Cha12] Eric Chaline. Fifty machines that changed the course of history. Firefly Books, Buffalo, NY, USA, 2012. ISBN 1-77085-090-2. 224 pp. LCCN T19 .C47 2012.

Campbell-Kelly:1988:CBT

[CK88] Martin Campbell-Kelly. Charles Babbage’s table of logarithms (1827). Annals of the History of Computing, 10(3): 159–169, July/September 1988. CODEN AHCOE5. ISSN 0164-1239. URL http://

                                        dlib.computer.org/an/books/
                                        an1988/pdf/a3159.pdf;   http:
                                        //www.computer.org/annals/
                                        an1988/a3159abs.htm.

Campbell-Kelly:1989:BRA

[CK89a] Martin Campbell-Kelly. Book review: Anthony Hyman, Editor, Science and Reform. Selected Works of Charles Babbage (1989) Cambridge University Press, 356 pp. Endeavour, 13(4):190, ???? 1989. CODEN ENDEAS. ISSN 0141-3058, 0013-7162, 0160-9327. URL http://www.

sciencedirect.com/science/
article/pii/S0160932789800328.

Campbell-Kelly:1989:CBT

[CK89b] Martin Campbell-Kelly. Charles Babbage’s table of logarithms (1823). In The works of Charles Babbage, Vol. 10: The Exposition of 1851 [CK89i], pages 159– 169. ISBN 1-85196-510-6, 1-85196-005-8 (set). LCCN ????

Campbell-Kelly:1989:EMM

[CK89c] Martin Campbell-Kelly, editor.

The Economy of Machinery and Manufactures, volume 8 of The Works of Charles Babbage. New York University Press, Washington Square, New York, NY, USA, 1989. ISBN 0-8147-1113-8 (set). 8 + xxviii + 280 pp. LCCN QA3 .B18 1989 vol. 8; TJ153. US$995.00.

Campbell-Kelly:1989:E

[CK89d] Martin Campbell-Kelly, editor.

The Exposition of 1851, volume 10 of The Works of Charles Babbage. New York University Press, New York, NY, USA, 1989. ISBN 0-8147-1113-8 (set). 5 + xi + 173 pp. LCCN QA3 .B18 1989 vol. 10; T690.B1. US$995.00.

Campbell-Kelly:1989:WCB

[CK89e] Martin Campbell-Kelly, editor.

The works of Charles Babbage.

William Pickering, London, UK, 1989. ISBN 1-85196-005-8, 1-85196-501-7 v.1, 1-85196-502-5 v.2, 1-85196-503-3 v.3, 1-85196-504-1 v.4, 1-85196-505-X v.5, 1-85196-506-8 v.6, 1-85196-507-6 v.7, 1-85196-508-4 v.8, 1-85196-509-2 v.9, 1-85196-510-6 v.10, 1-85196-511-4 v.11. 11 v. pp. LCCN QA37.2.

Campbell-Kelly:1989:WCBa

[CK89f] Martin Campbell-Kelly, editor. The works of Charles Babbage. The Pickering Masters. William Pickering, London, UK, 1989. ISBN 1-85196-005-8 (set). ???? pp. LCCN QA3 .B18 1989g.

Campbell-Kelly:1989:WCBb

[CK89g] Martin Campbell-Kelly, editor. The works of Charles Babbage. New York University Press, New York, NY, USA, 1989. ISBN 0-8147-1113-8 (set). ???? pp. LCCN QA3 .B18 1989; QA 3 .B18 1989.

Campbell-Kelly:1989:WCB-1

[CK89h] Martin Campbell-Kelly, editor. The works of Charles Babbage, Vol. 1, Mathematical papers. William Pickering, London, UK, 1989. ISBN 1-85196-501-7, 1-85196-005-8 (set). 456 pp. LCCN ????

Campbell-Kelly:1989:WCB-10

[CK89i] Martin Campbell-Kelly, editor. The works of Charles Babbage, Vol. 10: The Exposition of 1851. William Pickering, London, UK, 1989. ISBN 1-85196-510-6, 1-85196-005-8 (set). 173 pp. LCCN ????

Campbell-Kelly:1989:WCB-11

[CK89j] Martin Campbell-Kelly, editor.

The works of Charles Babbage, Vol. 11: Passages from the life of a philosopher. William Pickering, London, UK, 1989. ISBN 1-85196-511-4, 1-85196-005-8 (set). 425 pp. LCCN ????

Campbell-Kelly:1989:WCB-2

[CK89k] Martin Campbell-Kelly, editor.

The works of Charles Babbage, Vol. 2, The Difference Engine and table making. William Pickering, London, UK, 1989. ISBN 1-85196-502-5, 1-85196-005-8 (set). 233 pp. LCCN ????

Campbell-Kelly:1989:WCB-3

[CK89l] Martin Campbell-Kelly, editor.

The works of Charles Babbage, Vol. 3, The analytical engine and mechanical notation. William Pickering, London, UK, 1989. ISBN 1-85196-503-3, 1-85196-005-8 (set). 253 pp. LCCN ????

Campbell-Kelly:1989:WCB-4

[CK89m] Martin Campbell-Kelly, editor.

The works of Charles Babbage, Vol. 4, Scientific and miscellaneous papers, I. William Pickering, London, UK, 1989. ISBN 1-85196-504-1, 1-85196-005-8 (set). 217 pp. LCCN ????

Campbell-Kelly:1989:WCB-5

[CK89n] Martin Campbell-Kelly, editor.

The works of Charles Babbage, Vol. 5: Scientific and miscellaneous papers, II. William Pickering, London, UK, 1989. ISBN 1-85196-505-X, 1-85196-005-8 (set). 192 pp. LCCN ????

Campbell-Kelly:1989:WCB-6a

[CK89o] Martin Campbell-Kelly, editor. The works of Charles Babbage. Vol. 6. A comparative view of the various institutions for the assurance of lives, volume 6 of The Works of Charles Babbage. W. Pickering, London, UK, 1989. ISBN 1-85196-506-8. xx + 129 pp. LCCN QA3 .B18 1989 vol. 6HG9057.

Campbell-Kelly:1989:WCB-6b

[CK89p] Martin Campbell-Kelly, editor. The works of Charles Babbage. Vol. 6. A comparative view of the various institutions for the assurance of lives, volume 6 of The Works of Charles Babbage. New York University Press, New York, NY, USA, 1989. ISBN 0-8147-1113-8 (set). 6 + xx + 129 pp. LCCN QA3 .B18 1989 vol. 6HG9057. US$995.00.

Campbell-Kelly:1989:WCB-7a

[CK89q] Martin Campbell-Kelly, editor. The Works of Charles Babbage. Vol. 7. Reflections on the decline of science in England and on some of its causes: with a new appendix of correspondence by Charles Babbage and Peter M. Roget from the Philosophical Magazine: Decline of science, volume 7 of The Works of Charles Babbage; The Pickering masters. W. Pickering, London, UK, 1989. ISBN 1-85196-507-6. xvi + 133 pp. LCCN QA3 .B32 1989 V.7. With a new appendix of correspondence by Charles Babbage and Peter M. Roget from the Philosophical Magazine.

Campbell-Kelly:1989:WCB-7b

[CK89r] Martin Campbell-Kelly, editor.

The Works of Charles Babbage. Vol. 7. Reflections on the decline of science in England and on some of its causes: with a new appendix of correspondence by Charles Babbage and Peter M. Roget from the Philosophical Magazine: Decline of science, volume 7 of The Works of Charles Babbage. New York University Press, New York, NY, USA, 1989. ISBN 0-8147-1113-8 (set). 5 + xvi + 133 pp. LCCN QA3 .B18 1989 vol. 7Q127.G4. US$995.00.

Campbell-Kelly:1989:WCB-8

[CK89s] Martin Campbell-Kelly, editor.

The works of Charles Babbage, Vol. 8: The economy of machinery and manufactures. William Pickering, London, UK, 1989. ISBN 1-85196-508-4, 1-85196-005-8 (set). 280 pp. LCCN ????

Campbell-Kelly:1989:WCB-9

[CK89t] Martin Campbell-Kelly, editor.

The works of Charles Babbage. Vol. 9. The Ninth Bridgewater treatise: a fragment, volume 9 of The Works of Charles Babbage. New York University Press, New York, NY, USA, second edition, 1989. ISBN 0-8147-1113-8 (set). 7 + xx + 118 pp. LCCN QA3 .B32 1989 V.9. US$995.00.

Campbell-Kelly:1992:CBI

[CK92] Martin Campbell-Kelly. The Charles Babbage Institute Reprint Series for the history of computing. IEEE Annals of the History of Computing, 14(4): 44–48, October–December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        dlib.computer.org/an/books/
                                        an1992/pdf/a4008.pdf.

Campbell-Kelly:1994:CBA

[CK94] Martin Campbell-Kelly. Charles Babbage and the Assurance of lives. IEEE Annals of the History of Computing, 16(3):5–14, Fall 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        dlib.computer.org/an/books/
                                        an1994/pdf/a3005.pdf;   http:
                                        //www.computer.org/annals/
                                        an1994/a3005abs.htm.

Campbell-Kelly:2000:CBE

[CK00] Martin Campbell-Kelly. Charles Babbage eulogy. IEEE Annals of the History of Computing, 22(4):20–21, October–December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        dlib.computer.org/an/books/
                                        an2000/pdf/a4020.pdf;   http:
                                        //ieeexplore.ieee.org/iel5/
                                        85/19196/00887987.pdf.

Campbell-Kelly:2001:CBE

[CK01a] M. Campbell-Kelly. Charles Babbage and the engines of perfection [reviews]. IEEE Annals of the History of Computing, 23(3):78, July 2001. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://computer.org/
annals/an2001/a3078abs.htm;
 http://dlib.computer.org/
an/books/an2001/pdf/a3078.
pdf.

Campbell-Kelly:2001:BRB

[CK01b] Martin Campbell-Kelly. Book review: Bits and wheels: Doron Swade, The cogwheel brain: Charles Babbage and the quest to build the first computer. Little, Brown & Co., London, 2000. Pp. x + 342, £14.99 (hardback). ISBN 0-316-64847-7. Notes and Records of the Royal Society of London, 55(2):339–340, May 22, 2001. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/532109.

Campbell-Kelly:2001:CBI

[CK01c] Martin Campbell-Kelly. The Charles Babbage Institute reprint series for the history of computing. IEEE Annals of the History of Computing, 23(4): 44–48, October 2001. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL http://computer.org/
annals/an2001/a4044abs.htm;
 http://dlib.computer.org/

                                        an/books/an2001/pdf/a4044.
                                        pdf.

Campbell-Kelly:2003:HMT

[CKCFR03] Martin Campbell-Kelly, Mary Croarken, Raymond Flood, and Eleanor Robson, editors. The History of Mathematical Tables: From Sumer to Spreadsheets. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2003. ISBN 0-19-850841-7 (print). 372 pp. LCCN ????

Clayton:2003:CDC

[Cla03] Jay Clayton. Charles Dickens in cyberspace: the afterlife of the nineteenth century in postmodern culture. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2003. ISBN 0-19-516051-7. x + 270 pp. LCCN PR451 .C58 2003.

Cleveland:2009:CEH

[Cle09] Cutler J. Cleveland. Concise Encyclopedia of the History of Energy. Elsevier, Amsterdam, The Netherlands, 2009. ISBN 0-12-375117-9 (print), 0-12-375118-7 (e-book). ix + 354 pp. LCCN TJ163.16 .C657 2009.

Collier:1998:CBE

[CM98] Bruce Collier and James Henry MacLachlan. Charles Babbage and the engines of perfection. Oxford portraits in science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1998. ISBN 0-19-508997-9. 123 pp. LCCN QA29.B2 C65 1998. URL

http://www.loc.gov/catdir/
enhancements/fy0603/98017054-
d.html;   http://www.zentralblatt-
math.org/zmath/en/search/?
an=1013.01006.

Cassata:2011:TPE

[CM11] Francesco Cassata and Roberto Marchionatti. A transdisciplinary perspective on economic complexity. Marshall’s problem revisited. Journal of Economic Behavior & Organization, 80(1):122–136, September 2011. CODEN ???? ISSN

????  URL http://www.
sciencedirect.com/science/
article/pii/S0167268111000886.

Cohen:1988:BAN

[Coh88] I. Bernard Cohen. Babbage and Aiken: with notes on Henry Babbage’s gift to Harvard, and to other institutions, of a portion of his father’s Difference Engine. Annals of the History of Computing, 10(3):171–193, July/September 1988. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1988/pdf/a3171.pdf;   http:
//www.computer.org/annals/
an1988/a3171abs.htm.

Collier:1971:LEC

[Col71] Bruce Collier. The little engines that could’ve: the calculating machines of Charles Babbage. Ph.D. thesis, Harvard University, Cambridge, MA, USA, 1971. 140 pp. URL

http://hollis.harvard.edu/
?itemid=%7Clibrary/m/aleph%

                                        7C003811101;   http://search.
                                        proquest.com/docview/302590382.

Collier:1990:LEC

[Col90] Bruce Collier. The little engines that could’ve: the calculating machines of Charles Babbage. Harvard dissertations in the history of science. Garland Publishing, New York, NY, USA, 1990. ISBN 0-8240-0043-9. ix + 319 pp. LCCN QA75 .C635 1991.

Coldwell:1998:VDC

[Col98] Roger A. Coldwell. Viewpoint: did Chuck Babbage predict software piracy? Communications of the ACM, 41(8):25–27, August 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org:80/pubs/citations/
                                        journals/cacm/1998-41-8/p25-
                                        coldwelll/.

Colclasure:2006:MBM

[Col06] K. Colclasure. Mr. Babbage, meet Mr. Lego. IEEE Spectrum, 43(4):26–27, April 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comrie:1931:MBM

[Com31] L. J. Comrie. Modern Babbage machines. Bulletin of the Office Machinery Users’ Association Limited, ??(??):29–??, ???? 1931.

Comrie:1936:IIS

[Com36] L. J. Comrie. Inverse interpolation and scientific applications of the National Accounting Machine. Supplement to the Journal of the Royal Statistical Society, 3(2):87–114, ???? 1936. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.
org/stable/2983666.

Comrie:1946:BDC

[Com46] L. J. Comrie. Babbage’s dream come true. Nature, 158(4017): 567–568, October 26, 1946. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL https://doi.org/
10.1038/158567a0.

Comrie:1947:MOA

[Com47] L. J. Comrie. A manual of operation for the Automatic Sequence Controlled Calculator. Nature, 158(4017):567– 568, October 26, 1947. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/
nature/journal/v158/n4017/
pdf/158567a0.pdf.

Cook:2005:MME

[Coo05] Simon Cook. Minds, machines and economic agents: Cambridge receptions of Boole and Babbage. Studies in History and Philosophy of Science Part A, 36(2):331–350, June 2005. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S003936810500018X.

Cortada:1987:HDD

[Cor87] James W. Cortada. Historical Dictionary of Data Processing: Biographies. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1987. ISBN 0-313-25651-9 (lib. bdg.). xiii + 321 pp. LCCN QA76.15 .C66 1987.

Crowley:1985:DBD

[Cro85] Mary L. Crowley. The “Difference” in Babbage’s Difference Engine. The Mathematics Teacher, 78(5):366–372, May 1985. ISSN 0025-5769 (print), 2330-0582 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27964533.

Croarken:2004:BRB

[Cro04a] Mary Croarken. Book review: The Cogwheel Brain: Charles Babbage and the Quest to Build the First Computer by Doron Swade. British Journal for the History of Science, 37(3):351– 352, September 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028434.

Croarken:2004:BRD

[Cro04b] Mary Croarken. Book review: Doron Swade, The Cogwheel Brain: Charles Babbage and the Quest to Build the First Computer. London: Little, Brown and Company, 2000. Pp. x + 342. ISBN 0-316-64847-7. £14.99 (hardback). British Journal for the History of Science, 37(3):

351–352, September 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028132;
http://www.jstor.org/stable/
4028434.

Caruzzi:1990:RHE

[CSW90] Paul Caruzzi, K. W. Smillie, and Eric Weiss. Reviews: Hyman, (ed.): Science and Reform. Selected Works of Charles Babbage; Goldberg (ed.): A History of Personal Workstations. Annals of the History of Computing, 12 (4):291–293, October/December 1990. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1990/pdf/a4291.pdf;   http:
//www.computer.org/annals/
an1990/a4291abs.htm.

Curley:2010:MII

[Cur10] Robert Curley, editor. The 100 most influential inventors of all time. The Britannica guide to the world’s most influential people. Britannica Educational Publishers in association with Rosen Educational Services, New York, NY, 2010. ISBN 1-61530-003-1. 336 pp. LCCN T39 .A14 2010.

Curley:2012:AIA

[Cur12] Robert Curley. Architects of the Information Age, volume 1 of Computing and Connecting in the 21st Century. Britannica Educational Publishing, Chicago, IL, USA, 2012. ISBN 1-61530-717-6 (print), 1-61530-661-7. 178 (est.) pp. LCCN ???? 29.95 (NL).

Ceruzzi:1991:RPI

[CW91] Paul Ceruzzi and Eric A. Weiss. Reviews: Penzias: Ideas and Information: Managing in a Hightech World; Rose: West of Eden: The End of Innocence at Apple Computer; Lindgren: Glory & Failure: the Difference Engines of Johann Müller, Charles Babbage, and Edvard Scheutz; Jennings: The Devouring Fungus: Tales of the Computer Age; Norberg: High Technology Calculation in the Early 20th Century: Punched Card Machinery in Business and Government; Hall and Barry: Sunburst: The Ascent of Sun Microsystems; Heppenheimer: How von Neumann Showed the Way; Aspray: Back to Basics: The Stored Program Concept; Rosen: The Origins of Modern Computing; Taylor: In Memoriam: J. C. R. Licklider, 1915–1990; Mounier-Kuhn: Genese de l’informatique en France (1945–1965): Diffusion de l’innovation et transfert de technologie; Eisler: My Life with the Printed Circuit. Annals of the History of Computing, 13(2):231–234, April/ June 1991. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1991/pdf/a2231.pdf;   http:
//www.computer.org/annals/
an1991/a2231abs.htm.

Dasgupta:2014:IBB

[Das14] Subrata Dasgupta. It began with Babbage: the genesis of computer science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-930941-8 (hardcover), 0-19-930942-6 (ebook). x + 328 pp. LCCN QA76.17 .D36 2014.

DeAngelis:2005:CPD

[DB05] Gina DeAngelis and David J.

Bianco. Computers: processing the data. Innovators. Oliver Press, Minneapolis, MN, USA, 2005. ISBN 1-881508-87-0. 143 pp. LCCN QA76.17 D43 2005.

DeLiso:2006:CBT

[De 06] Nicola De Liso. Charles Babbage, technological change and the national system of innovation. Journal of Institutional and Theoretical Economics (JITE) / Zeitschrift für die gesamte Staatswissenschaft, 162(3):470– 485, September 2006. URL

http://www.ingentaconnect.
com/content/mohr/jite/2006/
00000162/00000003/art00005;
http://www.jstor.org/stable/
40752596.

DeFranco:2022:CCL

[DeF22] Joanna F. DeFranco. From calculations to computations: A look back at the first computer. Computer, 55(1):82–87, January 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Daintith:1999:DS

[DG99] John Daintith and Derek Gjertsen, editors. A Dictionary of Scientists. Oxford paperback reference. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-585-11047-6 (e-book), 0-19-280086-8 (paperback). 586 pp. LCCN Q141 .D52 1994. URL

http://www.loc.gov/catdir/
enhancements/fy0637/99488304-
d.html;   http://www.loc.gov/
catdir/enhancements/fy0637/
99488304-t.html;   http://
www.netLibrary.com/urlapi.
asp?action=summary&v=1&bookid=
12306.

Decker:1990:AEI

[DH90] Rick Decker and Stuart Hirshfield. The Analytical Engine: an introduction to computer science using HyperCard. Wadsworth Publishing Company, Pacific Grove, CA, USA, 1990. ISBN 0-534-12540-9. xxii + 410 pp. LCCN QA76.5 .D37 1990.

Decker:1994:AEI

[DH94] Rick Decker and Stuart Hirshfield. The Analytical Engine: an introduction to computer science using HyperCard 2.1. PWS Publishing, Boston, MA, USA, second edition, 1994. ISBN 0-534-93696-2. xxiv + 391 pp. LCCN QA76.5 .D37 1994.

Decker:1998:AEI

[DH98] Rick Decker and Stuart Hirshfield. The Analytical Engine: an introduction to computer science using the Internet. PWS Publishing, Boston, MA, USA, 1998. ISBN 0-534-95365-4 (paperback). xx + 360 pp. LCCN QA76 .D33344 1998.

Decker:2004:AEI

[DH04] Rick Decker and Stuart Hirshfield. The Analytical Engine: an introduction to computer science using the Internet. Thomson/Wadsworth, Belmont, CA, second edition, 2004. ISBN 0-534-39159-1. xx + 412 pp. LCCN QA76 .D33344 2004.

Dijkstra:1986:TCB

[Dij86] Edsger W. Dijkstra. A theorem of Charles Babbage’s extended. Circulated privately., October 1986. URL http:

//www.cs.utexas.edu/users/
EWD/ewd09xx/EWD990.PDF.

Disch:1991:BRQ

[Dis91] Thomas M. Disch. Book review: Queen Victoria’s computers: The Difference Engine, by William Gibson and Bruce Sterling. 429 pp. New York: Bantam Books. $19.95. New York Times, ??(??):BR5–BR6, March 10, 1991. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL

http://search.proquest.com/
docview/108838347.

Dodge:1874:CBE

[Dod74] N. S. Dodge. Charles Babbage [eulogy]. In Annual Report of the Smithsonian Institution for 1873, pages 169–197. Smithsonian Institution, Washington, DC, USA, 1874. Reprinted in [Dod00].

Dodge:2000:CBE

[Dod00] Nathaniel Shatswell Dodge. Charles Babbage eulogy (from the Smithsonian Annual Report of 1873). IEEE Annals of the History of Computing, 22(4):22–43, October–December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

dlib.computer.org/an/books/
an2000/pdf/a4022.pdf;   http:
//ieeexplore.ieee.org/iel5/
85/19196/00887988.pdf;   http:
//www.computer.org/annals/
an2000/a4022abs.htm. Reprint

of the 1873 original, With an introduction by Martin Campbell-Kelly.

Dolan:1998:RNC

[Dol98] Brian P. Dolan. Representing novelty: Charles Babbage, Charles Lyell, and experiments in Early Victorian geology. History of Science (UK), 36(3):299–327, September 1998. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/36/
3/299.full.pdf+html.

Donnelly:1992:MDE

[Don92] James Donnelly. A modern Difference Engine: software simulators for Charles Babbage’s Difference Engine No. 2. Armstrong Pub. Co., Corvallis, OR, USA, 1992. ISBN 1-879828-03-0. viii + 36 pp. LCCN QA75 .D66 1992.

Dotzler:1996:BRA

[Dot96] Bernhard Dotzler, editor. Babbages Rechen-Automate: ausgewählte Schriften. (German) [Babbage’s calculating machines: selected works], volume 6 of Computerkultur. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-211-82640-8. viii + 502 pp. LCCN ???? URL http:/

/www.zentralblatt-math.org/
zmath/en/search/?an=0853.01012.

Dotzler:1997:TBU

[Dot97] Bernhard J. Dotzler. Technotation: Babbage und die Macht der Zeichen. (German) [Technotation: Babbage and the power of signs]. Weimarer Beiträge: Zeitschrift für Literaturwissenschaft, Ästhetik und Kulturwissenschaften, 43(1):99– 109, ???? 1997. ISSN 0043-2199.

Durand-Richard:1992:CBE

[DR92a] Marie-José Durand-Richard. Charles Babbage (1791–1871): de l’école algébrique anglaise à la “machine analytique”. (French) [Charles Babbage (1791–1871): from the English algebraic school to the Analytical Engine]. Math. Inform. Sci. Humaines, 118(??):5– 31, 1992. ISSN 0995-2314. See erratum [DR92b].

Durand-Richard:1992:ECB

[DR92b] Marie-José Durand-Richard. Erratum: “Charles Babbage (1791– 1871): from the English algebraic school to the ‘Analytical Engine’ ”. Math. Inform. Sci.

Humaines, 120(??):79–82, 1992. ISSN 0995-2314. See [DR92a].

Dutz:2021:BGD

[DS21] Jonah Dutz and Dirk Schlimm. Babbage’s guidelines for the design of mathematical notations. Studies in History and Philosophy of Science Part A, 88(??): 92–101, August 2021. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0039368121000315.

Dubbey:1977:BPM

[Dub77]  J. M. Dubbey.  Babbage,

Peacock and modern algebra. Historia Mathematica, 4(3): 295–302, August 1977. CO-DEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/031508607790057X.

Dubbey:1978:MWC

[Dub78] John Michael Dubbey. The mathematical work of Charles Babbage. Cambridge University Press, Cambridge, UK, 1978. ISBN 0-521-21649-4 (hardback), 0-521-52476-8 (paperback), 0-511-62239-2 (e-book). viii + 235 pp. LCCN QA27.G7 D8; QA29.B2. £17.95: CIP entry (May). URL http://

                                        www.gbv.de/dms/bowker/toc/
                                        9780521524766.pdf.

Essinger:2019:CACe

[EB19] James Essinger and Lisa Noel Babbage. Charles and Ada: the Computer’s Most Passionate Partnership. The History Press, Chicago, IL, USA, 2019. ISBN 0-7509-9286-7, 0-7509-9286-7 (e-book). 256 pp. LCCN QA29.L72 E87 2019. URL http:

//www.vlebooks.com/vleweb/
product/openreader?id=none\
%26isbn=9780750992862.

Eckert:1940:PCM

[Eck40] W. J. (Wallace John) Eckert.

Punched card methods in scientific computation. The Thomas J. Watson Astronomical Computing Bureau, Columbia University, New York, NY, USA, 1940. ix + 136 pp. LCCN QA75 .E3.

Eckert:1984:PCM

[Eck84] W. J. (Wallace John) Eckert.

Punched card methods in scientific computation, volume 5 of Charles Babbage Institute reprint series for the history of computing. MIT Press, Cambridge, MA, USA, 1984. ISBN 0-262-05030-7. xv + ix + 136 + 4 pp. LCCN QB51.3.E43 E28 1984. With a new introduction by J. C. McPherson.

Editors:1992:CBD

[Edi92] Editors. Charles Babbage describes his calculating engines. Chance, 5(1–2):56–62, Winter– Spring 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Eames:2005:FCR

[EEM^{+}05] Charles. Eames, Ray. Eames, Shelley. Mills, Lucia Eames. Demetrios, Eames. Demetrios, and Gregory Peck. The films of Charles and Ray Eames, Volume 6 [videorecording]. Pyramid, Santa Monica, CA, USA, 2005. 1 videodisc (60 min.) pp. LCCN PN1995.9.E96 .F557

2005 v.6. URL http://www.
midwesttapes.com. Narrator,

Gregory Peck (introduction).

Enros:1978:BRB

[Enr78] Philip C. Enros. Book review: The Mathematical Work of Charles Babbage, by J. M. Dubbey. American Scientist, 66 (5):639, September 1978. CO-DEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (elec-

tronic).  URL http://www.
jstor.org/stable/27848947.

Enros:1980:BRB

[Enr80a] Philip C. Enros. Book review: The mathematical work of Charles Babbage: By John Michael Dubbey. New York/ London (Cambridge University Press). 1978. viii + 235 pp. $26.50. Historia Mathematica, 7 (1):98–100, February 1980. CO-DEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0315086080900853.

Enros:1980:BRM

[Enr80b] Philip C. Enros. Book review: The mathematical work of Charles Babbage: By John Michael Dubbey. New York/ London (Cambridge University Press). 1978. viii + 235 pp. $26.50. Historia Mathematica, 7 (1):98–100, February 1980. CO-DEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0315086080900853.

Enros:1983:ASP

[Enr83] Philip C. Enros. The analytical society (1812–1813): Precursor of the renewal of Cambridge mathematics. Historia Mathematica, 10(1):24–47, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0315086083900319.

Enros:1988:BRB

[Enr88] Philip C. Enros. Book review: Memoir of the Life and Labours of the Late Charles Babbage Esq. F.R.S., by H. W. Buxton, Anthony Hyman. Isis, 79 (3):544, September 1988. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic).  URL http://www.
jstor.org/stable/234735.

Essinger:2004:JWH

[Ess04]  James Essinger.  Jacquard’s

Web: How a Hand-Loom Led to the Birth of the Information Age. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2004. ISBN 0-19-280577-0. xi + 302 pp. LCCN QA76.17 .E88 2004. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0620/2004276200-
                                        d.html;   http://www.loc.gov/
                                        catdir/enhancements/fy0620/
                                        2004276200-t.html.

Essinger:2014:AAH

[Ess14a] James Essinger. Ada’s algorithm: how Lord Byron’s daughter Ada Lovelace launched the digital age. Melville House, Brooklyn, NY, USA, 2014. ISBN 1-61219-408-7 (hardcover), 1-61219-409-5 (ebook). xvi + 254 pp. LCCN QA29.L72 E87 2014.

Essinger:2014:FGH

[Ess14b] James Essinger. A female genius: how Ada Lovelace Lord Byron’s daughter, started the computer age. Gibson Square, London, UK, 2014. ISBN 1-908096-66-7 (hardcover). 256 pp. LCCN QA29.L72 E87 2014.

Fahey:1991:RSD

[Fah91] Fin Fahey. Review: See the difference: Making the Difference: Charles Babbage and the Birth of the Computer An exhibition at the Science Museum to 31 December. New Scientist, ??(1779):??, July 27, 1991. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (elec-

                                        tronic).  URL http://www.
                                        newscientist.com/article/mg13117795.
                                        800-review-see-the-difference.
                                        html.

Farr:1864:ELT

[Far64] William Farr. English Life Table: Tables of Lifetimes, Annuities and Premiums. HMSO, London, UK, 1864. ???? pp. Includes chapter on “Scheutz’s Calculating Machine and its Use in the Construction of the English Life Table No. 3”.

Fuegi:2003:LBC

[FF03a] John Fuegi and Jo Francis. Lovelace & Babbage and the creation of the 1843 ‘Notes’. IEEE Annals of the History of Computing, 25 (4):16–26, October/December 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

csdl.computer.org/dl/mags/
an/2003/04/a4016.htm;   http:
//csdl.computer.org/dl/mags/
an/2003/04/a4016.pdf. Women

and gender in the history of computing.

Fuegi:2003:DTP

[FF03b] John Fuegi and Jo Francis. To dream tomorrow: a portrait of Ada Byron Lovelace. Flare Productions, Adelphi, MD, USA, 2003. Video DVD (53m) pp. LCCN ????

Fuegi:2015:LBC

[FF15] J. Fuegi and J. Francis. Lovelace & Babbage and the creation of the 1843 ’Notes’. Ada User Journal, 36(2):89–98, June 2015. CO-DEN AUJOET. ISSN 1381-6551. Bicentennial Ada Lovelace articles.

Fuegi:2003:ABL

[FFT^{+}03] John Fuegi, Jo Francis, Maisie Tomlinson, Steven Crossley, Georgina Terry, et al. Ada Byron Lovelace [videorecording]: to dream tomorrow. Women of power series. Flare Productions, Adelphi, MD, USA, 2003. 1 videodisc (53 min.) pp. LCCN QA29.L72 A333 2003.

England:1852:LCB

[Fin52] Financial Reform Association Liverpool, England. A letter to Charles Babbage, Esq., in reply to his ‘Thoughts on the principles of taxation, with reference to a property tax, and its exceptions’, volume 3 of Financial reform tracts. New series. G. Philip, Liverpool, UK, 1852. 28 pp. LCCN ????

Fisher:2010:TFI

[Fis10] Richard Fisher. Toy-de-force: Inside the steampunk Lego lab. New Scientist, ??(2792): ??, December 22, 2010. CO-DEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (elec-

                                        tronic).  URL http://www.
                                        newscientist.com/article/dn19848-
                                        toydeforce-inside-the-steampunk-
                                        lego-lab.html.

Fisch:2014:BTL

[Fis14] Menachem Fisch. Babbage’s two lives. British Journal for the History of Science, 47(1):95–118, March 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Fisch:2017:CUT

[Fis17] Menachem Fisch. Creatively undecided toward a history and philosophy of scientific agency. University of Chicago Press, Chicago, IL, USA, 2017. ISBN 0-226-51448-X (hardcover), 0-226-51451-X (paperback), 0-226-51465-X (e-book). vii + 295 pp. LCCN Q175 .F5452 2017.

Fogg:2001:BR

[Fog01] G. E. Fogg, C.B.E., F.R.S. Book reviews. Notes and Records of the Royal Society of London, 55 (2):331–346, May 22, 2001. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Forti:1975:BRB

[For75] Umberto Forti. Book review:

Babbage, la macchina analitica [(Italian) [Babbage: the Analytical Engine]]: By Mario G. Losano. Milan (Etas Kompas), 1973, 191 p. Historia Mathematica, 2(3):350–353, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0315086075900853.

Fraser:1963:BRB

[Fra63a] W. Fraser. Book review: Charles Babbage and his Calculating Engines: Selected Writings by Charles Babbage and Others, by Philip Morrison, Emily Morrison. American Mathematical Monthly, 70(5):587, May 1963. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic).  URL http://www.
jstor.org/stable/2312104.

Fraser:1963:RPC

[Fra63b] W. Fraser. Recent publications: Charles Babbage and his calculating engines: Selected Writings by Charles Babbage and Others, by Philip Morrison and Emily Morrison. American Mathematical Monthly, 70 (5):587, May 1963. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Franksen:1981:MBD

[Fra81] Ole Immanuel Franksen. Mr. Babbage, the Difference Engine, and the problem of notation: An account of the origin of recursiveness and conditionals in computer programming. International Journal of Engineering Science, 19(12):1657–1694, ???? 1981. CODEN IJESAN. ISSN 0020-7225 (print), 1879-2197 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0020722581901580.

Special issue dedicated to Prof. K. Kondo.

Franksen:1984:MBS

[Fra84] Ole Immanuel Franksen. Mr. Babbage’s secret: the tale of a cypher — and APL. Strandbergs Forlag, Birkerød, Danmark, 1984. ISBN 87-87200-86-4. 320 pp. LCCN Z103.B2 F72 1984. The tale of a cypher—and APL, With a foreword by H. H. Goldstine.

Franksen:1985:MBS

[Fra85] Ole Immanuel Franksen. Mr.

Babbage’s secret: the tale of a cypher and APL. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1985. ISBN 0-13-604729-7. 319 pp. LCCN Z103.B2 F721 1985.

Franksen:1986:SHM

[Fra86] O. I. Franksen. The secret hobby of Mr. Babbage. Systems Anal. Modelling Simulation, 3(2):183– 194, 1986. ISSN 0232-9298.

Franksen:1993:BCM

[Fra93] Ole Immanuel Franksen. Babbage and cryptography. Or, the mystery of Admiral Beaufort’s cipher. Mathematics and Computers in Simulation, 35(4): 327–367, October 1993. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/037847549390063Z.

Freedgood:2003:FPN

[Fre03] Elaine Freedgood, editor. Factory production in nineteenthcentury Britain, volume 2 of The Victorian archives series. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2003. ISBN 0-19-514872-X (paperback), 0-19-516101-7. xiv + 321 pp. LCCN TS155 .F4855 2003. URL

http://www.loc.gov/catdir/
enhancements/fy0613/2002025781-
d.html;   http://www.loc.gov/

catdir/enhancements/fy0613/
2002025781-t.html.

Friedman:1992:BBB

[Fri92] Linda Weiser Friedman. From Babbage to Babel and beyond: a brief history of programming languages. Computer Languages, 17(1):1–17, ???? 1992. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009605519290019J.

Friedman:2005:EDC

[Fri05] Ted Friedman. Electric dreams: computers in American culture. New York University Press, New York, NY, USA, 2005. ISBN 0-8147-2739-5 (hardcover), 0-8147-2740-9 (paperback). x + 275 pp. LCCN QA76.9.C66 F745 2005. URL

http://www.loc.gov/catdir/
toc/ecip0514/2005017512.html.

Flood:2011:MVB

[FRW11] Raymond Flood, Adrian C. (Adrian Clifford) Rice, and Robin Wilson, editors. Mathematics in Victorian Britain. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-960139-

9. ix + 466 pp. LCCN QA26 .M37 2011. URL

http://www.loc.gov/catdir/
enhancements/fy1211/2011938072-
b.html;   http://www.loc.gov/
catdir/enhancements/fy1211/
2011938072-d.html;   http:/
/www.loc.gov/catdir/enhancements/
fy1211/2011938072-t.html.

Fu:2008:ELC

[FZZ08] Xu Dan Fu, Xia Zhou, and Xiao Dong Zhao. Extensions of Lucas’ congruence and Babbage’s congruence. Acta Math. Sinica (Chin. Ser.), 51(4):693– 698, 2008. CODEN SHHPBO. ISSN 0583-1431.

Garwig:1969:CB

[Gar69] Paul L. Garwig. Charles Babbage (1792–1871). American Documentation, 20(4):320–324, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Graham-Cumming:2010:LBB

[GC10] John Graham-Cumming. Let’s build Babbage’s Analytical Engine. New Scientist, 208(2791): 26–27, December 18, 2010. CO-DEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0262407910631004.

Grattan-Guinness:1979:BMT

[GG79a] I. Grattan-Guinness. Babbage’s mathematics in its time. British Journal for the History of Science, 12(1):82–88, March 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4025665.

Grattan-Guinness:1979:BRB

[GG79b] I. Grattan-Guinness. Book review: Babbage’s mathematics in its time: The Mathematical Work of Charles Babbage. British Journal for the History of Science, 12(1):82–88, March 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4025665.

Grattan-Guinness:1990:WHP

[GG90] I. Grattan-Guinness. Work for the hairdressers: The production of de Prony’s logarithmic and trigonometric tables. Annals of the History of Computing, 12(3):177–185, July/September 1990. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1990/pdf/a3177.pdf;   http:
//www.computer.org/annals/
an1990/a3177abs.htm.

Grattan-Guinness:1992:BFB

[GG92a] I. Grattan-Guinness. The Babbage/Faraday Bicentenary Conference: Cambridge University, England, July 5–7, 1991. Historia Mathematica, 19(2):207, May 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/
article/pii/031508609290083N.

Grattan-Guinness:1992:CBA

[GG92b] Ivor Grattan-Guinness. Charles Babbage as an algorithmic thinker. IEEE Annals of the History of Computing, 14(3): 34–48, July/September 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

dlib.computer.org/an/books/

an1992/pdf/a3034.pdf;   http:
//www.computer.org/annals/
an1992/a3034abs.htm.

Grattan-Guinness:2005:GBI

[GG05] I. Grattan-Guinness. George Boole, an investigation of the laws of thought on which are founded the mathematical theory of logic and probabilities (1854). In I. Grattan-Guinness, Roger Cooke, Leo Corry, Pierre Crépel, and Niccolo Guicciardini, editors, Landmark Writings in Western Mathematics 1640– 1940, chapter 36, pages 470–

479. Elsevier, Amsterdam, The Netherlands, ???? 2005. ISBN 0-444-50871-6. URL http://www.

sciencedirect.com/science/
article/pii/B9780444508713501170.

Gibson:2019:SIH

[Gib19] Susannah Gibson. The Spirit of Inquiry: How One Extraordinary Society Shaped Modern Science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2019. ISBN 0-19-883337-7 (hardcover). xxi + 377 pp. LCCN Q41.C194 G537 2019.

Gigerenzer:2001:DCI

[Gig01] G. Gigerenzer. Digital computer:

Impact on the social sciences. In Neil J. Smelser and Paul B. Baltes, editors, International Encyclopedia of the Social & Behavioral Sciences, pages 3684– 3688. Elsevier, Amsterdam, The Netherlands, ???? 2001. ISBN 0-08-043076-7. URL http://www.

sciencedirect.com/science/
article/pii/B0080430767001017.

Gleick:2011:IHT

[Gle11] James Gleick. The information: a history, a theory, a flood. Fourth Estate, London, UK, 2011. ISBN 0-00-722573-3, 0-00-742311-X. 526 pp. LCCN ????

Goldstine:1972:CPN

[Gol72] Herman Heine Goldstine. The Computer: From Pascal to von Neumann. Princeton University Press, Princeton, NJ, USA, 1972. ISBN 0-691-02367-0 (paperback), 0-691-08104-2. xii + 378 pp. LCCN TK7885.A5 G64

                                        1993. URL http://www.jstor.
                                        org/stable/j.ctt7rvp0. Sec-

ond printing, 1973. Paperback edition 1980. Fifth printing, 1993 with new preface. Reprint 2000 by Books on Demand, Ann Arbor, MI, USA.

Goldstine:1993:CPN

[Gol93] Herman H. (Herman Heine) Goldstine. The Computer: from Pascal to von Neumann. Princeton University Press, Princeton, NJ, USA, 1993. ISBN 0-691-02367-0 (paperback), 0-691-08104-2 (hardback). xii + 378 pp. LCCN TK7885.A5 G64 1993.

                                        URL http://www.jstor.org/
                                        stable/j.ctt7rvp0.

Gooch:2002:LWC

[Goo02] Joanna Gooch. The life and works of Charles Babbage. Mathematics Today, 38(4):119–123, 2002. ISSN 1361-2042.

Geary:2008:PPM

[GP08] Frank Geary and Renee Prendergast. Philosophers and practical men: Charles Babbage, Irish merchants and the economics of information. The European Journal of the History of Economic Thought, 15(4):571–597, 2008. ISSN 0967-2567 (print), 1469-5936 (electronic).

Grant:1871:NDE

[Gra71] George B. Grant. On a new Difference Engine. American Journal of Science and Arts, II(??): 1–5, ???? 1871.

Green:2001:CBA

[Gre01] Christopher D. Green. Charles Babbage, the analytical engine, and the possibility of a 19thcentury cognitive science. In Green et al. [GST01], page ?? ISBN 1-55798-776-9. LCCN BF103 .T73 2001.

GBPHC:2006:MBI

[Gre06] Great Britain.Parliament.House of Commons. Mr Babbage’s invention Copies of the correspondence between the lords commissioners of His Majesty’s Treasury and the president and council of the Royal Society, relative to an invention of Mr. Babbage.: 19th century House of Commons sessional papers: Parliamentary papers, volume (370) of [1823]; ACCOUNTS AND PAPERS. Proquest LLC, Cambridge, England, 2006. 8 pp. LCCN ???? URL http://

                                        hdl.library.upenn.edu/1017.
                                        12/939089.

Gridgeman:1978:BRA

[Gri78] N. T. Gridgeman. Book review: Another Lucasian Professor: The Mathematical Work of Charles Babbage. Science, 200(4342):667, May 12, 1978. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/1746963.

Grier:2006:CER

[Gri06] David Alan Grier. CBQ essay review: Computer pioneer: Biographies of and writings by Charles Babbage (1791–1870). Communication Booknotes Quarterly, 37(1):7–13, ???? 2006. CODEN ???? ISSN 1094-8007 (print), 1532-6896 (elec-

                                        tronic).  URL http://www.
                                        tandfonline.com/doi/abs/10.
                                        1207/s15326896cbq3701_1.

Grier:2010:IYC

[Gri10] David Alan Grier. The inconsistent youth of Charles Babbage. IEEE Annals of the History of Computing, 32(4):18–31, October/December 2010. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Groce:2022:Pb

[Gro22] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 47(2):3, April 2022. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3520273.3520274.

Groce:2023:Pd

[Gro23] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 48(4):13, October 2023. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3617946.3617948.

Gibson:1991:DE

[GS91] William Gibson and Bruce Sterling. The Difference Engine. Bantam Books, New York, NY, USA, 1991. ISBN 0-553-29461-X. 429 pp. LCCN PS3557.I2264 D54 1991; PR9199.3.G53 D54 1992. US$5.99.

Gibson:2011:DE

[GS11] William Gibson and Bruce Sterling. The Difference Engine. Ballantine Books, New York, NY, USA, 20th anniversary edition, 2011. ISBN 0-440-42362-7 (paperback). xi + 492 pp. LCCN PS3557.I2264 D55 2011. Reprint of [GS91], with a new introduction.

Green:2001:TPI

[GST01] Christopher D. Green, Marlene Gay Shore, and Thomas Teo, editors. The transformation of psychology: influences of 19th century philosophy, technology, and natural science. American Psychological Association, Washington, DC, USA, 2001. ISBN 1-55798-776-9. xvii + 245 pp. LCCN BF103 .T73 2001.

Goto:1980:SHM

[GT80] Eiichi Goto and Motoaki Terashima. Studies on hashing — 3. MTAC — Mathematical Tabulative Automatic Computing. Journal of the Information Processing Society of Japan, 3(1):23–30, ???? 1980. CODEN JIPRDE. ISSN 0387-6101.

Gullard:1980:CBI

[Gul80] Pamela Gullard. The Charles Babbage Institute for the History of Information Processing. Annals of the History of Computing, 2(1):71–74, January/ March 1980. CODEN AHCOE5. ISSN 0164-1239. URL http://

                                        dlib.computer.org/an/books/
                                        an1980/pdf/a1071.pdf;   http:
                                        //www.computer.org/annals/
                                        an1980/a1071abs.htm.

Gullard:1981:CBI

[Gul81] Pamela Gullard. The Charles Babbage Institute for the History of Information Processing (CBI). Isis, 72(2):262–264, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/230974.

Halacy:1970:CBF

[Hal70] Daniel Stephen Halacy. Charles Babbage, father of the computer. Crowell-Collier Press, New York, NY, USA, 1970. 170 pp. LCCN QA29.B2 H35.

Hall:2009:GSP

[Hal09] Derek Hall, editor. Great scientists. Philosophy, invention, and engineering. Facts at your fingertips. Brown Bear Books, Redding, CT, USA, 2009. ISBN 1-933834-48-X. 64 pp. LCCN T39 .P53 2009.

Hartree:1949:CIM

[Har49a] Douglas R. (Douglas Rayner)

Hartree. Calculating Instruments and Machines. University of Illinois Press, Urbana, IL, USA, 1949. ix + 138 pp. LCCN QA85 .H29.

Hartree:1949:CBA

[Har49b] Douglas R. (Douglas Rayner)

Hartree. Charles Babbage and the Analytical Engine. In Calculating Instruments and Machines [Har49a], page ?? LCCN QA85 .H29.

Hartree:2012:CIM

[Har12] Douglas R. Hartree. Calculating Instruments and Machines. Cambridge University Press, Cambridge, UK, 2012. ISBN 1-107-63065-7 (paperback). URL

https://www.cambridge.org/
us/universitypress/subjects/
mathematics/historical-mathematical-
texts/calculating-instruments-

and-machines. Reprint of [?].

Herschel:1813:EDT

[HB13] John F. W. Herschel and Charles Babbage. On equations of differences and their application to the determination of functions from given conditions. Memoirs of the Analytical Society [Cambridge], pages 65–114, 1813. Reprinted in [Sch81].

Babbage:1822:BOM

[HB22] John Hershel and Charles Babbage. Barometrical observations, made at the fall of the Staubbach. Brewster’s Edinburgh Journal of Science, 6(??):224– 227, ???? 1822. CODEN ???? ISSN ????

Hooper:1855:JSA

[HBC^{+}55] George N. Hooper, Charles Babbage, Hyde Clarke, Thomas Webster, M. S. Bentham, and S. A. Good. Journal of the society for arts. The Journal of the Society of Arts, 4(159):29–44, December 7, 1855.

Herschel:1831:RMB

[Her31a] John F. W. Herschel. Report on Mr. Babbage’s Calculating Engine. Journal of The Franklin Institute, 11(3):210–213, March 1831. CODEN JFINAB. ISSN 0016-0032 (print), 1879-2693 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0016003231907515.

Herschel:1831:SLE

[Her31b] Sir John Frederick William Herschel. Suppressed letter to the editor of the Edinburgh Journal of Science on Sir David Brewster’s observations on the decline of science in England. Edinburgh, Scotland, 1831. ?? pp. Also titled Babbage on the decline of science.

Herschel:1857:RRS

[Her57] John F. W. Herschel. Report of the Royal Society Babbage Engine Committee. Report, Royal Society, London, UK, 1857.

Herken:1988:UTM

[Her88] Rolf Herken. The Universal Turing machine: a half-century survey. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1988. ISBN 0-19-853741-7. xiv + 661 pp. LCCN QA267.U55 1988; QA267 .U941 1988.

Herschel:1989:RBE

[Her89] John F. W. Herschel. Report of the Babbage engine committee (1829). In Campbell-Kelly [CK89k], pages 108–114. ISBN 1-85196-502-5, 1-85196-005-8 (set). LCCN ????

Hershman:1999:DEI

[Her99] Lynn Hershman. Difference Engine #3. Leonardo (Oxford, England), 32(4):269–270, ???? 1999. CODEN LEONDP. ISSN 0024-094X (print), 1530-9282 (elec-

tronic).  URL http://www.
jstor.org/stable/1576718.

Hyman:1979:RDM

[HG79] Anthony Hyman and N. T.

Gridgeman. Reviews: D. L. Moore: Ada, Countess of Lovelace; J. M. Dubbey: The Mathematical Work of Charles Babbage. Annals of the History of Computing, 1 (1):75–76, July/September 1979. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1979/pdf/a1075.pdf;   http:
//www.computer.org/annals/
an1979/a1075abs.htm.

Huskey:1980:LLC

[HH80] Velma R. Huskey and Harry D. Huskey. Lady Lovelace and Charles Babbage. Annals of the History of Computing, 2 (4):299–329, October/December 1980. CODEN AHCOE5. ISSN 0164-1239. URL http://

                                        dlib.computer.org/an/books/
                                        an1980/pdf/a4299.pdf;   http:
                                        //www.computer.org/annals/
                                        an1980/a4299abs.htm.

Husbands:2008:MMH

[HHW08] Philip Husbands, Owen Holland, and Michael Wheeler, editors. The mechanical mind in history. MIT Press, Cambridge, MA, USA, 2008. ISBN 0-262-08377-9 (hardcover). viii + 458 pp. LCCN Q335 .M3956 2008.

Hill:1979:BRB

[Hil79] I. D. Hill. Book review: The Mathematical Work of Charles Babbage, by J. M. Dubbey. Journal of the Royal Statistical Society. Series A (General), 142 (1):63–64, ???? 1979. CO-DEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.
                                        org/stable/2344654.

Hill:1985:BRBa

[Hil85] I. D. Hill. Book review: Charles Babbage — Pioneer of the Computer, by Anthony Hyman. Journal of the Royal Statistical Society. Series A (General), 148 (2):171, ???? 1985. CO-DEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.
                                        org/stable/2981961.

Harman:2002:CSM

[HM02] P. M. (Peter Michael) Harman and Simon Mitton, editors. Cambridge scientific minds. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-78100-0 (hardback), 0-521-78612-6 (paperback). viii + 343 pp. LCCN Q141 .C1283 2002.

Hazzan:2022:BDS

[HMB22] Orit Hazzan, Koby Mike, and Herbert Bruderer. BLOG@CACM: Up with data science, and the first programmer. Communications of the ACM, 65(11): 10–11, November 2022. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3563107.

Hartree:1948:DCM

[HNW^{+}48] D. R. Hartree, M. H. A. Newman, Maurice V. Wilkes, F. C. Williams, James H. Wilkinson, and A. D. Booth. A discussion on computing machines. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 195(1042): 265–287 (2 plates), December 22, 1948. CODEN PRLAAZ. ISSN 0962-8444 (print), 2053-9177 (electronic). URL https:/

/ui.adsabs.harvard.edu/abs/
1948RSPSA.195..265H.

Hooper:2012:ALM

[Hoo12] Rowan Hooper. Ada Lovelace:

My brain is more than merely mortal. New Scientist, 216 (2886):29, October 13, 2012. CO-DEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0262407912626335.

Horsley:1909:DBM

[Hor09]  Victor Horsley.  Description

of the brain of Mr. Charles Babbage, F.R.S. Philosophical Transactions of the Royal Society of London. Series B, Containing Papers of a Biological Character, 200(3):117–131, ????

1909. URL http://www.jstor.
org/stable/91929.

Haigh:2016:HRW

[HP16] Thomas Haigh and Mark Priestley. Historical reflections: Where code comes from: architectures of automatic control from Babbage to Algol. Communications of the ACM, 59(1):39– 44, January 2016. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://cacm.acm.org/
magazines/2015/1/195721/fulltext.

Holzl:2003:BPA

[HR03] Werner Hölzl and Andreas Reinstaller. The Babbage Principle after evolutionary economics, volume 2003-013 of MERIT-Infonomics research memorandum series. MERIT, Maastricht, The Netherlands, 2003. 19 pp.

LCCN ???? URL http://www.
merit.unu.edu/publications/
rmpdf/2003/rm2003-013.pdf.

Hammerman:2016:AL

[HR16] Robin Hammerman and Andrew L. Russell. Ada’s legacy: cultures of computing from the Victorian to the digital age, volume 7 of ACM books. ACM Press, New York, NY 10036, USA, 2016. ISBN 1-970001-51-8 (hardcover), 1-970001-48-8 (paperback), 1-970001-49-6 (ebook), 1-970001-50-X (ePub). ISSN 2374-6777. xii + 249 pp. LCCN QA76.2.L72 H252 2016.

Hunter:1996:BRD

[Hun96] Lawrence Hunter. Book review: Digital days: The story of the computer, from the 17th Century to the Information Age. Computer: A History of the Information Machine. By Martin Campbell-Kelly and William Aspray. Illustrated. 342 pp. New York: Basic Books. $28. New York Times, ??(??):BR29, November 17, 1996. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL

http://search.proquest.com/
docview/109658057.

Hyman:1982:CBPa

[Hym82a] Anthony Hyman. Charles Babbage: pioneer of the computer. Oxford University Press paperback. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1982. ISBN 0-19-281491-5; 0-19-858170-X. xi + 287 + 32 pp. LCCN QA29.B2 H93 1982.

Hyman:1982:CBPb

[Hym82b] Anthony Hyman. Charles Babbage: pioneer of the computer. Princeton University Press, Princeton, NJ, USA, 1982. ISBN 0-691-08303-7, 0-691-02377-8 (paperback). xi + 287 + 33 pp. LCCN QA29.B2 H93 1982; Q143.B2 H9.

Hyman:1983:CB

[Hym83] Anthony Hyman. Charles Babbage. Princeton University Press, Princeton, NJ, USA, repr. edition, 1983. ISBN 0-691-08303-7. 287 pp. LCCN ????

Hyman:1987:CBP

[Hym87] Anthony Hyman. Charles Babbage: 1791–1871; Philosoph, Mathematiker, Computerpionier. (German) [Charles Babbage: 1791–1871]: Philosopher, Mathematician, Computer Pioneer. Klett-Cotta, Stuttgart, West Germany, 1987. ISBN 3-608-93095-7. 457 pp. LCCN ????

Hyman:1988:SRS

[Hym88] Anthony Hyman, editor. Science and reform: selected works of Charles Babbage. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-34311-9. vii + 356 + 18 pp. LCCN QA3

                                        .B2 1988. URL http://www.
                                        loc.gov/catdir/description/
                                        cam023/88010294.html.

Hyman:1992:CQD

[Hym92] Anthony Hyman. Comments, queries, and debate: The private Babbage. IEEE Annals of the History of Computing, 14(1):7, January–March 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

dlib.computer.org/an/books/
an1992/pdf/a1006.pdf.

Hyman:2002:CBS

[Hym02]  Anthony Hyman.  Charles

Babbage: science and reform. In Cambridge scientific minds, pages 79–93. Cambridge University Press, Cambridge, UK, 2002.

Isaacson:2014:IHG

[Isa14] Walter Isaacson. The Innovators: How a Group of Hackers, Geniuses and Geeks Created the Digital Revolution. Simon and Schuster, New York, NY, USA, 2014. ISBN 1-4711-3879-8 (hardcover), 1-4767-0869-X (cloth), 1-4711-3897-6 (paperback), 1-4104-7497-6 (cloth), 1-4767-0870-3. viii + 542 pp. LCCN QA76.2.A2.

Jarczyk:2003:BEC

[Jar03] Witold Jarczyk. Babbage equation on the circle. Publicationes Mathematicae Debrecen, 63(3):389–400, 2003. CODEN PUMAAR. ISSN 0033-3883 (print), 2064-2849 (electronic).

Johnson:1990:FHB

[JC90] W. Johnson and S. Chandrasekar. Forming holes in brittle materials: An antique technique and its report by Charles Babbage, 1792–1871. Journal of Materials Processing Technology, 21(1):43–50, January 1990. CODEN ???? ISSN

                                        ????  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/092401369090028S.

Jennings:1985:PCM

[Jen85] Humphrey Jennings, editor. Pandæmonium: the coming of the machine as seen by contemporary observers, 1660–1886. Free Press, New York, NY, USA, 1985. ISBN 0-02-916470-2. xxxviii + 376 pp. LCCN PR1111.I58 P3 1985. Edited by Mary-Lou Jennings and Charles Madge. Foreword by Frank Cottrell Boyce.

Jensen:2000:WWE

[Jen00] Peter R. Jensen. From the wireless to the web: the evolution of telecommunications, 1901–2001. University of New South Wales Press, Sydney, NSW, Australia, 2000. ISBN 0-86840-458-6. 306 pp. LCCN TK5102.2 .J46 2000.

Jerwood:1989:CPM

[Jer89] J. Jerwood. On the calculation of printing of mathematical tables by machinery (1861). In Campbell-Kelly [CK89h], pages 1–36. ISBN 1-85196-501-7, 1-85196-005-8 (set). LCCN ????

Jones:2011:DHC

[JL11] Cliff B. Jones and John L. Lloyd, editors. Dependable and Historic Computing: Essays Dedicated to Brian Randell on the Occasion of His 75th Birthday, volume 6875 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 3-642-24540-4 (paperback), 3-642-24541-2 (ebook). xii + 523 + 137 pp. LCCN QA76.17.

Johnson:1992:BMW

[Joh91] Paul Johnson. The birth of the modern: world society, 1815– 1830. HarperCollins Publishers, New York, NY, 1991. ISBN 0-06-092282-6. xx + 1095 pp. LCCN D383 .J64 1992.

Johnson:1992:DFT

[Joh92] W. Johnson. Developments in forming technology — an engineering educator’s approach. Journal of Materials Processing Technology, 31(1–2):1–26, May 1992. CODEN ???? ISSN

????  URL http://www.
sciencedirect.com/science/
article/pii/092401369290003B.

Johnson:1995:CCW

[Joh95] W. Johnson. Contents and commentary on William Moore’s A treatise on the motion of rockets and an essay on naval gunnery. International Journal of Impact Engineering, 16(3):499– 521, June 1995. CODEN ????

ISSN ???? URL http://www.
sciencedirect.com/science/
article/pii/0734743X9400052X.

Jones:1854:RO

[Jon54] T. Wharton Jones, F.R.S. Report on the opthalmoscope.

British and Foreign Medical Review, 14(??):551–??, October 1854. Contains short description of Mr. Babbage’s Opthalmoscope.

Jones:1991:LTC

[Jon91] Glyn Jones. The life and times of a computing pioneer: Charles Babbage invented all sorts of things besides automatic calculating machines, including ‘cow catchers’, flashing lighthouses and speedometers. Did he spread his talents too thinly? New Scientist, ??(1775):??, June 29, 1991. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic). URL http://www.

newscientist.com/article/mg13017755.
600-the-life-and-times-of-
a-computing-pioneer-charles-
babbage-invented-all-sorts-
of-things-besides-automatic-
calculating-machines-including-
cow-catchers-flashing-lighthouses-
and-speedometers-did-he-spread-
his-talents-too-thinly.html.

Jones:2016:RMC

[Jon16] Matthew L. (Matthew Laurence) Jones. Reckoning with Matter: Calculating Machines, Innovation, and Thinking About Thinking from Pascal to Babbage. University of Chicago Press, Chicago, IL, USA, 2016. ISBN 0-226-41146-X (hardcover), 0-226-41163-X (e-book). 331 pp. LCCN QA75 .J66 2016.

Kanelopoulos:1994:BET

[Kan94] Charles Kanelopoulos. Babbage: entre la technologie et l’économie. (French) [Babbage: between technology and economy]. Économies et sociétés: cahiers de l’ISMEA (Paris), 28 (4):5–68, 1994. CODEN ???? ISSN 0013-0567.

Kemp:1979:BRB

[Kem79] Emory Kemp. Book review:

John Scott Russell, by George S. Emmerson; The Mathematical Work of Charles Babbage, by J. M. Dubbey; Sir Charles Wheatstone, by Brian Bowers. Victorian Studies, 23(1):127–129, Autumn 1979. ISSN 0042-5222 (print), 1527-2052 (elec-

tronic).  URL http://www.
jstor.org/stable/3827448.

Kilgore:2010:DEA

[Kil10] De Witt Douglas Kilgore. Difference Engine: Aliens, robots, and other racial matters in the history of science fiction. Science Fiction Studies, 37(1):16– 22, March 2010. ISSN 0091-7729 (print), 2327-6207 (electronic).

Kirpal:20xx:BC

[Kirxx] Alfred Kirpal. Babbage, Charles (1791–1871). ????, ????, 20xx. LCCN ????

Krysa:2012:ALN

[KM12] Joasia Krysa and Luigi Federico Menabrea. Ada Lovelace: 100 Notes, 100 Thought, volume 055. Hatje Cantz Verlag, Kassel, Germany, 2012. ISBN 3-7757-2904-6 (paperback). 34 pp. LCCN QA29.L72 A33 2011.

Knott:1915:NTM

[Kno15] Cargill Gilston Knott, editor. Napier tercentenary memorial volume. Published for the Royal society of Edinburgh, by Longmans, Green and Company, London, UK, 1915. xi + 441 + 1 pp. LCCN QA29.N2 K5. URL http:

//resolver.library.cornell.
edu/math/2013807.

Koutalis:2014:MDB

[Kou14] Vangelis Koutalis. Making discoveries for a better life vs. bringing fruits to the national treasury: Davy, Babbage, Brewster and the (ongoing) struggle for the soul of science. Almagest: International Journal for the History of Scientific Ideas, 5(1): 52–109, 2014. CODEN ???? ISSN 1792-2593. URL https:/

/www.brepolsonline.net/doi/
epdf/10.1484/J.ALMAGEST.5.
102473.

Kim:1999:AFC

[KT99] Eugene Eric Kim and Betty Alexandra Toole. Ada and the first computer: The collaboration between Ada, Countess of Lovelace, and computer pioneer Charles Babbage resulted in a landmark publication that described how to program the world’s first computer. Scientific American, 280(5):76–81, May 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). See also [AAG^{+}00].

L:1990:BRB

[L.90] K. L. Book review: Computing Before Computers. William Aspray, ed. Iowa State University Press, Ames, 1990. x, 266 pp., illus. $27.95. Science, 248 (4963):1670, June 29, 1990. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.sciencemag.
org/content/248/4963/1670.
2.citation?.

LaCroix:1816:ETD

[LaC16] Sylvestre F. LaCroix. An Elementary Treatise on the Differential and Integral Calculus. J. Deighton, Cambridge, UK, 1816. ???? pp. Translation from French by Charles Babbage, John Herschel, and George Peacock.

Lardner:1834:BCE

[Lar34] Dionysius Lardner. Babbage’s calculating engines. Edinburgh Review, 59(??):263–327, ???? 1834.

Lardner:1989:BCE

[Lar89] Dionysius Lardner. Babbage’s calculating engine. In Campbell-Kelly [CK89k], pages 118–186. ISBN 1-85196-502-5, 1-85196-005-8 (set). LCCN ????

Lavington:2014:BRI

[Lav14] Simon Lavington. Book review:

It began with Babbage. Resurrection: The Computer Conservation Society Journal, ?? (65):??, Spring 2014. ISSN 0958-7403. URL https://

computerconservationsociety.
org/resurrection/res65.htm#
f.

Leatherdale:2020:BRC

[Lea20]  Dik Leatherdale.  Book re-

view: Charles & Ada: The computer’s most passionate partnership. Resurrection: The Journal of the Computer Conservation Society, ??(89):??, Spring 2020. ISSN 0958-7403. URL https://

                                        computerconservationsociety.
                                        org/resurrection/res89.htm#
                                        e.

Lee:1995:BKH

[Lee95]  J. A. N. Lee.  On “Bab-

bage and Kings” and “How sausage was made”: and now for the rest of the story. IEEE Annals of the History of Computing, 17(4):7–23, Winter 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

                                        //ieeexplore.ieee.org/iel4/
                                        85/10129/00477433.pdf.

Lewis:2003:BBO

[Lew03] Michael A. Lewis. Babbage and the beginnings of OM: reappraising on the economy of machinery and manufactures. In M. A. Lewis and N. Slack, editors, Operations management: Critical perspectives on business and management, chapter 6, pages 54–72. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2003. LCCN ????

Lewis:2007:CBR

[Lew07]  Michael A. Lewis.  Charles

Babbage: Reclaiming an operations management pioneer.

Journal of Operations Management, 25(2):248–259, March 2007. CODEN ???? ISSN 0272-6963 (print), 1873-1317 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0272696306000805.

Lewis:2021:ICF

[Lew21] Harry R. Lewis. Ideas That Created the Future: Classic Papers of Computer Science. MIT Press, Cambridge, MA, USA, 2021. ISBN 0-262-04530-3. xxii + 495 pp. LCCN Q124.6-127.2.

Lindgren:1987:GFD

[Lin87] Michael Lindgren. Glory and failure: the difference engines of Johann Müller, Charles Babbage and Georg and Edvard Scheutz, volume 9; 2017 of Linköping studies in arts and science; Stockholm papers in history and philosophy of technology. Department of Technology and Social Change, Linköping University, Linköping, Sweden, 1987. ISBN 91-7870-146-5. 415 pp. LCCN QA76.2.A2 L51 1987. Originally presented as the author’s doctoral thesis, Linköping University, 1987. Translation from Swedish by Craig G. McKay.

Lindgren:1990:GFD

[Lin90] Michael Lindgren. Glory and failure: the difference engines of Johann Müller, Charles Babbage and Georg and Edvard Scheutz. The MIT Press series in the history of computing. MIT Press, Cambridge, MA, USA, second edition, 1990. ISBN 0-585-35107-4 (e-book), 0-262-12146-8. 414 pp. LCCN QA75 .L493 1990. Translation from German by Craig G. McKay.

Lindgren:1982:SFD

[LL82] Michael Lindgren and Svante Lindqvist. Scheutz’s first Difference Engine rediscovered. Technology and Culture, 23(2):207– 213, April 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

                                        URL http://www.jstor.org/
                                        stable/3104133.

Losano:1972:CBP

[Los72] Mario G. Losano. Charles Babbage e la programmazione delle macchine da calcolo. Atti Accad. Sci. Torino Cl. Sci. Fis. Mat. Natur., 106:25–37, 1972. ISSN 0001-4419.

Lovelace:1843:NTL

[Lov43] Augusta Ada Lovelace. Notes [on translation of Luigi Federico Menebrae’s paper on Babbage’s Analytical Engine]. In Taylor’s Scientific Memoirs, page ???? ????, ????, August 1843. LCCN ???? Translation to English, with notes about three times the length of the original French article, of [Men43]. See [Woo99, pages 260–282] and [Swa01b, pages 162–164] for accounts of how those notes were written, and despite vigorous opposition from Charles Babbage [Bab43a], finally published in this book. The author was listed only as the initials ‘A. A. L.’, because it was then considered inappropriate for a woman, particularly one of high rank, to author a scientific paper.

Lovelace:2012:AL

[Lov12] Ada Lovelace. Ada Lovelace.

Hatje Cantz Verlag, Kassel, Germany, 2012. ISBN 3-7757-2904-6, 3-7757-3084-2. 34 pp. LCCN

????  URL http://ebooks.
ciando.com/book/index.cfm/
bok_id/304705;   http://www.
ciando.com/img/books/width167/
3775730842_k.jpg;   http://
www.ciando.com/pictures/bib/
3775730842bib_t_1.jpg.

Lieberman:1967:EDC

[LR67] Henry D. Lieberman and Louis Robinson. The electronic digital computer: How it started, how it works and what it does. New York Times, ??(??):136–137, January 9, 1967. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL

http://search.proquest.com/
docview/118029058?accountid=
14677.

Lyons:1940:CBO

[Lyo40] H. G. Lyons. Charles Babbage and the ophthalmoscope. Notes and Records of the Royal Society of London, 3(1):146– 148, April 1, 1940. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic).  URL http://www.
jstor.org/stable/531152.

Macfarlane:1919:LTB

[Mac19] Alexander Macfarlane. Lectures on ten British physicists of the nineteenth century, volume 20 of Mathematical monographs. Wiley, New York, NY, USA, 1919. 144 + 10 pp. LCCN ????

Mach:2007:SFE

[Mac07] Andrzej Mach. On some functional equations involving Babbage equation. Results Math., 51 (1-2):97–106, 2007. ISSN 1422-6383.

Maddox:1986:BRB

[Mad86]  Brenda Maddox.  Book re-

view: Byronic, despite everything: The Calculating Passion of Ada Byron, by Joan Baum. Illustrated. 133 pp. Hamden, Conn. Archon Books/The Shoe String Press. $21.50. New York Times, ??(??):519, October 5, 1986. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL

                                        http://search.proquest.com/
                                        docview/110969134.

Martin:1925:RIE

[Mar25] Ernst Martin. Die Rechenmaschinen und ihre Entwicklungsgeschichte. Teil 1. Rechenmaschinen mit automatischer Zehnerübertragung. (German) [The Calculating Machines and the history of their development. Part 1. Calculating machines with automatic tens carry]. Meyer, Pappenheim, Germany, 1925. 477 pp. LCCN

???? US$. URL http://www.
xnumber.com/xnumber/cmisc_

martin.htm. English translation in [MKW92].

Martin:1986:RIE

[Mar86] Ernst Martin. Die Rechenmaschinen und ihre Entwicklungsgeschichte: Rechenmaschinen mit automatischer Zehnerübertragung. (German) [The Calculating Machines and the history of their development. Part 1. Calculating machines with automatic tens carry], volume 1. Köntopp, Leopoldshöhe, West Germany, 1986. ISBN 3-9801485-2-1. LCCN ???? Reprint of [Mar25].

Markoff:2008:SCC

[Mar08] John Markoff. 1800s-style computer comes to U.S. New York Times, ??(??):C6, May 5, 2008. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.
proquest.com/docview/897116572.

Marsh:2022:CBD

[Mar22] Allison Marsh. Charles Babbage’s Difference Engine turns 200: Error-riddled astronomical tables inspired the first computer — and the first vaporware. IEEE Spectrum, ??(??):??, May 27, 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic). URL https:/

/spectrum.ieee.org/charles-
babbage-difference-engine.

Matthews:1990:VNC

[Mat90] R. Matthews. Victorian number cruncher finally gets past the drawing board. Sunday Correspondent, ??(??):5, September 16, 1990.

Matthews:1991:CRG

[Mat91] Charles Matthews. Computer revolutionaries: A guide to the literature on pioneers in computing. Science & Technology Libraries, 11(4):43–74, ???? 1991. CODEN ???? ISSN 0194-262X (print), 1541-1109 (electronic). URL http://www.

                                        tandfonline.com/doi/abs/10.
                                        1300/J122v11n04_03.

Muller:1989:EDJ

[MB89] J. H. Müller and Ralf Bülow. Ein Entwurf für eine Differenzmaschine aus dem Jahr 1784. (German) [A design for a difference engine from the year 1784]. Sudhoffs Archiv: Zeitschrift für Wissenschaftsgeschichte, 73(2): 219–222, ???? 1989. CO-DEN SUARAH. ISSN 0039-

                                        4564. URL http://www.jstor.
                                        org/stable/20777242.

McClain:1997:EC

[McC97] Dylan Loeb McClain. The evolution of the calculator. New York Times, ??(??):D3, September 1, 1997. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL

                                        http://search.proquest.com/
                                        docview/109786292.

Maistrov:1978:CBD

[MÈ78] L. E. Maĭstrov and I. S. Èdlin. C.

Babbage’s difference engine. In History and methodology of the natural sciences, No. XX (Russian), pages 99–105. Moskov. Gos. Univ., Moscow, USSR, 1978.

Menabrea:1842:NMA

[Men43] Luigi F. Menabrea. Notions sur la machine analytique de M. Charles Babbage. Bibliothèque Universelle de Genève, 41:352–376, 1842–1843. CODEN ???? ISSN ???? URL http:

//www.fourmilab.ch/babbage/

sketch.html. Translated, with additional notes by Augusta Ada, Countess of Lovelace, as Sketch of the Analytical Engine. Reprinted in [CK89l, pages 62– 82]. Goldstine [Gol72, page 19] cites this paper as number 82, reprinted in Taylor’s Scientific Memoirs 111, note A, page 348. Woolley [Woo99, page 400] says it appears in that journal in volume 3, pages 666–731, 1843.

Menabrea:2021:SAEa

[Men21a] L. F. Menabrea. Sketch of the Analytical Engine (1843) / with notes by the translator, Ada Agusta, Countess of Lovelace. In Ideas That Created the Future: Classic Papers of Computer Science [Lew21], chapter 3, pages 9– 26. ISBN 0-262-04530-3. LCCN Q124.6-127.2.

Menabrea:2021:SAEb

[Men21b] Luigi Federico Menabrea. Sketch of the Analytical Engine Invented by Charles Babbage: Translation and Notes by Ada Lovelace. Quaternion Books, ????, 2021. 99 pp. LCCN ???? URL https://

quaternionbooks.com/product/
sketch-of-the-analytical-engine-
translation-and-notes-by-ada-
lovelace/.

Merzbach:1977:GSF

[Mer77] Uta C. Merzbach. Georg Scheutz and the first printing calculator, volume 36 of Smithsonian studies in history and technology. Smithsonian Institution Press, Washington, DC, USA, 1977. iii + 74 pp. LCCN QA75 .M46.

Merrifield:1982:RCL

[Mer82] C. W. Merrifield. Report of the committee, consisting of Professor Cayley, Dr. Farr, Mr. J. W. L. Glaisher, Dr. Pole, Professor Fuller, Professor A. B. W. Kennedy, Professor Clifford, and Mr. C. W. Merrifield, appointed to consider the advisability and to estimate the expense of constructing Mr. Babbage’s Analytical Machine, and of printing tables by its means (1879). In Randell [Ran82b], pages 55–

65. ISBN 0-387-11319-3, 3-540-11319-3. LCCN TK7885.A5 O741 1982.

Misa:2007:ANC

[Mis07] Thomas J. Misa. Arthur Norberg, the Charles Babbage Institute, and the history of computing. IEEE Annals of the History of Computing, 29(4):8–15, October/December 2007. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Martin:1992:CMR

[MKW92] Ernst Martin, Peggy Aldrich Kidwell, and Michael R. (Michael Roy) Williams, editors. The Calculating Machines (Die Rechenmaschinen): Their History and Development, volume 16 of Charles Babbage Institute reprint series for the history of computing. MIT Press, Cambridge, MA, USA, 1992. ISBN 0-262-13278-8. xvii + 367 + 25 pp. LCCN QA75 .M3513 1992.

Morrison:1952:SLC

[MM52] Philip Morrison and Emily Morrison. The strange life of Charles Babbage. Scientific American, 186(4):66–73, April 1952. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/
scientificamerican/journal/
v186/n4/pdf/scientificamerican0452-
66.pdf.

Morrison:1961:CBH

[MM61a] Philip Morrison and Emily Morrison, editors. Charles Babbage and his calculating engines: selected writings, volume T12 of Dover books. Dover Publications, Inc., New York, NY, USA, 1961. xxxviii + 400 pp. LCCN QA75 .B112.

Morrison:1961:CBP

[MM61b] Philip Morrison and Emily Morrison, editors. Charles Babbage on the principles and development of the calculator: and other seminal writings. Dover Publications, Inc., New York, NY, USA, 1961. ISBN 0-486-24691-4 (paperback). xxxviii + 400 pp. LCCN QA75 .C52 1961. US$7.95.

Moore:2011:NC

[MM11] Christopher Moore and Stephan Mertens. The nature of computation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-923321-7. xvii + 985 pp. LCCN ????

Maulucci:1988:HCC

[MN88] Ruth A. Maulucci and Arthur L. Norberg. Happenings: Computing in the 21st Century: A Charles Babbage Institute Symposium. Annals of the History of Computing, 10(2):127– 132, April/June 1988. CO-DEN AHCOE5. ISSN 0164-1239.

URL http://dlib.computer.
org/an/books/an1988/pdf/a2127.
pdf;   http://www.computer.
org/annals/an1988/a2127abs.
htm.

Morrison:1995:NTW

[Mor95] Philip Morrison. Nothing is too wonderful to be true, volume 11 of Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1995. ISBN 1-56396-363-9 (hardcover). xi + 446 pp. LCCN Q173.

Moseley:1964:IGLa

[Mos64a] Maboth Moseley. Irascible genius: a life of Charles Babbage, inventor. Hutchinson, London, UK, 1964. 287 + 4 pp. LCCN QA 29 .B3M6 1964A; T40.B25 M6. Foreword by B. V. Bowden.

Moseley:1964:IGLb

[Mos64b] Maboth Moseley. Irascible genius; the life of Charles Babbage. H. Regnery Co., Chicago, IL, USA, 1964. 287 pp. LCCN T40.B11 M898 1970.

Moseley:1965:IGL

[Mos65] Maboth Moseley. Irascible genius; a life of Charles Babbage, inventor. Augustus M. Kelley, New York, NY, USA, 1965. 287 pp. LCCN T/40/B25/M67 SS. Reprints of economic classics.

Mosconi:1983:CBV

[Mos83] Jean Mosconi. Charles Babbage: vers une théorie du calcul mécanique. (French) [Charles Babbage: toward a theory of mechanical calculation]. Revue d’Histoire des Sciences et de Leurs Applications, 36(1):69– 107, 1983. CODEN RHSAAM. ISSN 0048-7996, 0151-4105, 1969-6582.

Moulton:1915:ILG

[Mou15]  Lord Moulton.  The inven-

tion of logarithms, its genesis and growth. In Knott [Kno15], pages 1–24. LCCN QA29.N2 K5. URL http:/

/resolver.library.cornell.
edu/math/2013807.

Morrell:1981:GSE

[MT81] Jack Morrell and Arnold Thackray. Gentlemen of science: early years of the British Association for the Advancement of Science. Clarendon Press, New York, NY, USA, 1981. ISBN 0-19-858163-

7. xxiii + 592 + 16 pp. LCCN Q41.B85 M67 1981. US$49.95 (U.S.).

Mobbs:2013:ABB

[MU13] Tessa Mobbs and Robert William Unwin. Additions to the Babbage bibliography. Notes and Records of the Royal Society of London, 67(1):91–93, March 20, 2013. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Mullett:1948:CBS

[Mul48] Charles F. Mullett. Charles Babbage: a scientific gadfly. The Scientific Monthly, 67(5): 361–371, November 1948. CO-DEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (elec-

tronic).  URL http://www.
jstor.org/stable/22300.

Mullins:2002:UDR

[Mul02] Justin Mullins. The ultimate domino rally. New Scientist, ??(2367):??, November 2, 2002. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (elec-

tronic).  URL http://www.
newscientist.com/article/mg17623671.
300-the-ultimate-domino-rally.
html.

Musselman:2006:NCS

[Mus06] Elizabeth Green Musselman. Nervous conditions: science and the body politic in early industrial Britain. SUNY series, studies in the long nineteenth century; SUNY series in science, technology, and society. State University of New York, Albany, NY, USA, 2006. ISBN 0-7914-6679-5 (hardcover). xi + 276 pp. LCCN RC464.A1 G74 2006. URL

http://www.loc.gov/catdir/
toc/ecip0512/2005014025.html.

Metropolis:1972:TEH

[MW72] N. Metropolis and J. Worlton. A trilogy on errors in the history of computing. In AFIPS [AFI72], pages 683–691. LCCN QA76 .U2 1972.

Metropolis:1980:TEH

[MW80] N. Metropolis and J. Worlton. A trilogy on errors in the history of computing. Annals of the History of Computing, 2(1):49–59, January/ March 1980. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1980/pdf/a1049.pdf;   http:
//www.computer.org/annals/
an1980/a1049abs.htm.  An

earlier version is published in [MW72].

Nagler:1980:ANB

[Nag80a] Harry Nagler. Anecdotes:

Napier and Babbage. Annals of the History of Computing, 2(2):186–187, April/ June 1980. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1980/pdf/a2186a.pdf;   http:
//www.computer.org/annals/
an1980/a2186aabs.htm.

Nagler:1980:NB

[Nag80b] Harry Nagler. Napier and Babbage. Annals of the History of Computing, 2(2):186–187, 1980. CODEN AHCOE5. ISSN 0164-1239.

Nicolas:1843:SCA

[Nic43a] Sir Harris Nicolas. Statement of the circumstances attending the invention and construction of Mr. Babbage’s calculating engines. Philosophical Magazine, 23 (series 3)(151):235–239, ???? 1843. CODEN PHMAA4. ISSN 0031-8086. Addition to the memoir of M. Menabrea on the Analytical Engine. Scientific Memoirs, Vol. III, Part XII, p. 666.

Nicolas:1843:SCR

[Nic43b] Sir Harris Nicolas. Statement of the circumstances respecting Mr. Babbage’s calculating engines’, 1843. ?? pp. LCCN ???? Privately printed, London, UK.

Niman:2008:CBI

[Nim08] Neil B. Niman. Charles Babbage’s influence on the development of Alfred Marshall’s theory of the firm. Journal of the History of Economic Thought, 30(4): 479–49, 2008.

Norris:2005:DAI

[NJ05] Sigrid Norris and Rodney H. Jones, editors. Discourse in action: introducing mediated discourse analysis. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2005. ISBN 0-415-36617-8 (paperback), 0-415-35429-3 (hardcover). xii + 229 pp. LCCN P302.84 .D574 2005. URL

http://www.loc.gov/catdir/
enhancements/fy0653/2004023214-
d.html;   http://www.loc.gov/
catdir/toc/ecip051/2004023214.
html.

Nofre:2015:IBB

[Nof15] David Nofre. It Began with Babbage: The Genesis of Computer Science by Subrata Dasgupta (review). Technology and Culture, 56(2):537–538, April 2015. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/
582325.

Norberg:2001:PHC

[Nor01] Arthur L. Norberg. A perspective on the history of the Charles Babbage Institute and the Charles Babbage Foundation. IEEE Annals of the History of Computing, 23(4):12–23, October 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

dlib.computer.org/an/books/
an2001/pdf/a4012.pdf;   http:
//www.computer.org/annals/
an2001/a4012abs.htm.

Nudds:1969:CB

[Nud69] Donald Nudds. Charles Babbage (1791–1871). In Mid-nineteenth century scientists, pages 1–34. Pergamon Press, New York, NY, USA, 1969.

Ng:2006:AEL

[NZ06] Che Tat Ng and Weinian Zhang. An algebraic equation for linear operators. Linear Algebra and its Applications, 412(2–3): 303–325, January 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0024379505003484.

OBrien:1993:BRB

[O’B93] D. P. O’Brien. Book review: Pioneers in Economics 12: Adam Smith (1723–1790), Volume I, by Mark Blaug; Pioneers in Economics 12: Adam Smith (1723– 1790), Volume II, by Mark Blaug; Pioneers in Economics 13: Henry Thornton, Jeremy Bentham, James Lauderdale Simonde de Sismondi, by Mark Blaug; Pioneers in Economics 14: David Ricardo (1772–1823), by Mark Blaug; Pioneers in Economics 15: Jean-Baptiste Say (1776–1832), by Mark Blaug; Pioneers in Economics 16: Thomas Robert Malthus (1766–1834) and John Stuart Mill (1806–1873), by Mark Blaug; Pioneers in Economics 17: Ramsay McCulloch (1789–1864), Nassau Senior (1790–1864), Robert Torrens (1780–1864), by Mark Blaug; Pioneers in Economics 18: Thomas Tooke (1774–1858), Mountifort Longfield (1802–1884), Richard Jones (1790–1855), by Mark Blaug; Pioneers in Economics 19: William Whewell (1794– 1866), Dionysius Lardner (1793– 1859), Charles Babbage (1792– 1871), by Mark Blaug; Pioneers in Economics 20: George Scrope (1797–1876), Thomas Attwood (1783–1856), Edwin Chadwick (1800–1890), John Cairnes (1823–1875); Pioneers in Economics 21: James Mill (1773–1836), John Rae (1796– 1872), Edward West (1782– 1882), Thomas Joplin (1790– 1847), by Mark Blaug; Pioneers in Economics 22: James Wilson (1805–1860), Issac Butt (1813–1879), T. E. Cliffe Leslie (1827–1882), by Mark Blaug; Pioneers in Economics 23: Karl Marx (1818–1883), by Mark Blaug. The Economic Journal, 103(416):233–237, January 1993.

URL http://www.jstor.org/
stable/2234352.

ODonnell:1985:BRB

[O’D85a] John T. O’Donnell. Book review: Charles Babbage: Pioneer of the Computer, by Anthony Hyman. American Mathematical Monthly, 92(7):522–525, August/September 1985. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic).  URL http://www.
jstor.org/stable/2322530.

ODonnell:1985:RCB

[O’D85b] John T. O’Donnell. Reviews:

Charles Babbage: Pioneer of the Computer. American Mathematical Monthly, 92(7):522– 525, 1985. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ogborn:1965:BRB

[Ogb65]  M. E. Ogborn.  Book re-

view: Irascible Genius. A life of Charles Babbage, Inventor, by Maboth Moseley. Journal of the Institute of Actuaries (1886–1994), 91(1):100–102, June 1965. URL http://www.

                                        jstor.org/stable/41139814.

OHanlon:1982:BRB

[O’H82]  Redmond O’Hanlon.  Book

review: A Calculating Man: Charles Babbage, Pioneer of the Computer. By Anthony Hyman. Illustrated. 287 pp. Princeton: Princeton University Press. $25. New York Times, ??(??):BR26, November 14, 1982. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL

                                        http://search.proquest.com/
                                        docview/122073847.

ONeill:1991:HSP

[OLCJ91] Judy O’Neill, J. A. N. Lee, Paul Ceruzzi, and Gail Jennes. Happenings. Sumptuous piece of engineering sculpture. Babbage stamp issued. Fifty years of program-controlled calculating machines. Celebration at the Deutsches Museum. People and computers exhibition opens at the computer museum. Deutsches Museum workshop on the technical history of electrical information technology. International congress of history of science. Technological competitiveness conference, Center for the History of Electrical Engineering. Adelle and Erwin Tomash Fellowship in the history of information processing new book in progress. Annals of the History of Computing, 13 (4):353–358, October/December 1991. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1991/pdf/a4353.pdf;   http:
//www.computer.org/annals/
an1991/a4353abs.htm.

ONeill:1991:RBD

[O’N91] Bill O’Neill. Review: Better described as a ‘great uncle’: Doron Swade, Charles Babbage and his Calculating Engines, (Science Museum, pp 48, £3.50). New Scientist, ??(1785):??, November 1991. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic). URL http://www.

newscientist.com/article/mg13117855.
000-review-better-described-
as-a-great-uncle.html.

Oppenheimer:1971:BRB

[Opp71] Jane Oppenheimer. Book review: Irascible Genius, The Life of Charles Babbage, by Maboth Moseley. Quarterly Review of Biology, 46(2):170, June 1971. CODEN QRBIAK. ISSN 0033-5770 (print), 1539-7718 (elec-

tronic).  URL http://www.
jstor.org/stable/2822076.

ORegan:2012:BHC

[O’R12] Gerard O’Regan. A brief history of computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2. edition, 2012. ISBN 1-4471-2359-X (e-book), 1-4471-2358-1. LCCN ????

Ortiz:2007:BFI

[Ort07a] Eduardo L. Ortiz. Babbage and French idéologie: functional equations, language, and the analytical method. In Jeremy J. Gray and Karen Hunger Parshall, editors, Episodes in the history of modern algebra (1800– 1950), volume 32 of Hist. Math., pages 13–47. American Mathematical Society, Providence, RI, USA, 2007.

Ortiz:2007:LTS

[Ort07b] Eduardo L. Ortiz. Languages, the theory of signs and French “ideology” in the mathematical works of Charles Babbage. Rev. Bras. Hist. Mat., ??(Special Issue 1):117–126, 2007. ISSN 1519-955X.

Ord-Smith:1965:BRB

[OS65] R. J. Ord-Smith. Book review: Irascible Genius (Charles Babbage). The Computer Journal, 7(4):277, January 1965. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

                                        tronic). URL http://comjnl.
                                        oxfordjournals.org/content/
                                        7/4/277.full.pdf+html. See

comments and corrections [Phi65].

Ozgur:2010:CBI

[Ozg10] Erdem Ozgur. Charles Babbage: an inadvertent development economist. History of economic ideas, 18(3):11–31, ????

2010. ISSN 1122-8792 (print), 1724-2169 (electronic).

Padua:2015:TAL

[Pad15] Sydney Padua. The Thrilling Adventures of Lovelace and Babbage. Pantheon Books, New York, NY, USA, 2015. ISBN 0-307-90827-5 (hardcover), 0-307-90828-3 (e-book). 315 pp. LCCN PN6737.P34 T48 2015.

Padua:2017:PLB

[Pad17]  Sydney Padua.  Picturing

Lovelace, Babbage, and the Analytical Engine: a cartoonist in mathematical biography. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(3): 214–220, 2017. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1080/17498430.2017.
1315520.

Panov:1963:BRB

[Pan63]  D. Yu. Panov.  Book re-

view: Charles Babbage and his calculating engines: Morrison, P. and Morrison, E., New York, 1961. U.S.S.R. Computational Mathematics and Mathematical Physics, 3(2):550– 551, ???? 1963. CODEN CMMPA9. ISSN 0041-5553, 0502-9902. URL http://www.

sciencedirect.com/science/
article/pii/0041555363900429.

Park:1996:OHW

[Par96]  Edwards Park.  The object

at hand: What a difference the Difference Engine made: from Charles Babbage’s calculator emerged today’s computer. Smithsonian, 26(11):20– ??, February 01, 1996. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Patterson:1960:FEC

[Pat60] George W. Patterson. The first electric computer, a magnetological analysis. Journal of The Franklin Institute, 270(2): 130–137, August 1960. CO-DEN JFINAB. ISSN 0016-0032 (print), 1879-2693 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0016003260901150.

Payne:1966:BRB

[Pay66] N. P. Payne. Book review: The Analytical Engine, by Jeremy Bernstein. Mathematical Gazette, 50(372):191–192, May 1966. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL http://www.

                                        jstor.org/stable/3611982.

Peacock:1820:CEA

[PHB20] George Peacock, John Frederick William Herschel, and Charles Babbage. A collection of examples of the applications of the differential and integral calculus. Cambridge University Press, Cambridge, UK, 1820. ???? pp. LCCN ????

Phillips:1965:IG

[Phi65] William Phillips. “Irascible genius”. The Computer Journal, 8(1):56, April 1965. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/content/
8/1/56.full.pdf+html.  See

[OS65].

Pilgrim:1995:TTT

[Pil95] Robert A. Pilgrim. TIC-TAC-TOE: introducing expert systems to middle school students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):340–344, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pincock:2006:CHC

[Pin06] Stephen Pincock. Codebreaker:

the history of codes and ciphers, from the ancient pharaohs to quantum cryptography. Walker, New York, NY, USA, 2006. ISBN 0-8027-1547-8. 176 pp. LCCN Z104 .P56 2006.

Plotkin:2012:CSM

[Plo12] Robert Plotkin. Computers in science and mathematics. Computers, internet, and society. Facts on File, New York, NY, USA, 2012. ISBN 0-8160-7757-6. xxiii + 168 pp. LCCN Q183.9 .P56 2012.

Poitou:1982:PBA

[Poi82] Jean-Pierre Poitou. Prony et Babbage: aperçus sur l’histoire de la division du travail mental. (French) [Prony and Babbage: views on the history of the division of mental labor]. History of European Ideas, 3(3):295– 302, ???? 1982. CODEN ????

                                        ISSN ???? URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0191659982900183.

Porter:2007:CBG

[Por07] Dorothy Porter. Charles Babbage and George Birkbeck: science, reform and radicalism. Palgrave Macmillan, Basingstoke, England, 2007. 58–69 pp. LCCN ????

Petrenko:1979:BMO

[PP79] A. K. Petrenko and O. L. Petrenko. The Babbage machine and the origins of programming. Istor.-Mat. Issled., 24(??):340– 360, 389, 1979. CODEN IMA-IAS. ISSN 0136-0949.

Purbrick:1993:DMC

[Pur93] Louise Purbrick. The dream machine: Charles Babbage and his imaginary computers. Journal of Design History, 6(1):9– 21, ???? 1993. ISSN 0952-4649 (print), 1741-7279 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/1315933.

Randell:1973:ODC

[Ran73] Brian Randell, editor. The Origins of Digital Computers: Selected Papers. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., first edition, 1973. ISBN 0-387-06169-X (New York), 3-540-06169-X. xvi + 464 pp. LCCN TK7888.3 .R36.

Randell:1975:ODC

[Ran75] Brian Randell, editor. The Origins of Digital Computers: Selected Papers. Texts and monographs in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1975. ISBN 0-387-07114-8, 3-540-07114-8, 3-642-96244-0, 3-642-96242-4 (e-book). xvi + 464 pp. LCCN ?TK7888.3 .R36 1975.

Randell:1976:C

[Ran76] Brian Randell. The COLOS-SUS. Technical Report 90, Computing Laboratory, University of Newcastle upon Tyne, Newcastle upon Tyne, UK, 1976.

48 pp.  URL http://www.
cs.ncl.ac.uk/publications/
trs/papers/90.pdf;   http:
//www.ncl.ac.uk/computing/
research/publication/160056.

Randell:1979:ABO

[Ran79] Brian Randell. An annotated bibliography on the origins of digital computers. Annals of the History of Computing, 1 (2):101–207, October/December 1979. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1979/pdf/a2101.pdf;   http:
//www.computer.org/annals/
an1979/a2101abs.htm.

Randell:1982:AEE

[Ran82a] Brian Randell. From Analytical Engine to electronic digital computer: The contributions of Ludgate, Torres and Bush. Annals of the History of Computing, 4(4):327– 341, October/December 1982. CODEN AHCOE5. ISSN 0164-1239. URL http://

                                        dlib.computer.org/an/books/
                                        an1982/pdf/a4327.pdf;   http:
                                        //www.computer.org/annals/
                                        an1982/a4327abs.htm.

Randell:1982:ODC

[Ran82b] Brian Randell, editor. The Origins of Digital Computers: Selected Papers. Texts and monographs in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1982. ISBN 0-387-11319-3, 3-540-11319-3. xvi + 580 pp. LCCN TK7885.A5 O741 1982.

Randell:1983:MA

[Ran83]  Brian Randell.  A mysteri-

ous advertisement. Annals of the History of Computing, 5 (1):60–63, January/March 1983. CODEN AHCOE5. ISSN 0164-1239. URL http://

                                        dlib.computer.org/an/books/
                                        an1983/pdf/a1060.pdf;   http:
                                        //www.computer.org/annals/
                                        an1983/a1060abs.htm.

Randell:1991:PBB

[Ran91] Brian Randell. Programming: from Babbage to Backus. Technical report 327, University of Newcastle upon Tyne, Computing Laboratory, Newcastle upon Tyne, UK, 1991. 17 pp.

Rathbone:1952:CBS

[Rat52] Robert R. Rathbone, editor.

Charles Babbage — scientist and philosopher: excerpts describing his work on mechanical computation and his reflections on human behavior, volume R-206 of Paper. Digital Computer Laboratory, Massachusetts Institute of Technology, Cambridge, MA, USA, January 31, 1952. 28 pp. LCCN QA76.5.M41 D57 no.R-206.

Reingold:1968:BMS

[Rei68] Nathan Reingold. Babbage and Moll on the state of science in Great Britain: a note on a document. British Journal for the History of Science, 4 (1):58–64, June 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4024987.

Reinstaller:2007:DLF

[Rei07] Andreas Reinstaller. The division of labor in the firm: agency, near-decomposability and the Babbage Principle. Journal of Institutional Economics, 3(3): 293–322, December 2007.

Richardson:1839:DWV

[Ric39] Charles James Richardson. Description of the Warming and Ventilation Apparatus at the Residence of Charles Babbage. ????, London, UK, 1839. 14 pp. LCCN ????

Richards:1983:BRA

[Ric83] Joan L. Richards. Book review: Anthony Hyman: Charles Babbage: Pioneer of the Computer. Isis, 74(2):292, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic).  URL http://www.
jstor.org/stable/233151.

Rice:1997:IDA

[Ric97] Adrian Rice. Inspiration or desperation? Augustus de Morgan’s appointment to the chair of mathematics at London University in 1828. British Journal for the History of Science, 30(3):257–274, September 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4027861.

Richardson:2002:BO

[Ric02] Ruth Richardson. Babbage’s ophthalmoscope. The Lancet (London, England), 359(9304): 450, February 2, 2002. CO-DEN LANCAO. ISSN 0140-6736 (print), 1474-547X (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0140673602076109.

Rima:2009:DEA

[Rim09] Ingrid Rima. Development of economic analysis. Routledge & Kegan Paul, London, UK and New York, NY, USA, seventh edition, 2009. ISBN 0-415-77293-1, 0-415-77292-3. xii + 606 pp. LCCN HB75 .R46 2009.

Ritt:1916:CRS

[Rit16] J. F. Ritt. On certain real solutions of Babbage’s functional equation. Annals of Mathematics (2), 17(3):113–122, 1916. CO-DEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic).

Robinson:1979:BRJ

[Rob79] Stephen M. Robinson. Book review: J. M. Dubbey: The Mathematical Work of Charles Babbage. Isis, 70(2):326–327, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230856.

Roberts:1987:CQD

[Rob87a] C. J. D. (Jim) Roberts. Comments, queries, and debate: Babbage’s Difference Engine No. 1 and the production of sine tables. Annals of the History of Computing, 9(2): 210–212, July/September 1987. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1987/pdf/a2210.pdf;   http:
//www.computer.org/annals/
an1987/a2210abs.htm.

Roberts:1987:FPU

[Rob87b] C. J. D. (Jim) Roberts. Comments, queries, and debate: Babbage’s Difference Engine No. 1 and the production of sine tables. Annals of the History of Computing, 9(2): 210–212, July/September 1987. CODEN AHCOE5. ISSN 0164-1239. URL http://

                                        dlib.computer.org/an/books/
                                        an1987/pdf/a2210.pdf;   http:
                                        //www.computer.org/annals/
                                        an1987/a2210abs.htm.

Roberts:1993:CQD

[Rob93] C. J. D. Roberts. Comments, queries, and debate: Babbage and Clement. IEEE Annals of the History of Computing, 15(2):6–??, April–June 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        dlib.computer.org/an/books/
                                        an1993/pdf/a2006.pdf.

Roegel:2009:APF

[Roe09] Denis Roegel. Anecdotes: Prototype fragments from Babbage’s first Difference Engine. IEEE Annals of the History of Computing, 31(2):70–75, April/June 2009. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Rojas:2021:CPC

[Roj21] R. Rojas. The computer programs of Charles Babbage. IEEE Annals of the History of Computing, 43(1):6–18, 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Rosenberg:1988:CBC

[Ros88] Nathan Rosenberg. Charles Babbage in a complex world. In Complexity and the history of economic thought: perspectives on the history of economic thought: selected papers from the History of Economics Society Conference, 1998, pages 47–57. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1988. LCCN ????

Rosmaita:1990:BRB

[Ros90] Brian J. Rosmaita. Book review: Science and Reform: Selected Works of Charles Babbage by Anthony Hyman, Charles Babbage. Albion: A Quarterly Journal Concerned with British Studies, 22(4):694–695, Winter

1990. URL http://www.jstor.
org/stable/4051416.

Rosenberg:1994:CBP

[Ros94a] Nathan Rosenberg. Charles Babbage: pioneer economist. In Exploring the black box: technology, economics, and history [Ros94b], chapter 2, page ?? ISBN 0-521-45270-8 (hardback), 0-521-45955-9 (paperback). LCCN T173.8 .R66 1994. URL http://www.

loc.gov/catdir/description/
cam025/93004681.html;   http:
//www.loc.gov/catdir/toc/cam021/
93004681.html.

Rosenberg:1994:EBB

[Ros94b] Nathan Rosenberg. Exploring the black box: technology, economics, and history. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-45270-8 (hardback), 0-521-45955-9 (paperback). ix + 274 pp. LCCN T173.8 .R66 1994. URL http://www.

loc.gov/catdir/description/
cam025/93004681.html;   http:

                                        //www.loc.gov/catdir/toc/cam021/
                                        93004681.html.

Ralston:1993:ECS

[RR93] Anthony Ralston and Edwin D. Reilly, Jr., editors. Encyclopedia of Computer Science and Engineering. Van Nostrand Reinhold, New York, NY, USA, third edition, 1993. ISBN 0-442-27679-6. xxv + 1558 pp. LCCN QA76.15 .E48 1993.

Swade:2002:DEH

[SB02] Doron Swade and Charles Babbage. The Difference Engine: Charles Babbage and the quest to build the first computer. Recording for the Blind and Dyslexic, Princeton, NJ, USA, 2002. ISBN 0-14-200144-9. 1 sound disc + x + 342 + 16 pp. LCCN ????

Smillie:1985:RFM

[SBET85] K. W. Smillie, F. L. Bauer, Ralph Erskine, and Henry S. Tropp. Reviews: O. I. Franksen, Mr. Babbage’s Secret; F. H. Hinsley, British Intelligence in the Second World War; T. M. Thompson, From Error-Correcting Codes Through Sphere Packings to Simple Groups; capsule reviews. Annals of the History of Computing, 7(2):185–191, April/ June 1985. CODEN AHCOE5. ISSN 0164-1239. URL http://

                                        dlib.computer.org/an/books/
                                        an1985/pdf/a2185.pdf;   http:
                                        //www.computer.org/annals/
                                        an1985/a2185abs.htm.

Simoes:2003:TLG

[SCD03] Ana Simões, Ana Carneiro, and Maria Paula Diogo, editors. Travels of learning: a geography of science in Europe, volume 233 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 1-4020-1259-4, 94-017-3584-0. ISSN 0068-0346. xiv + 353 pp. LCCN Q174 .B67 v. 233. URL https:/

/link.springer.com/book/10.
1007/978-94-017-3584-1.

Schweber:1981:ALT

[Sch81] S. S. (Silvan S.) Schweber, editor. Aspects of the life and thought of Sir John Frederick Herschel. The Development of science. Arno Press, New York, NY, USA, 1981. ISBN 0-405-13898-9 (vol. 1), 0-405-13899-7 (vol. 2). ???? pp. LCCN Q113 .H58 1981. US$95.00 (set). Contains a complete facsimile of Babbage, Herschel, and Peacock’s Memoirs of the Analytical Society.

Schwarz:2002:FB

[Sch02] K. K. Schwarz. Faraday and Babbage. Notes and Records of the Royal Society of London, 56 (3):367–381, September 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Schneider:20xx:VLK

[Schxx] Dieter Schneider. Vorläufer der Lehre von den Kernkompetenzen: von Xenophons Kyrupaedie über Babbage’s Prinzip zu Lists “Gesetz der Kraftvereinigung”. (German) [Progenitor of the theory of core competence: from Xenophon Kyrupaedi to Babbage’s principle to channels “Law of force unification”]. ????, ????, 20xx. LCCN ????

Scollon:2005:BBC

[Sco05] Suzie Scollon. Bentham, Babbage and the census. In Norris and Jones [NJ05], page ?? ISBN 0-415-36617-8 (paperback), 0-415-35429-3 (hardcover). LCCN P302.84 .D574 2005. URL

http://www.loc.gov/catdir/
enhancements/fy0653/2004023214-
d.html;   http://www.loc.gov/
catdir/toc/ecip051/2004023214.
html.

Segala:2003:BAE

[Seg03] Marco Segala. Babbage, the Analytical Engine and the Turin Academy of Sciences. In Simões et al. [SCD03], chapter 10, pages 233–247. ISBN 1-4020-1259-4, 94-017-3584-0. ISSN 0068-0346. LCCN Q174 .B67 v. 233.

URL https://link.springer.
com/chapter/10.1007/978-94-
017-3584-1_10.

Segala:2008:BRA

[Seg08]  Marco Segala.  Book re-

view: Andrea Villa (ed.), Charles Babbage. Passaggi dalla vita di uno scienziato. Autobiografia dell’inventore del computer. Torino: Utet, 2007. xiii + 414 pp., ISBN 978-88-02-07715-4. Nuncius, 23(1):184– 185, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.
com/content/10.1163/182539108x00517.

Seidel:1995:CBI

[Sei95] R. W. Seidel. Charles Babbage Institute Director’s column. IEEE Annals of the History of Computing, 17(4):5, Winter 1995. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//ieeexplore.ieee.org/iel4/
85/10129/00477431.pdf.

Sendov:1971:CBO

[Sen71] Blagovest Sendov. Charles Babbage (1791–1871) (on the occasion of the 100th anniversary of his death). Fiz.-Mat. Spis. B′′lgar. Akad. Nauk., 14(47): 331–339, 1971. ISSN 0015-3265.

Severance:2013:JCH

[Sev13] Charles Severance. John C.

Hollar: History of computing. Computer, 46(9):13–14, September 2013. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schneider:2007:ICS

[SGM07] G. Michael Schneider, Judith L.

Gersting, and Keith Miller. Invitation to computer science: Java version. Thomson Course Technology, Boston, MA, USA, third edition, 2007. ISBN 1-4188-3654-0, 1-4239-0143-6. xxiii + 720 pp. LCCN QA76.73.J38 S36 2007.

Shapcott:2010:PU

[Sha10] Thomas W. Shapcott. Parts of us. University of Queensland Press, St Lucia, QLD, Australia, 2010. ISBN 0-7022-3769-8 (paperback). 124 pp. LCCN PR9619.3.S47 P37 2010. US$24.95. URL http://www.

uqp.com.au.

Sheepshanks:1854:LBV

[She54] R. (Richard) Sheepshanks. A letter to the Board of Visitors of the Greenwich Royal Observatory in reply to the calumnies of Mr. Babbage at their meeting in June 1853, and in his book entitled The Exposition of 1851. Printed by G. Barclay, London, UK, 1854. iv + 5–92 pp. LCCN ????

Sheepshanks:1860:LBV

[She60] R. (Richard) Sheepshanks. A letter to the Board of Visitors of the Greenwich Royal Observatory in reply to the calumnies of Mr. Babbage at their meeting in June 1853, and in his book entitled The Exposition of 1851. Printed by G. Barclay, London, UK, 1860. 37 pp. LCCN QB36.S55 A3.

Sherman:2006:CBS

[She06]  Josepha Sherman.  Charles

Babbage and the story of the first computer. Uncharted, unexplored, and unexplained. Mitchell Lane Publishers, Newark, DE, USA, 2006. ISBN 1-58415-372-5. 48 pp. LCCN QA29.B2. URL http:/

/www.gbv.de/dms/bowker/toc/

9781584153726.pdf;   http:/
/www.loc.gov/catdir/toc/ecip052/
2004024613.html.

Strachey:1981:AIP

[SHH81] C. Strachey, Velma R. Huskey, and Harry D. Huskey. Anecdotes: An impossible program; Charles Babbage and Lady Lovelace. Annals of the History of Computing, 3 (4):414–415, October/December 1981. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1981/pdf/a4414.pdf.

Sibbett:1995:BRB

[Sib95] Trevor Sibbett. Book review:

Charles Babbage and the Assurance of Lives by Martin Campbell-Kelly. British Actuarial Journal, 1(3):592, August

1995. URL http://www.jstor.
org/stable/41141237.

Singh:1999:CBE

[Sin99] Simon Singh. The code book:

the evolution of secrecy from Mary, Queen of Scots, to quantum cryptography. Doubleday, New York, NY, USA, 1999. ISBN 0-385-49531-5. xiii + 402 pp. LCCN Z103 .S56 1999. See also [AAG^{+}00].

Slater:1987:WIC

[Sla87] Robert Slater. Who invented the computer? Computers in Physics, 1(1):44–49, November 1987. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

                                        //aip.scitation.org/doi/10.
                                        1063/1.4903432.

Smith:1971:ECI

[Smi71] William D. Smith. The electronic computer’s inventors: Mauchly and Eckert to mark anniversary. New York Times, ??(??):F7, August 1, 1971. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL

                                        http://search.proquest.com/
                                        docview/119265082.

Smillie:1983:BRB

[Smi83]  K. W. Smillie.  Book re-

view: Charles Babbage, Pioneer of the Computer, by Anthony Hyman. SIAM Review, 25(3):422–423, July 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/2029408.

Smith:2006:C

[Smi06] Ludovic Smith. Crash. Corrillium Press, Staines, UK, 2006. ISBN 0-9552634-0-9 paperback. 301 pp. LCCN ????

Smiley:2010:MWI

[Smi10] Jane Smiley. The man who invented the computer: the biography of John Atanasoff, digital pioneer. Doubleday, New York, NY, USA, 2010. ISBN 0-385-52713-6, 0-385-53372-1 (e-book), 1-299-11995-6 (e-book). 246 + 8 pp. LCCN QA76.2.A75 S64 2010. US$25.95.

Snyder:2011:PBC

[Sny11] Laura J. Snyder. The philosophical breakfast club: four remarkable friends who transformed science and changed the world. Broadway Books, New York, NY, USA, 2011. ISBN 0-7679-3048-7. 439 + 4 pp. LCCN ????

Spencer:1999:RAT

[Spe99] Nicholas Spencer. Rethinking ambivalence: Technopolitics and the Luddites in William Gibson and Bruce Sterling’s “The Difference Engine”. Contemporary Literature, 40(3):403–429, Autumn

1999. URL http://www.jstor.
org/stable/1208884.

Spiller:2002:CR

[Spi02]  Neil Spiller.  Cyber reader.

Phaidon, London, UK, 2002. ISBN 0-7148-4071-8 hardcover. 319 pp. LCCN QA76.9.C66 C894 2002.

Scheutz:1857:STC

[SS57] Georg Scheutz and Edvard Scheutz. Specimens of Tables, Calculated, Stereomoulded, and Printed by Machinery. Longman, London, UK, 1857. ???? pp.

Stanoyevitch:2011:ICM

[Sta11] Alexander Stanoyevitch. Introduction to cryptography with mathematical foundations and computer implementations. Discrete mathematics and its applications. Chapman and Hall/CRC, Boca Raton, FL, USA, 2011. ISBN 1-4398-1763-4 (hardcover). xix + 649 pp. LCCN QA268 .S693 2011.

Stein:1984:LLN

[Ste84]  Dorothy K. Stein.  Lady

Lovelace’s notes: Technical text and cultural context. Victorian Studies, 28(1):33–67, Autumn 1984. ISSN 0042-5222 (print), 1527-2052 (electronic).

URL http://www.jstor.org/
stable/3826758.

Steel:1992:BD

[Ste92] D. I. Steel. Babbage’s dates. Nature, 356(6371):653, April 23, 1992. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/
v356/n6371/pdf/356653b0.pdf.

Comments that Babbage’s birthdate was 26 December 1791, even though Babbage himself long believed it was 1792.

Stevenson:1995:CSP

[Ste95] Richard W. Stevenson. Christie’s sells part of a pre-computer. New York Times, ??(??):D4, October 9, 1995. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL

http://search.proquest.com/
docview/109507261.

Stigler:1991:CBI

[Sti91] George J. Stigler. Charles Babbage (1791 + 200 = 1991 ). Journal of Economic Literature, 29 (3):1149–1152, September 1991. CODEN JECLB3. ISSN 0022-0515 (print), 1547-1101 (elec-

tronic).  URL http://www.
jstor.org/stable/2727614.

Stokes:1855:RCA

[Sto55] George G. Stokes. Report of a Committee appointed by the Council to examine the calculating machine of M. Scheutz. Proceedings of the Royal Society, VII (7):499–509, ???? 1855.

Spufford:1996:CBT

[SU96] Francis Spufford and Jennifer S.

Uglow, editors. Cultural Babbage: technology, time and invention. Faber and Faber, London, UK, 1996. ISBN 0-571-17242-3, 0-571-17243-1 (paperback). 313 pp. LCCN CB478 .C85 1996. 0571172431 (pbk)L.15.99: Formerly CIP.

Sussman:1994:CMC

[Sus94] Herbert Sussman. Cyberpunk meets Charles Babbage: “The Difference Engine” as alternative Victorian history. Victorian Studies, 38(1):1–23, Autumn 1994. ISSN 0042-5222 (print), 1527-2052 (electronic).

URL http://www.jstor.org/
stable/4618879.

Sussman:2000:BRB

[Sus00a] Herbert Sussman. Book review:

Machine dreams: The culture of technology. Victorian Literature and Culture, 28(1):197– 204, ???? 2000. ISSN 1060-1503 (print), 1470-1553 (elec-

tronic).  URL http://www.
jstor.org/stable/25058498.

Sussman:2000:BRM

[Sus00b] Herbert Sussman. Book review: Machine dreams: The culture of technology: The Economy of Machinery and Manufactures. Victorian Literature and Culture, 28(1):197–204, ???? 2000. ISSN 1060-1503 (print), 1470-1553 (elec-

tronic).  URL http://www.
jstor.org/stable/25058498.

Sussman:2011:BRS

[Sus11] Herbert Sussman. Book review: Steampunk at Oxford: Steampunk, by Art Donovan. Steampunk Neuromancer, by William Gibson. The Difference Engine by William Gibson and Bruce Sterling. League of Extraordinary Gentlemen, by Stephen Norrington. The War Lord of the Air by Michael Moorcock. League of Extraordinary Gentlemen, Vol. 1, by Alan Moore and Kevin O’Neill; The Thrilling Adventures of Lovelace and Babbage by Sydney Padua, The Museum of the History of Science. Victorian Literature and Culture, 39(1): 278–284, ???? 2011. ISSN 1060-1503 (print), 1470-1553 (electronic).

Swade:1989:UCM

[Swa89] Doron Swade. The unerring certainty of mechanical agency: Machines and table making in the Nineteenth Century. In Campbell-Kelly et al. [CKCFR03], pages 143–174. ISBN 0-19-850841-7 (print). LCCN ????

Swade:1991:BRW

[Swa91a] Doron Swade. Book review: The world reduced to number: The Works of Charles Babbage. Isis, 82(3):532–536, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic).  URL http://www.
jstor.org/stable/233232.

Swade:1991:BBD

[Swa91b] Doron Swade. Building Babbage’s dream machine. New Scientist, ??(1775):37–39, June 1991. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

Swade:1991:CBH

[Swa91c] Doron Swade. Charles Babbage and his calculating engines. Science Museum, London, UK, 1991. ISBN 0-901805-45-9. xii + 48 pp. LCCN QA76.2.B3 S8 1991; QA76.2 .B33 S93 1991.

Swade:1991:CCB

[Swa91d] Doron Swade. The construction of Charles Babbage’s difference engine. Annals of the History of Computing, 13(1): 82–83, January/March 1991. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1991/pdf/a1081.pdf;   http:
//www.computer.org/annals/
an1991/a1081abs.htm.

Swade:1993:RCB

[Swa93]  Doron D. Swade.  Redeem-

ing Charles Babbage’s mechanical computer. Scientific American, 268(2):86–91, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/
scientificamerican/journal/
v268/n2/pdf/scientificamerican0293-
86.pdf.

Swade:1995:PBH

[Swa96a] Doron Swade. Piece of Babbage history sold. Resurrection: The Computer Conservation Society Journal, ??(14): ??, Winter 1995–1996. ISSN 0958-7403. URL https://

computerconservationsociety.
org/resurrection/res14.htm#
c.

Swade:1996:CBD

[Swa96b] Doron Swade. Charles Babbage’s Difference Engine No. 2: Technical Description, volume 5 of Science Museum Papers in the History of Technology. Science Museum, London, UK, 1996. 147 + 22 pp.

Swade:2000:CBC

[Swa00a] Doron Swade. The Cogwheel Brain: Charles Babbage and the Quest to Build the First Computer. Little, Brown and Co., Boston, MA, USA, 2000. ISBN 0-316-64847-7. x + 342 + 16 pp. LCCN QA75 .S93 2000. URL http://hdl.

library.upenn.edu/1017.12/
366197;   http://www.gbv.de/
dms/bsz/toc/bsz086755331inh.
pdf.

Swade:2000:DEC

[Swa00b] Doron Swade. The Difference Engine: Charles Babbage and the quest to build the first computer. Penguin Books, Harmondsworth, UK, 2000. ISBN 0-14-200144-9 (paperback). x + 342 + 16 pp. LCCN QA75 .S954 2002.

Swade:2001:CBC

[Swa01a] Doron Swade. The cogwheel brain: Charles Babbage and the quest to build the first computer. Abacus, London, UK, 2001. ISBN 0-349-11239-8 (paperback). 352 + 16 pp. LCCN QA75 .S954 2001.

Swade:2001:DEC

[Swa01b] Doron Swade. The Difference Engine: Charles Babbage and the Quest to Build the First Computer. Viking, New York, NY, USA, 2001. ISBN 0-670-91020-1, 0-14-200144-9 (paperback). x + 342 + 16 pp. LCCN QA75 .S954 2001. URL http:/

/www.gbv.de/dms/bowker/toc/
9780670910205.pdf;   http:/
/www.loc.gov/catdir/toc/fy034/
00068643.html;   http://www.
zentralblatt-math.org/zmath/
en/search/?an=1017.01006.

Swade:2001:EN

[Swa01c] Doron Swade. The enchantress of number. In Swade-DE [Swa01b], chapter 8, pages 155– 171. ISBN 0-670-91020-1, 0-14-200144-9 (paperback). LCCN QA75 .S954 2001. URL http:/

/www.gbv.de/dms/bowker/toc/
9780670910205.pdf;   http:/

                                        /www.loc.gov/catdir/toc/fy034/
                                        00068643.html;   http://www.
                                        zentralblatt-math.org/zmath/
                                        en/search/?an=1017.01006.

Swade:2002:DEC

[Swa02] Doron Swade. The Difference Engine: Charles Babbage and the quest to build the first computer. Penguin Books, Harmondsworth, UK, 2002. ISBN 0-14-200144-9 (paperback). x + 342 + 16 pp. LCCN QA75 .S954 2002.

Swade:2003:CTC

[Swa03] Doron Swade. Calculation and Tabulation in the 19th Century: George Biddell Airy versus Charles Babbage. Ph.D. thesis, University College London, London, UK, 2003. 353 pp.

                                        URL https://www.proquest.
                                        com/pqdtglobal/docview/1752660629/
                                        1261A6EDAC064B26PQ/1.

Swade:2004:STA

[Swa04] Doron Swade. The shocking truth about Babbage and his calculating engines. Resurrection: The Computer Conservation Society Journal, ??(32): ??, New Year 2004. ISSN 0958-7403. URL https://

                                        computerconservationsociety.
                                        org/resurrection/res32.htm#
                                        d.

Swade:2005:CCB

[Swa05] Doron D. Swade. The construction of Charles Babbage’s Difference Engine No. 2. IEEE Annals of the History of Computing, 27 (3):70–88, July/September 2005.

CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Swade:2009:BC

[Swa09] Doron Swade. Babbage, Charles (1791–1871). In Oxford Dictionary of National Biography, volume 3, page ???? Oxford University Press, Walton Street, Oxford OX2 6DP, UK, May

21, 2009. URL https://www.
oxforddnb.com/view/10.1093/
ref:odnb/9780198614128.001.
0001/odnb-9780198614128-e-
962.

Swade:2010:ACC

[Swa10] Doron D. Swade. Automatic computation: Charles Babbage and computational method. Rutherford Journal, 3(??):??, ???? 2010. CODEN ???? ISSN 1177-1380. URL http://

rutherfordjournal.org/article030106.
html.

Swade:2011:PPA

[Swa11a] Doron Swade. Pioneer profile:

Ada Lovelace. Resurrection: The Computer Conservation Society Journal, ??(53):31–33, 2011. ISSN 0958-7403. URL https://

computerconservationsociety.
org/resurrection/res53.htm#
g.

Swade:2011:PEC

[Swa11b] Doron Swade. Pre-electronic computing. In Jones and Lloyd [JL11], pages 58–83. ISBN 3-642-24540-4 (paperback), 3-642-24541-2 (e-book). LCCN QA76.17. URL https://

                                        link.springer.com/chapter/
                                        10.1007/978-3-642-24541-1_
                                        7.

Swade:2013:ODC

[Swa13] Doron Swade. Origins of digital computing: Alan Turing, Charles Babbage, and Ada Lovelace. In Zenil [Zen13], pages 23–43. ISBN 981-4374-29-6. LCCN QA267.7 .C676 2013. URL http:

                                        //www.worldscientific.com/
                                        worldscibooks/10.1142/8306.

Foreword by Roger Penrose.

Swade:2017:TLB

[Swa17] Doron Swade. Turing, Lovelace, and Babbage. In Copeland et al. [CBSW17], chapter 24, pages 249–262. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Sydell:2009:CMF

[Syd09] Laura Sydell. A 19th-Century mathematician finally proves himself: Charles Babbage’s ‘Difference Engine’. NPR Web site, December 10, 2009.

                                        URL https://www.npr.org/
                                        templates/story/story.php?
                                        storyId=121206408.

Taylor:1992:CBM

[Tay92] Nicholas K. Taylor. Charles Babbage’s mini-computer—Difference Engine No. 0. Bulletin of the Institute of Mathematics and its Applications, 28(6-8):112–114, 1992. CODEN IMTABW. ISSN 0905-5628.

Tee:1982:CBM

[Tee82] Garry J. Tee. Charles Babbage materials in New Zealand and Australia. Historia Mathematica, 9(3):344–345, August 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0315086082901306.

Tee:1983:HCB

[Tee83] Garry J. Tee. The heritage of Charles Babbage in Australasia. Annals of the History of Computing, 5(1):45–59, January/ March 1983. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1983/pdf/a1045.pdf;   http:
//www.computer.org/annals/
an1983/a1045abs.htm.

Tee:1984:CBH

[Tee84] Garry J. Tee. Charles Babbage (1791–1871) and his New Zealand connections. Roy. Soc. New Zealand Bull., 21(??):81–90, 1984. ISSN 0370-6559.

Tee:1986:CBH

[Tee86] Garry J. Tee. Charles Babbage (1791–1871) and his New Zealand connections. New Zealand Math. Mag., 22(3):112– 123, 1986. ISSN 0549-0510.

Tee:1994:MAC

[Tee94]  G. J. Tee.  More about

Charles Babbage’s difference engine no. 0. Bulletin of the Institute of Mathematics and its Applications, 30(9-10):134–137, 1994. CODEN IMTABW. ISSN 0905-5628.

Tee:1995:BP

[Tee95] Garry Tee. Babbage’s pocket. New Scientist, ??(1985):??, July 8, 1995. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic). URL http://

www.newscientist.com/article/
mg14719856.500-babbages-pocket.
html.

Thimbleby:2000:CNB

[Thi00] Harold Thimbleby. Calculators are needlessly bad. International Journal of Human-Computer Studies, 52(6):1031– 1069, June 2000. CODEN ????

ISSN ???? URL http://www.
sciencedirect.com/science/
article/pii/S1071581999903415.

Tomayko:1988:AAT

[THWV88] James E. Tomayko, Peter Hilton, Richard Louis Weis, and Alfred Van Sinderen. Anecdotes: Alan Turing in the Home Guard; Overcoming Murphy’s Law; Babbage and the Scheutz Machine at Dudley Observatory. Annals of the History of Computing, 10(2):133–139, April/ June 1988. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1988/pdf/a2133b.pdf;   http:
//www.computer.org/annals/
an1988/a2133babs.htm.

Tomas:2002:IHL

[Tom02] David Tomas. On the imagination’s horizon line: Uchronic histories, protocybernetic contact, and Charles Babbage’s calculating engines. In Clarke and Henderson [CH02], pages 217–234. ISBN 0-8047-4176-X, 0-8047-4210-3 (paperback). LCCN T15 .F78 2002. URL

http://www.loc.gov/catdir/
enhancements/fy0710/2002070594-
b.html;   http://www.loc.gov/
catdir/enhancements/fy0710/
2002070594-d.html;   http:/
/www.loc.gov/catdir/toc/fy033/
2002070594.html.

Tricomi:1972:PDM

[Tri72] Francesco Giacomo Tricomi. Un precursore delle moderne macchine calcolatrici: Charles Babbage (1792–1871). Atti Accad. Sci. Torino Cl. Sci. Fis. Mat. Natur., 106:17–24, 1972. ISSN 0001-4419.

Tropp:1984:CBP

[Tro84] Henry S. Tropp. Charles Babbage: Pioneer of the Computer by Anthony Hyman (review). Technology and Culture, 25(3): 651–653, July 1984. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL http://www.jstor.org/
stable/3104217;   https://
muse.jhu.edu/pub/1/article/
890005/pdf.

Tukey:1972:HCS

[Tuk72] John W. Tukey. How computing and statistics affect each other. In The Babbage Memorial Meeting: Report of Proceedings, pages 21–37. British Computer Society and Royal Statistical Society, London, UK, 1972.

Turvey:1991:SJH

[Tur91] Peter J. Turvey. Sir John Herschel and the abandonment of Charles Babbage’s Difference Engine No. 1. Notes and Records of the Royal Society of London, 45(2):165–176, July 1991. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Tropp:1987:ABE

[TW87] Henry S. Tropp and Maurice V. Wilkes. Anecdotes: Babbage’s Expectations for the Difference Engine. Annals of the History of Computing, 9(2):203–204, April/ June 1987. CODEN AHCOE5. ISSN 0164-1239. URL http://

                                        dlib.computer.org/an/books/
                                        an1987/pdf/a2203.pdf;   http:
                                        //www.computer.org/annals/
                                        an1987/a2203abs.htm.

Tweedale:1989:BRB

[Twe89a] Geoffrey Tweedale. Book review: The Works of Charles Babbage, by Martin Campbell-Kelly, Charles Babbage. British Journal for the History of Science, 22(4):481–482, December 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4026950.

Tweedale:1989:BRM

[Twe89b] Geoffrey Tweedale. Book review: Martin Campbell-Kelly (ed.). The Works of Charles Babbage. London: Pickering & Chatto (Publishers) Ltd, 1989. 11 vols. ISBN 1-85196-005-8. £500. British Journal for the History of Science, 22(4):481– 482, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4026950.

Tweedale:1991:BRM

[Twe91] Geoffrey Tweedale. Book review: Michael Lindgren. Glory and Failure: The Difference Engines of Johann Müller, Charles Babbage and Georg and Edvard Scheutz, translated by Craig C. McKay. Cambridge, Mass, and London: MIT Press, 1990. Pp. 414. ISBN 0-262-12146-8. £40.50. British Journal for the History of Science, 24(2): 261–263, June 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4027184.

Tweedale:1992:BRJ

[Twe92] Geoffrey Tweedale. Book review:

John Napier, Rabdology, translated by W. F. Richardson, introduction by R. E. Rider. Charles Babbage Institute Series for the History of Computing, 15. Cambridge, Mass, and London: MIT Press/Los Angeles and San Francisco: Tomash Publishers, 1990. Pp. xxxvii + 135. ISBN 0-262-14046-2. £35.95. British Journal for the History of Science, 25(4):462–463, December 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4027054.

Tweedale:1993:BRE

[Twe93] Geoffrey Tweedale. Book review: Ernst Martin, The Calculating Machines (Die Rechenmaschinen): Their History and Development, translated and edited by Peggy Aldrich Kidwell and Michael R. Williams. Volume 16 in the Charles Babbage Institute Reprint Series for the History of Computing. Cambridge, Mass, and London: MIT Press; Los Angeles and San Francisco: Tomash Publishers, 1992. Pp. xvii + 367, illus. ISBN 0-262-13278-8. £44.95. British Journal for the History of Science, 26 (1):126–127, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4027388.

Uglow:1996:CBT

[Ugl96] Jennifer S. Uglow, editor. Cultural Babbage: technology, time, and invention. Faber and Faber, London, UK, 1996. ISBN 0-571-17242-3, 0-571-17243-1 (paperback). 313 pp. LCCN CB478 .C85 1996.

Vamos:1985:BRB

[Vam85] T. Vamos. Book review: Mr. Babbage’s secret. The tale of a cypher — and APL: Ole Immanuel Franksen. Automatica, 21(5):616, September 1985. CODEN ???? ISSN

????  URL http://www.
sciencedirect.com/science/
article/pii/0005109885900135.

VanSinderen:1980:PPC

[Van80] Alfred W. Van Sinderen. The printed papers of Charles Babbage. Annals of the History of Computing, 2(2):169–185, April/ June 1980. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1980/pdf/a2169.pdf;   http:
//www.computer.org/annals/
an1980/a2169abs.htm.

VanSinderen:1983:BLQ

[Van83] Alfred W. Van Sinderen. Babbage’s letter to Quetelet, May 1835. Annals of the History of Computing, 5(3):263– 267, July/September 1983. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1983/pdf/a3263.pdf;   http:
//www.computer.org/annals/
an1983/a3263abs.htm.

VanSinderen:1983:RHC

[VBCC83] Alfred W. Van Sinderen, Allan G. Bromley, Bernard Cohen, and Paul Ceruzzi. Reviews: A. Hyman: Charles Babbage; N. Stern: from ENIAC to UNIVAC I; C. Baum: The System Builders; capsule reviews. Annals of the History of Computing, 5(1):76–85, January/ March 1983. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1983/pdf/a1076.pdf;   http:

                                        //www.computer.org/annals/
                                        an1983/a1076abs.htm.

VanSinderen:1991:HBM

[VW91] Alfred W. Van Sinderen and Michael R. Williams. Happy birthday Mr. Babbage. Annals of the History of Computing, 13(2):125–139, April/ June 1991. CODEN AHCOE5. ISSN 0164-1239. URL http://

                                        dlib.computer.org/an/books/
                                        an1991/pdf/a2125.pdf;   http:
                                        //www.computer.org/annals/
                                        an1991/a2125abs.htm.

W:1962:BRB

[W.62] J. W. W. Book review: Charles Babbage and His Calculating Engines, by Philip Morrison, Emily Morrison. Mathematics of Computation, 16(80):509, October 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

                                        jstor.org/stable/2003156.

W:1986:BRB

[W.86] J. W. W. Book review: Charles Babbage, Pioneer of the Computer, by Anthony Hyman. Mathematics of Computation, 46 (174):759, April 1986. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                        URL http://www.jstor.org/
                                        stable/2008013.

Watson:2012:UMH

[Wat12] Ian Watson. The Universal Machine: From the Dawn of Computing to Digital Consciousness. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-28101-X, 3-642-28102-8 (print). xiv + 353 pp. LCCN QA75.5-76.95.

Wilmot-Buxton:1976:MLL

[WB76] H. J. (Harry John) Wilmot-Buxton, editor. Memoir of the life and labours of Charles Babbage: a descriptive and historical account of his analytical and difference engines. Museum of the History of Science, Oxford, UK, 1976. 13 pp. LCCN QA75.B112 A3 1976.

Weiss:1988:BOP

[Wei88] Eric A. Weiss. Biographies:

Oh, pioneers! Annals of the History of Computing, 10 (4):348–361, October/December 1988. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1988/pdf/a4348.pdf;   http:
//www.computer.org/annals/
an1988/a4348abs.htm.

Weld:1989:ECH

[Wel89] C. R. Weld. The eleventh chapter of the history of the Royal Society (1849). In Campbell-Kelly [CK89i], pages 149–163. ISBN 1-85196-510-6, 1-85196-005-8 (set). LCCN ????

Wendling:2009:KMT

[Wen09] Amy E. Wendling. Karl Marx on technology and alienation. Palgrave Macmillan, Basingstoke, UK, 2009. ISBN 0-230-22440-7 (hardback). x + 252 pp. LCCN B3305.M74 W45 2009.

Wilkes:1964:BDA

[Wil64] Maurice V. Wilkes. Babbage’s dream in action: [book reviews: Charles Babbage and His Calculating Engines, edited by Philip Morrison and Emily Morrison, New York: Dover, London: Constable, 1961]. Nature, 202(4933): 631–632, May 16, 1964. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

                                        URL http://www.nature.com/
                                        nature/journal/v202/n4933/
                                        pdf/202631a0.pdf.

Wilkes:1975:HBD

[Wil75] Maurice V. Wilkes. How Babbage’s dream came true. Nature, 257(5527):541–544, October 16, 1975. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

                                        URL http://www.nature.com/
                                        nature/journal/v257/n5527/
                                        pdf/257541a0.pdf.

Williams:1976:DE

[Wil76] M. R. Williams. The Difference Engines. The Computer Journal, 19(1):82–89, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

                                        tronic). URL http://comjnl.
                                        oxfordjournals.org/content/
                                        19/1/82.full.pdf+html.

Wilkes:1977:BCP

[Wil77] M. V. Wilkes. Babbage as a computer pioneer. Historia Mathematica, 4(4):415–440, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0315086077900799.

Williams:1980:SLC

[Wil80] M. R. Williams. The scientific library of Charles Babbage. Report 80/46/4, University of Calgary, Calgary, AB, Canada, 1980.

Wilkes:1981:DCU

[Wil81a] M. V. Wilkes. The design of a control unit — reflections on reading Babbage’s notebooks. Annals of the History of Computing, 3(2):116–120, April/ June 1981. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1981/pdf/a2116.pdf;   http:
//www.computer.org/annals/
an1981/a2116abs.htm.

Williams:1981:SLC

[Wil81b] M. R. Williams. The scientific library of Charles Babbage. Annals of the History of Computing, 3(3):235–240, July/September 1981. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1981/pdf/a3235.pdf;   http:
//www.computer.org/annals/
an1981/a3235abs.htm.

Williams:1985:HCT

[Wil85] Michael R. (Michael Roy)

Williams. A history of computing technology. Prentice-Hall series in computational mathematics. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1985. ISBN 0-13-389917-9. xi + 432 pp. LCCN QA71 .W661 1985.

Williams:1986:BRBa

[Wil86] M. R. Williams. Book review: Mr. Babbage’s Secret: The Tale of a Cypher — and APL by Ole Immanuel Franksen. Isis, 77(1):156, March 1986. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/232545.

Williams:1987:BBT

[Wil87] Michael R. Williams. Babbage and Bowditch: a transatlantic connection. Annals of the History of Computing, 9 (3/4):283–290, July/September 1987. CODEN AHCOE5. ISSN 0164-1239. URL http://

                                        dlib.computer.org/an/books/
                                        an1987/pdf/a3283.pdf;   http:
                                        //www.computer.org/annals/
                                        an1987/a3283abs.htm.

Williams:1988:BBT

[Wil88] Michael R. Williams. Babbage and Bowditch: a transatlantic connection. Annals of the History of Computing, 9(3-4):283– 290, 1988. CODEN AHCOE5. ISSN 0164-1239.

Wilkes:1990:HPB

[Wil90] M. V. Wilkes, F.R.S. Herschel, Peacock, Babbage and the development of the Cambridge curriculum. Notes and Records of the Royal Society of London, 44 (2):205–219, July 1990. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Wilkes:1991:BCP

[Wil91a] Maurice V. Wilkes. Babbage as a computer pioneer. British Computer Society and the Royal Statistical Society, ??(??):??, ???? 1991.

Wilkes:1991:BEH

[Wil91b] Maurice V. Wilkes. Babbage’s expectations for his Engines. Annals of the History of Computing, 13(2):141–145, April/ June 1991. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1991/pdf/a2141.pdf;   http:
//www.computer.org/annals/
an1991/a2141abs.htm.

Wilkes:1991:PMB

[Wil91c]  Maurice V. Wilkes.  Pray

Mr. Babbage — a character study in dramatic form. Annals of the History of Computing, 13(2):147–154, April/ June 1991. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1991/pdf/a2147.pdf;   http:
//www.computer.org/annals/
an1991/a2147abs.htm.

Wilkes:1992:CPC

[Wil92a] Maurice V. Wilkes. Computing perspectives: Charles Babbage: The Great Uncle of computing? Communications of the ACM, 35 (3):15–16, 21, March 1992. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.
acm.org/pubs/toc/Abstracts/
0001-0782/214839.html.

Williams:1992:BJC

[Wil92b] M. R. Williams. Biographies — Joseph Clement: the first computer engineer. IEEE Annals of the History of Computing, 14 (3):69–76, July/September 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). Joseph Clement was Charles Babbage’s superb machinist who worked for years on building Babbage’s Difference Engine.

Wilkes:1995:BRU

[Wil95] Maurice V. Wilkes. Book review: The ubiquitous Mr Babbage: Passages from the Life of a Philosopher. Notes and Records of the Royal Society of London, 49(2):341–342, July 1995. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Williams:1997:HCT

[Wil97] Michael R. Williams. A History of Computing Technology. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, second edition, 1997. ISBN 0-8186-7739-2 (hardcover). xi + 426 pp. LCCN QA76.17 .W55 1997.

                                        URL http://catalogue.bnf.
                                        fr/ark:/12148/cb37514972q.

Williams:1998:LWC

[Wil98] M. R. (Michael R.) Williams. The “Last word” on Charles Babbage. IEEE Annals of the History of Computing, 20 (4):10–14, October/December 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:/

/ieeexplore.ieee.org/iel4/
85/15706/00728225.pdf.

Wilkes:2000:CBC

[Wil00a] M. V. Wilkes. The cogwheel brain: Charles Babbage and the quest to build the first computer [book review]. IEEE Annals of the History of Computing, 22(4):62–63, October–December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:/

/ieeexplore.ieee.org/iel5/
85/19196/00887990.pdf.

Wilkes:2000:CQD

[Wil00b] Maurice Wilkes. Comments, queries, and debates: Was Babbage caught in the act? IEEE Annals of the History of Computing, 22(4):69– 70, October–December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

dlib.computer.org/an/books/
an2000/pdf/a4069.pdf.

Wilkes:2000:IBA

[Wil00c] Maurice V. Wilkes. Introduction to “Babbage’s Analytical Engine Plans 28 and 28a — The Programmer’s Interface” by Allan Bromley. IEEE Annals of the History of Computing, 22(4):4–5, October–December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

                                        dlib.computer.org/an/books/
                                        an2000/pdf/a4004.pdf;   http:
                                        //ieeexplore.ieee.org/iel5/
                                        85/19196/00887985.pdf;   http:
                                        //www.computer.org/annals/
                                        an2000/a4004abs.htm.

Wilkes:2002:CBH

[Wil02] Maurice V. Wilkes. Charles Babbage and his world. Notes and Records of the Royal Society of London, 56(3):353–365, September 22, 2002. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/3557738.

Wolfram:2016:IMP

[Wol16] Stephen Wolfram. Idea makers: personal perspectives on the lives and ideas of some notable people. Wolfram Media, Inc., Champaign, IL, USA, 2016. ISBN 1-57955-003-7 (hardcover), 1-57955-005-3 (ebook), 1-57955-011-8. 250 (est.) pp. LCCN Q141 .W678562

                                        2016.  URL http://www.
                                        wolfram-media.com/products/
                                        idea-makers.html.

Woolley:1999:BSR

[Woo99] Benjamin Woolley. The bride of science: romance, reason, and Byron’s daughter. Mc-Graw-Hill, New York, NY, USA, 1999. ISBN 0-07-137329-2. viii + 416 pp. LCCN QA29.L72 W66 1999. URL

                                        http://www.loc.gov/catdir/
                                        bios/mh041/2001274424.html;
                                         http://www.loc.gov/catdir/
                                        description/mh021/2001274424.

html;   http://www.loc.gov/
catdir/toc/mh021/2001274424.
html.

Wortmann:1987:BRB

[Wor87] J. C. Wortmann. Book review: Mr. Babbage’s secret: The tale of a Cypher — and APL: Strandberg, Birkerød, (Denmark) 1984, 319 pages. European Journal of Operational Research, 29(2):216, May 1987. CO-DEN EJORDT. ISSN 0377-2217 (print), 1872-6860 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0221787901184.

Wrottesley:1856:PLS

[Wro56] Lord Wrottesley. Proceedings of learned societies: Royal Society: “Report of a Committee appointed by the Council to examine the Calculating Machine of M. Scheutz”. Philosophical Magazine, 11 (series 4)(75 (supplement)):540– 548, July 1856. CODEN PH-MAA4. ISSN 0031-8086. URL

http://hdl.handle.net/2027/
uc1.31822016397200;   http:
//www.tandfonline.com/doi/
abs/10.1080/14786445608642120.

The committee consisted of Professors George G. Stokes, W. H. Miller, Charles Wheatstone, and Rev. R. Willis.

Williams:1988:MAI

[WSC^{+}88] M. R. Williams, Oliver B. R.

Strimpel, Paul Ceruzzi, William Jacobs, Doron Swade, Joachim Fisher, W. Aspray, Joyce E.

Bedi, Nancy R. McGovern, Geoffrey Tweedale, Walter Carlson, and Pierre E. Mounier-Kuhn. Museums and archives: Introduction; The Computer Museum, Boston; Beyond the Limits: A New Gallery at the National Air and Space Museum, Smithsonian Institution, Washington, DC; The Information Age: Visions and Realities; Deutsches Museum, Munich Computer Science and Automation Hall; Charles Babbage Institute: Center for the History of Information Processing; The IEEE Center for the History of Electrical Engineering; Records in the National Archives Relating to the Early Involvement of the U.S. Government in Data Processing, 7880s to 7950s; British National Archive for the History of Computing; AFIPS History of Computing Committee; History of Computing in France. Annals of the History of Computing, 10 (4):305–329, October/December 1988. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1988/pdf/a4305.pdf;   http:
//www.computer.org/annals/
an1988/a4305abs.htm.

Weiss:1992:BEA

[WWCP92] Eric A. Weiss, Michael R. Williams, John W. Carr III, and Noah Prywes. Biographies: Eloge: Arthur Lee Samuel (1901– 1990); Joseph Clement: The first computer engineer; Saul Gorn (1913–1992). Annals of the History of Computing, 14 (3):55–77, July/September 1992. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1992/pdf/a3055.pdf;   http:
//www.computer.org/annals/
an1992/a3055abs.htm.

Yntema:1977:CB

[Ynt77] L. Yntema. Charles Babbage (1792/1871). Verzekerings-Arch., 54(3):189–205, 1977.

Yost:2002:BGR

[Yos02] Jeffrey R. Yost. A Bibliographic Guide to Resources in Scientific Computing, 1945–1975, volume 15 of Bibliographies and indexes in library and information science, 0742-6879. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 2002. ISBN 0-313-31681-3. 263 pp. LCCN Z7405.D37 Y67 2002; Q183.9. URL

http://www.loc.gov/catdir/
toc/fy035/2002069622.html.

Yost:2021:CBI

[Yos21] J. R. Yost. Charles Babbage Institute for Computing, Information, and Culture: a major interdisciplinary turn. IEEE Annals of the History of Computing, 43(1):92–93, 2021. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Zenil:2013:CUU

[Zen13] Hector Zenil, editor. A computable universe: understanding and exploring nature as computation. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2013. ISBN 981-4374-29-6. xliv + 810 pp. LCCN QA267.7 .C676 2013. URL http:

                                        //www.worldscientific.com/
                                        worldscibooks/10.1142/8306.

Foreword by Roger Penrose.